### v2

# Fourier dimension at $C^{1+\alpha}$ regularity

Abstract.
We prove power Fourier decay for equilibrium states of nonlinear $C^{1+\alpha}$ iterated function systems, giving first examples of truly $C^{1+\alpha}$ IFSs with attractors of positive Fourier dimension that are not conjugated to a linear IFS previously out of reach of existing techniques. The key new ingredient is a Measure Non-Linearity (MNL) criterion tailored to $C^{1+\alpha}$ IFSs, coupling oscillations in geometric and symbolic directions and allowing us to prove the necessary multiscale non-linearity condition for Fourier decay via an induction-on-scales scheme. This bypasses the classical obstructions that require either large dimension enabling stationary phase comparisons (Kaufman measures), or smoothness or cohomological conditions on the roof function $\tau$ such as $\tau\in C^{1}$ (Dolgopyat’s method), $\tau\in C^{\alpha}$ but with smooth underlying dynamics (Tsujii–Zhang strategy), or non-flat $C^{\alpha}$ conjugacy to a constant (fractal van der Corput lemma). The method also applies to parabolic systems including limit sets of geometrically finite groups, Manneville-Pommeau maps, and Lorenz-type maps.
The authors are all supported by the Research Council of Finland’s Academy Research Fellowship “Quantum chaos of large and many body systems”, grant Nos. 347365, 353738.
1. Introduction
1.1. Background and summary of the main results
The study of Fourier decay of dynamically defined measures has seen rapid progress in recent years. Given a probability measure $\mu$ on $\mathbb{R}$, we are interested in the high-frequency asymptotics of its Fourier transform
$\widehat{\mu}(\xi)=\int e^{-2\pi i\xi x}\,d\mu(x),\qquad\xi\in\mathbb{R},$
with particular emphasis on singular measures $\mu$ arising in hyperbolic and parabolic dynamics, fractal geometry, and geometric group theory. Examples include self-conformal measures, equilibrium states for uniformly and non-uniformly hyperbolic maps and flows, and Patterson–Sullivan measures for isometric group actions generating fractal limit sets. Fourier decay for such stationary measures for chaotic dynamical systems plays a central role in several areas, including bounding resonance free regions for scattering resonances in quantum chaos [13, 22], exponential mixing of flows and the decay of correlations (e.g. [31, 45]), equidistribution and Diophantine properties of numbers when applied to Gauss map [39], dispersive estimates for quasiperiodic Schrödinger operators [26] and stable intersections and homoclinic tangencies [7]. We refer to [40] for a broader survey on the recent developments.
Most of the literature on Fourier decay in dynamical settings has focused on systems with at least $C^{2}$ regularity or analyticity, where powerful tools such as Dolgopyat’s method [21] apply. This method, and its later refinements, exploits smooth holonomies and the structure of hyperbolic codings with finite alphabet to establish exponential mixing and Fourier decay [25, 13, 35, 2, 4, 41, 8, 10, 27, 28]. However, many natural dynamical systems of interest have only $C^{1+\alpha}$ regularity, such as codings of hyperbolic diffeomorphisms on basic sets due to $C^{1+\alpha}$ holonomies [44, 26], or exhibit parabolic features leading to infinite alphabet codings, including Manneville–Pommeau maps [18], Bunimovich billiards [16, 34, 14], geodesic flows on surfaces of non-positive curvature and Lorenz flows [33, 19], or hyperbolic surfaces with cusps [31].
The recent breakthrough of Tsujii and Zhang [44] on exponential mixing of 3D Anosov flows extended Dolgopyat’s method to rougher holonomies via new UNI conditions, and was adapted in [26] to prove power Fourier decay for equilibrium states of various Axiom $A$ diffeomorphisms, e.g. ones coming from the trace map of the Fibonacci Hamiltonian. However, these methods still fundamentally depend on the underlying smoothness of the dynamics (as there the roof function is smooth and one compares the temporal distance function to polynomials) and do not apply to equilibrium states of general $C^{1+\alpha}$ IFSs where both the roof and temporal distance are not smooth.” See Section 1.2 for further discussion comparing to [44, 26].
In this paper we introduce an induction-on-scales approach at the $C^{1+\alpha}$ regularity threshold that allows us to prove power Fourier decay for truly $C^{1+\alpha}$ IFSs. The method relies on verifying a multiscale nonlinearity condition Quantitative Non-Linearity (QNL) which yields power Fourier decay for oscillatory integrals with slowly varying amplitudes and phases (Theorem 1.1). In $C^{2}$ regularity, this is already more or less well-established route originating from the work of Bourgain and Dyatlov [13] and series of works after that [30, 41, 27, 10, 9, 26]. However, the main novelty of our work are the ways we verify QNL in rough settings where existing techniques do not apply. We develop two complementary routes which both base on induction-on-scales.
First, in the genuinely $C^{1+\alpha}$, non-conjugated IFS setting we introduce Measure Non-Linearity (MNL), a condition that couples symbolic and geometric oscillations; using an induction-on-scales tree scheme that avoids spectral gaps for twisted transfer operators, we show MNL implies QNL in the finite-branch case (Theorem 1.3). Then via a fixed-point argument on a moduli space of IFSs that may be of independent interest, we construct explicit truly $C^{1+\alpha}$ examples with power Fourier decay (Theorem 1.4). We expect this approach to apply more broadly in low-regularity dynamical systems. Second, in smoother categories we prove that uniform non-integrability (UNI) implies QNL (Theorem 1.2) via a $C^{\alpha}$ tree argument rather than Dolgopyat estimates. Because the framework is compatible with Young towers, this yields parabolic applications such as Manneville–Pommeau and Lorenz models, and via Li–Pan’s coding [31] gives Fourier decay for Patterson–Sullivan measures on geometrically finite hyperbolic surfaces with cusps (Theorems 1.5 and 1.6). Let us now describe our results in detail in the following Sections 1.2 and 1.3.2.
1.2. Results in $C^{1+\alpha}$ regularity
To state our results, we will need some notation from symbolic dynamics and thermodynamical formalism of subshifts of iterated function systems. The framework is that of equilibrium states for (possibly infinitely branched) uniformly hyperbolic $C^{1+\alpha}$ Markov systems with bounded distortions. This includes finite-branch $C^{1+\alpha}$ hyperbolic systems (by allowing equilibrium states that assign zero mass to the tail), and it is chosen specifically so as to also cover parabolic examples via Young tower constructions and codings that represent a finite-branch parabolic system as an infinite-branch hyperbolic IFS.
Symbolic coding. Let $\{I_{a}\}_{a\in\mathcal{A}}$ be pairwise disjoint open intervals in $[0,1]$, where $\mathcal{A}$ is finite or countable, and set $I=\bigcup_{a\in\mathcal{A}}I_{a}$. Suppose $T:\overline{I}\to\mathbb{R}$ is a $C^{1}$ map with the Markov property: whenever $T(I_{b})\cap I_{a}\neq\emptyset$, we have $T(\overline{I_{b}})\supset\overline{I_{a}}$. Define the transition matrix $A$ by $A(a,b)=1$ if $T(\overline{I_{a}})\supset\overline{I_{b}}$ and $A(a,b)=0$ otherwise. A finite (or infinite) word $\mathbf{a}=a_{1}\cdots a_{n}$ is admissible if $A(a_{i},a_{i+1})=1$ for all $i$. For each admissible pair $ab$ we write $g_{ab}:I_{b}\to I_{a}$ for the inverse branch; the collection $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$ is a (Markov) IFS. The symbolic space is $\Sigma^{\mathbb{N}}:=\{\mathbf{a}\in\mathcal{A}^{\mathbb{N}}\mid\forall j\in\mathbb{N}:A(a_{j},a_{j+1})=1\}$; we denote $\Sigma^{n}$ the set of admissible words of length $n$ and $\Sigma^{*}=\bigcup_{n\geq 1}\Sigma^{n}$. We define the notation $\mathbf{a}\to x$ to simply mean that $g_{\mathbf{a}}$ is defined at $x$. By denoting the last letter of $\mathbf{a}$ with $b(\mathbf{a})$, the condition $\mathbf{a}\to x$ satisfies $\mathbf{a}\to x\Longleftrightarrow x\in I_{b\left(\mathbf{a}\right)}$.
We assume:
  • (1)
    Uniform $\kappa_{+}$-contraction: there exists $\kappa_{+}\in(0,1)$ such that for every admissible $ab\in\Sigma^{2}$ there is $\kappa_{-}(ab)\in(0,1)$ with
    $\kappa_{-}(ab)\leq|g_{ab}^{\prime}(x)|\leq\kappa_{+}\quad\text{for all }x\in I_{b}.$
  • (2)
    Bounded $(C,\alpha)$-distortion: there exist $C>0$ and $\alpha>0$ such that, for all admissible $ab$ and all $x\neq y\in I_{b}$,
    $\big|\log|g_{ab}^{\prime}(x)|-\log|g_{ab}^{\prime}(y)|\big|\leq C|x-y|^{\alpha}.$
We say that $\Phi$ is $C^{1+\alpha}$ regular. When $\Phi$ is $C^{2}$, the bounded $(C,1)$-distortion follows from the Adler condition $\big\|\tfrac{T^{\prime\prime}}{(T^{\prime})^{2}}\big\|_{\infty}<\infty$ by the mean value theorem. Uniform contraction guarantees the existence of a unique attractor $F\subset I$, namely
$F=F_{\Phi}=\bigcap_{n\geq 1}\ \bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}},\qquad I_{\mathbf{a}}:=g_{\mathbf{a}}(I_{b(\mathbf{a})}).$
Gibbs measures and equilibrium states. We work with Gibbs measures $\mu_{\phi}$ supported on $F$. A function $\phi:F\to\mathbb{R}$ has summable variations if
$\sum_{n\geq 2}\ \sup_{\mathbf{a}\in\Sigma^{n}}\ \sup_{x,y\in I_{\mathbf{a}}}|\phi(x)-\phi(y)|<\infty.$
A potential $\phi$ with summable variations is normalized if $\phi<0$ and $1$ is an eigenfunction of the transfer operator
$\mathcal{L}_{\phi}f(x):=\sum_{a\in\mathcal{A},\,ab\in\Sigma^{2}}e^{\phi(g_{ab}(x))}\,f(g_{ab}(x)),\qquad x\in I_{b},$
for all continuous $f:F\to\mathbb{C}$; we write $w_{ab}(x):=e^{\phi(g_{ab}(x))}$. The associated Gibbs measure $\mu_{\phi}$ is the unique probability measure on $F$ satisfying the $\phi$-conformality:
$\int_{F}f(x)\,d\mu_{\phi}(x)\;=\;\sum_{ab\in\Sigma^{2}}\int_{I_{b}}e^{\phi(g_{ab}(x))}\,f(g_{ab}(x))\,d\mu_{\phi}(x)$
for all continuous $f$. In this setting $\mu_{\phi}$ is non-atomic and $T:F\to F$ is ergodic with respect to $\mu_{\phi}$.
Light tail (infinite alphabet). To cover parabolic systems (hence infinite alphabets), we prevent $\mu_{\phi}$ from concentrating on words with exceptionally small derivatives:
  • (3)
    Light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail: for some $\gamma_{0}>0$,
    $C_{\phi,\gamma_{0}}:=\sum_{ab\in\Sigma^{2}}\ \sup_{x\in I_{b}}e^{\phi(g_{ab}(x))}\,|g_{ab}^{\prime}(x)|^{-\gamma_{0}}<\infty.$
Non-linearity and temporal distance. The key input for Fourier decay (when the $g_{\mathbf{a}}^{\prime}$ are not constant) is a statistical non-concentration property for the distribution of derivatives along the symbolic tree. For $\mathbf{a},\mathbf{b}\in\Sigma^{n}$ and $x,y\in F$, define
$\Delta_{n}\big((\mathbf{a},x),(\mathbf{b},y)\big):=\big(\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{b}}^{\prime}(x)\big)-\big(\log g_{\mathbf{a}}^{\prime}(y)-\log g_{\mathbf{b}}^{\prime}(y)\big),$
which is a finite-time analogue of the temporal distance function associated to the roof $\tau:=-\log T^{\prime}$:
$\Delta_{\infty}\big((w,x),(v,y)\big):=\sum_{n=0}^{\infty}\Big(\tau(g_{w_{1}\cdots w_{n}}x)-\tau(g_{v_{1}\cdots v_{n}}x)\Big)-\Big(\tau(g_{w_{1}\cdots w_{n}}y)-\tau(g_{v_{1}\cdots v_{n}}y)\Big),$
so that $\Delta_{\infty}((w,x),(v,y))=\lim_{n\to\infty}\Delta_{n}((w|n,x),(v|n,y))$. In the flow picture (e.g. [44]), fluctuations of temporal distance govern mixing rates; in the present IFS setting the words $\mathbf{a},\mathbf{b}$ mimic the stable direction while $x,y$ play the role of the unstable direction under the expansion of $T$. To capture non-linearity at low regularity we couple motion in both symbolic and geometric directions. We assume the following multiscale incidence bound we call Quantitative Non-Linearity (QNL): there exist $\rho,\Theta\in(0,1)$ such that for every interval $I\subset\mathbb{R}$ and $n\in\mathbb{N}$,
$\sum_{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\in\Sigma^{n}}\mu(I_{\mathbf{a}})\mu(I_{\mathbf{b}})\mu(I_{\mathbf{c}})\mu(I_{\mathbf{d}})\,\mathbf{1}_{I}\!\left(\Delta_{n}\big((\mathbf{a},x_{\mathbf{c}}),(\mathbf{b},x_{\mathbf{d}})\big)\right)\ \lesssim\ |I|^{\Theta}+\rho^{n},$
where $x_{\mathbf{a}},x_{\mathbf{b}},x_{\mathbf{c}},x_{\mathbf{d}}$ denote the centers of $I_{\mathbf{a}},I_{\mathbf{b}},I_{\mathbf{c}},I_{\mathbf{d}}$. Under QNL we obtain power decay for oscillatory integrals with slowly varying amplitudes and phases.
Theorem 1.1.
Let $\Phi$ be a $C^{1+\alpha}$ IFS satisfying (1) uniform $\kappa_{+}$-contraction, (2) bounded $(C,\alpha)$-distortion, and let $\phi$ be a normalized potential satisfying the light-tail condition (3). Let $\mu_{\phi}$ be the associated Gibbs measure. If (QNL) holds, then there exists $\beta>0$ such that
$\Big|\int\chi(x,\xi)\,e^{i\psi(x,\xi)}\,d\mu_{\phi}(x)\Big|\ \lesssim\ |\xi|^{-\beta},\qquad|\xi|\to\infty,$
for every pair $\chi:I\times\mathbb{R}\to\mathbb{R}$ and $\psi:I\times\mathbb{R}\to\mathbb{R}$ with
$\|\chi(\cdot,\xi)\|_{C^{\alpha}}\lesssim|\xi|^{\varepsilon_{\chi}},\qquad\|\log\partial_{x}\psi(\cdot,\xi)\|_{C^{\alpha}}\lesssim|\xi|^{1+\tilde{\varepsilon}_{\psi}},\qquad\inf_{x}|\partial_{x}\psi(x,\xi)|\gtrsim|\xi|^{1-\varepsilon_{\psi}},$
for some sufficiently small $\varepsilon_{\chi},\varepsilon_{\psi},\tilde{\varepsilon}_{\psi}>0$.
Theorem 1.1 follows ultimately similar lines as the works of Bourgain-Dyatlov [13] and several subsequent works [30, 41, 27, 10, 9] in higher generality. However, the multiscale QNL condition in Theorem 1.1 is in general very difficult to verify, it requires often quite delicate non-concentration analysis, Diophantine or spectral gap assumptions (see e.g. [40] for details and historical aspects), and these themselves are the main culprit that require smoothness from the dynamics. The main novelty of our work, which we discuss now, comes from the induction-on-scales scheme that allows us to prove QNL condition for systems that also involve truly $C^{1+\alpha}$ systems and in smoother categories completely avoiding the use of spectral gaps of twisted transfer operators.
First of all, in the $C^{2}$ category (finite or infinite branch), prior works such as [41, 27, 10, 5] employ the following (UNI) condition going back to Chernov and Dolgopyat [21]:
  • (UNI)
    $(c_{0},\alpha)$-Uniform Non-Integrability: $\tau\in C^{1+\alpha}$ and there exists $c_{0}>0$ such that for infinitely many $N\in\mathbb{N}$ and every $c\in\mathcal{A}$ there are $\mathbf{a},\mathbf{b}\in\Sigma^{N}$ with $b(\mathbf{a})=b(\mathbf{b})=c$ and, for all $x\neq y\in F\cap I_{c}$,
    $|X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)|\ \geq\ c_{0}\,|x-y|^{\alpha},$
where
$X_{\mathbf{a}}(x,y):=S_{n}\tau\circ g_{\mathbf{a}}(x)-S_{n}\tau\circ g_{\mathbf{a}}(y)=\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{a}}^{\prime}(y),\qquad\mathbf{a}\to x,y,$
measures the spatial distortions of $g_{\mathbf{a}}$. (Typically checked with $\alpha=1$, but the arguments extend formally to $\alpha\in(0,1]$.) Under (UNI) we give a proof of QNL:
Theorem 1.2.
If the IFS has a finite or countable set of branches, then (UNI) implies (QNL).
In the finite-branch case, Theorem 1.2 appears, e.g., in [41, 10, 30, 27, 5], via spectral gaps for twisted transfer operators derived from Dolgopyat estimates. Here to showcase the method also in smoother category, we give a proof avoiding Dolgopyat’s estimate, instead building a $C^{\alpha}$ version of the tree strategy previously employed by the first author in the smoother setting [28] using an induction-on-scales scheme. Theorem 1.2 is new for parabolic IFSs and underpins our later applications to Patterson–Sullivan measures in the parabolic case and intermittent systems such as Manneville–Pommeau maps, where UNI holds. We also note that for analytic (finite-branch) IFSs, power Fourier decay is already guaranteed by the existence of a branch with non-vanishing second derivative, by combining [5, 10, 1, 8] (though only for the specific phase $\psi(x,\xi)=x\xi$).
While $(c_{0},\alpha)$–(UNI) is formally stated for $C^{1+\alpha}$, verifying a uniform lower bound of order $|x-y|^{\alpha}$ on the entire attractor is hard without at least $C^{2}$ smoothness. For truly $C^{1+\alpha}$ systems one needs a condition that leverages the fractal geometry of $F_{\Phi}$ and couples symbolic and geometric directions (words $\mathbf{a},\mathbf{b}$ and points $x,y$). This is inspired by the work of Tsujii and Zhang [44], where the unstable and stable directions are coupled to capture the non-linearity. However, the main difference with [44, 26] is that in our setting there is no smooth dynamics we can rely on. We therefore introduce:
  • (MNL)
    Measure Non-Linearity for $\Phi$: Let $\delta:=\dim_{H}F_{\Phi}$, let $\mu$ be the equilibrium state for the potential $\delta\tau$, and assume $\tau\in C^{\delta}$. There exists $C_{\text{(MNL)}}$ such that, for all $N\geq 1$, there are $\mathbf{a},\mathbf{b}\in\mathcal{A}^{N}$ with
    $\forall x_{0}\in F_{\Phi},\ \forall I\subset\mathbb{R},\quad\big(X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})\big)_{*}\mu(I)\ \leq\ C_{\text{(MNL)}}\,|I|.$
Using an induction-on-scales scheme, we then show that MNL indeed yields QNL in the finite-branch setting:
Theorem 1.3.
If the $C^{1+\alpha}$ IFS has finitely many branches, then (MNL) implies (QNL).
Thus by Theorem 1.1, MNL suffices to obtain power Fourier decay for the conformal measure $\mu$ on $F_{\Phi}$. We now construct $C^{1+\alpha}$ IFSs satisfying MNL, yielding fractal equilibrium states with power Fourier decay. As far as we know, these are the first examples of $C^{1+\alpha}$ IFSs not conjugated to linear IFSs admitting power Fourier decay.
The construction proceeds via fixed point theory in the moduli space of IFSs parametrized by equilibrium states, ensuring the required non-concentration to implement the proof of Theorem 1.1. The key input is the existence of a branch whose log-derivative has $\alpha$–Hölder lower regularity, as furnished by a Cantor-type function with jumps precisely on the attractor.
Let us first define the space of Frostman measures $\mu$ in [1,1][-1,1] for given constants $C\geq 1$ and $\varepsilon>0$ by
$\mathcal{M}_{C,\varepsilon}:=\Big\{\mu\in\mathcal{P}([-1,1])\ :\ \mu(I)\leq C|I|^{\varepsilon}\ \text{for all intervals }I\text{ with }|I|\leq\tfrac{1}{100}\Big\}.$
Given $\mu\in\mathcal{M}_{C,\varepsilon}$, we can define an IFS $\Phi_{\mu}$ depending on $\mu$ with branches
$g_{\mu,1}(x):=\kappa\int_{0}^{x}\exp(\mu(-\infty,t])\,dt+b_{1},\qquad g_{\mu,0}(x):=\kappa_{0}x+b_{0},\qquad x\in[0,1],$
where $0<\kappa_{0}\leq\kappa<1/10$ and $b_{0},b_{1}\in\mathbb{R}$ are chosen so that $\Phi_{\mu}=\{g_{\mu,0},g_{\mu,1}\}$ satisfies the strong separation condition. By the Frostman property, $\Phi_{\mu}$ is $C^{1+\alpha}$ for any $\alpha\leq\varepsilon$. Let $T_{\mu}$ be the associated expanding map on $F_{\Phi_{\mu}}$ and write $\tau_{\mu}(x):=-\log|T_{\mu}^{\prime}(x)|$ as the $C^{\alpha}$ roof function.
Theorem 1.4.
For any $C>0$ and $\varepsilon>0$, there exist $\mu\in\mathcal{M}_{C,\varepsilon}$ and $0<\delta_{\mu}<1$ such that the Gibbs measure associated to the potential $\delta_{\mu}\tau_{\mu}$ for the (truly) $C^{1+\varepsilon}$ IFS $\Phi_{\mu}$ has power Fourier decay.
The proof performs a fixed-point analysis for $\mu$–controlled” transfer operators
$\mathcal{L}_{\mu}f(x):=\sum_{i}|g^{\prime}_{\mu,i}(x)|^{\delta_{\mu}}\,f(g_{\mu,i}(x)),$
which depend on the equilibrium state $\mu$, to produce measures with the Measure Non-Linearity (MNL) property; Theorem 1.1 then applies. Let us now discuss on the MNL condition and prospects of the theory.
Remark.
  • (1)
    MNL property and autosimilar temporal distance function. The Cantor-type structure of $\tau$ in our examples of $C^{1+\alpha}$ IFSs was chosen precisely to verify the Measure Non-Linearity condition. In light of recent progress on exponential mixing for 33D flows [44] and Axiom $A$ maps [26], it is natural to ask what axiomatic properties of $\tau$ are sufficient to guarantee Quantitative Non-Linearity. More concretely, let $F_{\mu}(x):=\mu(-\infty,x)$ denote the Cantor staircase associated to $\mu$. Since $(F_{\mu})_{*}\mu$ is Lebesgue when $\mu$ is non-atomic, one expects MNL to hold at least when $\tau$ behaves like a Cantor staircase on $F_{\Phi}$; this motivates the example in Theorem 1.4. Generally, MNL quantifies the spread of the values $X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})$ on $F_{\Phi}$. Under (MNL) one can show the following variant of UNI: for all $N\geq 1$, there exist $\widehat{\mathbf{a}},\widehat{\mathbf{b}}\in\mathcal{A}^{N}$ such that for all $\mathbf{c}\in\mathcal{A}^{*}$ and all $y\in I_{\mathbf{c}}$ there exists $x\in I_{\mathbf{c}}$ with
    $|X_{\widehat{\mathbf{a}}}(x,y)-X_{\widehat{\mathbf{b}}}(x,y)|\ \geq\ \frac{1}{3C_{\text{(MNL)}}}\,\mu(I_{\mathbf{c}})\qquad\text{(MNL–UNI)}.$
    This is slightly weaker than $((3C_{\text{(MNL)}})^{-1},\delta)$–UNI, and better adapted to truly $C^{1+}$ settings. The non-emptiness of (MNL) suggests prospects for a $C^{1+}$ theory of power Fourier decay and Fractal Uncertainty Principles in the spirit of [13]. It is natural to seek Fourier decay for $C^{1+}$ IFS under UNI-type conditions inspired by (MNL–UNI), and to explore $\tau$ with autosimilar oscillations (e.g. Brownian/Weierstrass-type) which could be used as basis for an axiomatic definition needed for QNL.
  • (2)
    Spectral gaps and Fourier integral operators in low regularity. The motivation for proving Fourier decay is closely tied to its spectral implications, including scattering resonances [23, 13]. Since Theorem 1.1 yields Fourier decay bounds for general phases and slowly growing amplitudes, it is interesting to explore how these methods interact with the spectral theory of transfer and Fourier integral operators in low regularity. In particular, combining our results with ideas from [29] could pave the way to establishing spectral gap properties for twisted transfer operators associated to the $C^{1+\alpha}$ IFSs constructed in Theorem 1.4.
  • (3)
    On quantitative aspects. The decay exponent $\beta$ can be traced explicitly to the exponent in the discretized sum–product theorem 3.2. This depends on a non-concentration exponent that can be written in terms of the (QNL) constants, which in turn can be expressed via the (MNL) or (UNI) constants and the light-tail exponent $\gamma_{0}$. Developing the multiscale induction in the proof of (QNL) under (MNL) or (UNI) is an interesting problem to improve the Fourier decay rates.
  • (4)
    Overlapping IFSs. Our current approach relies on symbolic coding under a separation condition. However, previous work has shown that Fourier decay can also be obtained in overlapping settings via a random Dolgopyat operator method and co-cycle versions of Dolgopyat’s argument [10, 5]. It seems plausible that a variant of the induction-on-scales scheme here (Section 6) could be adapted to handle such overlaps, providing a complementary path to Fourier decay beyond separation.
  • (5)
    Theorem 1.3 is stated for finitely many branches. Extending it to countably many branches (to directly address parabolic examples) is an interesting problem, and would likely require adapting the proof of Theorem 1.2 to infinite alphabets.
Finally, let us remark on the opposite case when the system is conjugated to a linear IFS. The analytic and $C^{2}$ settings (e.g. [5, 10, 8, 1, 11, 3]), power Fourier decay is typically classified through the presence of non-linear branches with non-vanishing second derivative or non-flatness condition of the map [1]. In our rough $C^{1+\alpha}$ framework, autosimilarity of the distortion function might play the role of the second derivative giving the necessary non-flatness property needed in [1]. Establishing such a replacement would move the theory toward a genuine characterisation of power Fourier decay at the $C^{1+\alpha}$ regularity threshold.
1.3. Results in parabolic dynamics
Since Theorem 1.1 applies whenever (QNL) holds, and (QNL) follows from (UNI) in the $C^{2}$ category by Theorem 1.2, our tree strategy avoids using spectral gaps of twisted transfer operators (as in [41, 27, 10, 5]) to obtain power Fourier decay. This framework therefore extends many known results to parabolic systems without having to establish spectral gap estimates. The interface we establish between Fourier–decay estimates and the inducing scheme appears to be new, and it may have further applications in parabolic dynamics in relation to spectral theory; it also treats phases and amplitudes that exhibit slow frequency growth.
In Section 1.3.1, using the coding of Li–Pan [31], we obtain Fourier dimension bounds for Patterson–Sullivan measures on geometrically finite hyperbolic surfaces with cusps. In Section 1.3.2, we treat more general parabolic situations modeled by Young towers, focusing on conformal measures for Manneville–Pommeau maps; the same method applies to other induced systems under analogous distortion and non-integrability assumptions.
1.3.1. Patterson–Sullivan measures on parabolic limit sets
[…]
$\int_{\Lambda_{\Gamma}}f(x)\,d\mu_{\Gamma}(x)=\int_{\Lambda_{\Gamma}}f(\gamma(x))\,|\gamma^{\prime}(x)|_{\mathbb{S}^{1}}^{\,\delta}\,d\mu_{\Gamma}(x),$
where $|\gamma^{\prime}(x)|_{\mathbb{S}^{1}}=\frac{1+x^{2}}{1+\gamma(x)^{2}}\,|\gamma^{\prime}(x)|$.
Theorem 1.5.
Let $\Gamma<\mathrm{PSL}(2,\mathbb{R})$ be torsion-free, geometrically finite, with parabolic elements. Then there exists $\beta>0$ such that
$\Big|\int\chi(x,\xi)\,e^{i\psi(x,\xi)}\,d\mu_{\Gamma}(x)\Big|\ \lesssim\ |\xi|^{-\beta},\qquad|\xi|\to\infty,$
for any $\chi:I\times\mathbb{R}\to\mathbb{R}$ and $\psi:I\times\mathbb{R}\to\mathbb{R}$ with
$\|\chi(\cdot,\xi)\|_{C^{\alpha}}\lesssim|\xi|^{\varepsilon_{\chi}},\quad\|\log\partial_{x}\psi(\cdot,\xi)\|_{C^{\alpha}}\lesssim|\xi|^{\tilde{\varepsilon}_{\psi}},\quad\inf_{x}|\partial_{x}\psi(x,\xi)|\gtrsim|\xi|^{1-\varepsilon_{\psi}},$
for some sufficiently small $\varepsilon_{\chi},\varepsilon_{\psi},\tilde{\varepsilon}_{\psi}>0$.
This follows from Li and Pan [31, Prop. 4.1] together with Theorem 1.1. Indeed Li-Pan construct a symbolic coding in which $\mu_{\Gamma}$ is a Gibbs measure with a light-tail property for a subshift of finite type of an IFS $\Phi$ satisfying (1)–(3) in our setting, and $\Lambda_{\Gamma}$ coincides with the attractor $F_{\Phi}$. More precisely, there exists a fundamental interval $I\subset\dot{\mathbb{R}}$ for the parabolic fixed point at $\infty$ (denoted $\Delta_{\infty}$ in [31, §2.3]), constants $C_{1}>0$, $\kappa_{+}\in(0,1)$, $\gamma_{0}>0$, a countable family of disjoint open sets $\{I_{a}\}_{a\in\mathcal{A}}\subset I$, and an expanding map $T:\bigcup_{a}I_{a}\to I$ such that:
  • (1)
    $\sum_{a\in\mathcal{A}}\mu_{\Gamma}(I_{a})=\mu_{\Gamma}(I)$;
  • (2)
    for each $a$ there exists $\gamma_{a}\in\Gamma$ with $I_{a}=\gamma_{a}(I)$ and $T|_{I_{a}}=\gamma_{a}^{-1}$;
  • (3)
    (uniform contraction) $|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for $x\in I$;
  • (4)
    (bounded distortion) $|\partial_{x}\log|\gamma_{a}^{\prime}(x)||<C_{1}$ on $I$;
  • (5)
    (finite exponential moments) with $\tau(x)=\log|DT(x)|$, one has $\int e^{\gamma_{0}\tau}\,d\mu_{\Gamma}<\infty$ for some $\gamma_{0}>0$.
Thus $\Phi=\{\gamma_{a}:a\in\mathcal{A}\}$ is an IFS with uniform $\kappa_{+}$–contraction and bounded $(C_{1},1)$–distortion, and [31] shows
$C_{\phi,\gamma_{0}}=\sum_{a\in\mathcal{A}}\sup_{x\in I}|\gamma_{a}^{\prime}(x)|^{\delta-\gamma_{0}}<\infty.$
With $\phi(x):=-\delta\log|\gamma_{a}^{\prime}(x)|_{\mathbb{S}^{1}}$ for $x\in I_{a}$, equivariance yields that $\mu_{\Gamma}$ is a Gibbs measure for $\phi$, and the $(C_{\phi,\gamma_{0}},\gamma_{0})$–tail condition holds. Finally, Li–Pan prove a UNI statement [31, Lemma 4.5]: writing $\Phi^{n}=\{\gamma_{\mathbf{a}}:\mathbf{a}\in\mathcal{A}^{n}\}$ and $\tau_{n}(x)=\sum_{k=0}^{n-1}\tau(T^{k}x)$, there exist $r,c_{0}>0$ such that for all large $N$ there are words $\{\gamma_{mj}\}_{m=1,2;\,1\leq j\leq j_{0}}$ with
$\big|\partial_{z}\big(\tau_{N}(\gamma_{1j}(z))-\tau_{N}(\gamma_{2j}(z))\big)\big|_{z=y}\big|\ \geq c_{0},\qquad\forall\,y\in B(x,r),\ \forall\,x\in\Lambda_{\Gamma}\cap\overline{I},$
which implies $(c_{0},1)$–UNI by the mean value theorem. Therefore Theorem 1.1 applies, proving Theorem 1.5. The decay exponent in Theorem 1.5 can be traced explicitly in terms of the contraction/distortion data, the UNI constants, and the tail exponent (all available from [31]).
Remark (FUP and resonances). Bourgain–Dyatlov [13] used Fourier decay to prove a Fractal Uncertainty Principle (FUP) for convex co-compact groups ($\delta_{\Gamma}<1/2$), with consequences for the distribution of scattering resonances. In the parabolic case one has $\delta_{\Gamma}\geq 1/2$, and the strategy in [13] does not apply directly. It would be interesting to investigate whether Theorem 1.5 has implications for the meromorphic continuation of the resolvent on geometrically finite manifolds with cusps (cf. [24]).
1.3.2. Parabolic non-linear systems
Our axioms in Theorem 1.1 were designed with Young tower models in mind, obtained by inducing parabolic systems such as Lorenz maps (arising from the Lorenz flow) [33, 19] and, in higher dimensions, the Chernov–Markarian–Zhang structure for slowly mixing billiards (e.g. the Bunimovich stadium) [16, 34, 14]. We give two applications: Manneville–Pommeau maps and Lorenz maps; the assumptions can be verified in broader induced settings.
  • (a)
    Manneville–Pommeau maps. We use the Liverani–Saussol–Vaienti parametrization [32]:
    $T_{\alpha}(x)=\begin{cases}x\big(1+2^{\alpha}x^{\alpha}\big),&x<\tfrac{1}{2},\\ 2x-1,&x\geq\tfrac{1}{2},\end{cases}$
    with $T^{\prime}_{\alpha}(0)=1$ (parabolic fixed point). Let $f_{0},f_{1}$ be the inverse branches (with $f_{1}$ the left, nonlinear one, and $f_{0}(x)=(x+1)/2$), set $x_{0}=1$, $x_{1}=\tfrac{1}{2}$, $x_{n+1}=f_{1}(x_{n})$, and $I=[\tfrac{1}{2},1]$, $I_{n}=[x_{n+1},x_{n})$. This gives a Markov partition of $0,1]=T_{\alpha}(I)$ with $I_{n-1}=T_{\alpha}(I_{n})$. Let $R:I\to\mathbb{N}$ be the first return time to $I$, and define the induced map $T_{A}:I\to I$ by $T_{A}(x)=T_{\alpha}^{R(x)}(x)$. Then $T_{A}$ is an infinitely branched expanding Markov map with inverse branches $g_{a}:I\to A_{a}$, $a\geq 0$ (where $A_{0}=[3/4,1]$ and $A_{a}:=(T_{\alpha}^{-1}I_{a})\cap I$).
  • (b)
    Lorenz maps. Consider the interval map $f:J\setminus\{0\}\to J$, J=[1/2,1/2]J=[-1/2,1/2], defined by [6, §3]
    $f(x)=\begin{cases}-a(-x)^{\alpha}+\tfrac{1}{2},&x<0,\\ ax^{\alpha}-\tfrac{1}{2},&x>0,\end{cases}\qquad\alpha:=-\lambda_{3}/\lambda_{1}\in(0,1),\ 0<a\leq 2^{\alpha},$
    arising from the Poincaré map of the geometric Lorenz flow ($\lambda_{2}<\lambda_{3}<0<-\lambda_{3}<\lambda_{1}$). The derivative of $f$ blows up near 0, and inverse branches have zero derivative there. Diaz-Ordaz [19, Thm. 2] constructs an induced expanding Markov map $T_{A}:I\to I$ (for some neighborhood $I\subset[-1,1]$ of 0) with bounded distortion and first return time $R:I\to\mathbb{N}$ having exponential tails, under uniform contraction (1) and bounded distortion (2).
Induced potentials and summable variations. Given a base potential $\phi_{0}$ for the original system (e.g. $\phi_{0}=-\delta\log|T^{\prime}_{\alpha}|$ for Manneville–Pommeau or $\phi_{0}=-\delta\log|f^{\prime}|$ for Lorenz), define the induced potential on $I$ by $\phi(x):=S_{R(x)}\phi_{0}(x)$. For our results we require summable variations for $\phi$:
$\sum_{n\geq 1}V_{n}(\phi)<\infty,\qquad V_{n}(\phi):=\sup_{\mathbf{a}\in\Sigma^{n}}\ \sup_{x,y\in I_{\mathbf{a}}}|\phi(x)-\phi(y)|,$
which ensures that the Sarig theory applies to the countable Markov map $T_{A}$ [43, 42]. A sufficient condition is $\sum_{n}n\,V_{n}(\phi_{0})<\infty$ [15, §2.4].
Pushing down Gibbs measures. Let $\mu_{\phi}$ be a Gibbs measure for the induced map $T_{A}$. Define the projected measure on the original system by
$\nu_{\phi}(E):=\sum_{a=0}^{\infty}\ \sum_{k=0}^{a}\mu_{\phi}\big(T_{\alpha}^{-k}(E)\cap A_{a}\big),\qquad E\subset I,$
and normalize by $\int R\,d\mu_{\phi}$ when finite. This is the standard way (e.g. for SRB/conformal measures) to transfer statistical properties from the Young tower to the base system; see [15, 18].
Theorem 1.6.
Assume $\phi:I\to\mathbb{R}$ has summable variations and, for some $\gamma_{0}>0$,
$\sum_{a\in\mathcal{A}}\ \sup_{x\in I}w_{a}(x)\,|g_{a}^{\prime}(x)|^{-\gamma_{0}}<\infty,$
where $w_{a}(x)=e^{\phi(g_{a}(x))}$. Let $\mu_{\phi}$ be the Gibbs measure for $T_{A}$ arising from the Manneville–Pommeau map $T_{\alpha}$. Then the projected measure $\nu_{\phi}$ for $T_{\alpha}$ satisfies, for some $\beta>0$,
$\Big|\int\chi(x)\,e^{i\xi\psi(x)}\,d\nu_{\phi}(x)\Big|\ \lesssim\ |\xi|^{-\beta},$
for any $\chi,\psi:I\to\mathbb{R}$ with $\operatorname{supp}\chi$ disjoint from the parabolic fixed point, $\|\chi\|_{C^{\alpha}}<\infty$, $\|\psi\|_{C^{1+\alpha}}<\infty$, and $\inf|\psi^{\prime}|>0$.
For Manneville–Pommeau maps $T_{\alpha}$ with $0<\delta<1$ and $0<\alpha<1$, the choice $\phi_{0}=-\delta\log T^{\prime}_{\alpha}$ yields an induced potential $\phi=S_{R}\phi_{0}$ with summable variations and a light-tail condition for any $\gamma_{0}>0$ (see [15, 18]); hence Theorem 1.6 gives that the $\delta$–conformal measure of $T_{\alpha}$ has positive Fourier dimension. The same argument applies to Lorenz maps once (UNI) is verified for the induced model (see Section 8 for examples of parameters, where this holds).
Remark.
The assumption that $\operatorname{supp}\chi$ avoids the parabolic fixed point can be relaxed to a local decay condition; we keep the support condition for clarity of statement.
Heavy-tailed measures. Finally, an intriguing direction is to relax the light-tail assumption. Because Theorem 1.1 is formulated in terms of exponential moment conditions, it is well-suited to being adapted to heavier tails, where one might expect sub-polynomial (e.g. polylogarithmic) Fourier decay. In the context of Furstenberg measures, such log–log decay has already been demonstrated in [20]. Extending our framework to this regime would provide a natural bridge between strong polynomial decay and much weaker forms of Fourier decay, potentially revealing new structures in dynamical systems with weaker hyperbolicity.
Organisation of the article
The paper is dedicated to the proof of Theorem 1.1, Theorem 1.2, Theorem 1.3, Theorem 1.4 and Theorem 1.6. In Section 2 we first give the preliminary thermodynamic notations and derive the needed moment bounds and upper regularity using the light tail condition where the uniform contraction, bounded distortions and light tail conditions are crucially used. In Section 3 we show how the moments arise from the discretised sum-product bounds if we allow the multiplicative random variables have unbounded support. In Section 4 we then begin the proof of Theorem 1.1 by reducing the Fourier transform of the equilibrium state $\mu$ to exponential sums of multiplicative random variables with moment bounds. To then apply the moment bounds and upper regularity derived in Section 2 and the sum-product bound from Section 3, in Section 5 we prove the needed non-concentration bound from (QNL). In section 6, we prove that (QNL) holds under (MNL) and under $(c_{0},\alpha)$-(UNI), which is Theorem 1.2 and Theorem 1.3. In section 7 we construct a $C^{1+}$ IFS satisfying (MNL) and prove Theorem 1.4. Finally, in section 8 we prove Theorem 1.6.
2. Moment bounds and upper regularity
In this section we introduce some supplementary notations and prove two preliminary statements that will prove useful to prove Theorem 1.1. The first is a precise bound coming from the properties of the pressure. The second is an upper-regularity estimate for $\mu_{\phi}$ on this setting.
2.1. Symbolic coding, Gibbs measures and transfer operators
Recall the notation from Section 1.2, where we are given open disjoint intervals $I_{a}\subset[0,1]$, $a\in\mathcal{A}$ and a $C^{1}$ map $T:\overline{I}\to\mathbb{R}$ for $I=\bigcup_{a\in\mathcal{A}}I_{a}$ satisfying the Markov property: for all $a,b$ such that we have that $T(\overline{I_{b}})\cap(I_{a})\neq\emptyset$, then $T(\overline{I_{b}})\supset\overline{I_{a}}$, which gives us a $\{0,1\}$ transition matrix $A$ satisfying $A(a,b)=1$ if and only if $T(\overline{I_{a}})\supset\overline{I_{b}}$. This gives us a coding $\Sigma^{\mathbb{N}}:=\{\mathbf{a}\in\mathcal{A}^{\mathbb{N}}:A(a_{j},a_{j+1})=1\}$, where length $n$ words are denoted by $\Sigma^{n}$ and finite words are denoted $\Sigma^{*}=\bigcup_{n\in\mathbb{N}}\Sigma^{n}$. Then a word $\mathbf{a}:=a_{1}\dots a_{n}\in\mathcal{A}^{n}$ satisfying $A(a_{i},a_{i+1})=1$ for all $i$ is called admissible and for each admissible $ab\in\mathcal{A}^{2}$, we have an associated inverse branch $g_{ab}:I_{b}\to I_{a}$, which form an iterated function system $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$. Recall that the notation $\mathbf{a}\to x$ simply means that $g_{\mathbf{a}}$ is defined at $x$, i.e. $x\in I_{b\left(\mathbf{a}\right)}$ with $b\left(\mathbf{a}\right)$ denoting the last letter of $\mathbf{a}$.
We then assume the IFS $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$ satisfies the conditions (1)-(3) from Section 1.2, though in this section we only need uniform $\kappa_{+}$ contraction (1) and bounded $(C,\alpha)$-distortions (2). The uniform contraction of the maps $g_{ab}$ guarantees that there exists a unique attractor $F\subset I$ associated to $\Phi$, that is, for $\mathbf{a}\in\Sigma^{n+1}=a_{0}\dots a_{n}$ the map $g_{\mathbf{a}}:=g_{a_{0}a_{1}}\circ\dots\circ g_{a_{n-1}a_{n}}:I_{a_{n}}\rightarrow I_{a_{0}}$, and the interval $I_{\mathbf{a}}:=g_{\mathbf{a}}(I_{a_{n}})$, then $F$ is the compact set given by $F=\bigcap_{n\in\mathbb{N}}\bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}}$.
Let $\mu=\mu_{\phi}$ be the Gibbs measure given by some normalized potential $\phi$, which is uniquely defined to satisfy the $\phi$-conformality condition:
$\int_{F}f(x)\,d\mu_{\phi}(x)=\sum_{ab\in\Sigma^{2}}\int_{I_{b}}e^{\phi(g_{ab}(x))}f(g_{ab}(x))\,d\mu_{\phi}(x)$
for all continuous $f:F\to\mathbb{C}$, where $b(\mathbf{a})$ is the last letter of a word, so that we can write $g_{\mathbf{a}}:I_{b(\mathbf{a})}\rightarrow I_{\mathbf{a}}$. We also assume $\mu_{\phi}$ satisfies the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail condition (4). Now the $\phi$-conformality condition can be iterated to yields the formula:
$\forall f\in C^{0}(F,\mathbb{C}),\ \int_{F}fd\mu=\sum_{\mathbf{a}\in\Sigma^{n+1}}\int_{b(\mathbf{a})}e^{S_{n}\phi(g_{\mathbf{a}}(x))}f(g_{\mathbf{a}}(x))d\mu(x),$
where $S_{n}\phi(x):=\sum_{k=0}^{n-1}\phi(T^{k}(x))$ is a Birkhoff sum. It is then convenient to denote $w_{\mathbf{a}}(x):=e^{S_{n}\phi(g_{\mathbf{a}}(x))}$. The invariance formula can then be rewritten as
$\int fd\mu=\int\mathcal{L}_{\varphi}^{n}(f)d\mu,$
where $\mathcal{L}_{\phi}:f\mapsto\mathcal{L}_{\phi}f$ is the transfer operator associated to $\phi$, given by the formula
$\mathcal{L}_{\phi}(f)(x):=\underset{I_{b}\ni x}{\sum_{ab\in\Sigma^{2}}}w_{ab}(x)f(g_{ab}(x)).$
Iterating yields
$\forall x\in I_{b},\ \mathcal{L}_{\phi}^{n}(f)(x):=\underset{b(\mathbf{a})=b}{\sum_{\mathbf{a}\in\Sigma^{n+1}}}w_{\mathbf{a}}(x)f(g_{\mathbf{a}}(x)).$
Notice that the normalisation assumption on $\phi$ implies that $\mathcal{L}_{\phi}^{n}(1)=1$. Notice further that the $\phi$-conformality assumption (and summable variations of $\phi$) implies the following Gibbs estimate: there exists a Gibbs constant $C_{\phi}\geq 1$ such that
$\forall\mathbf{a}\in\Sigma^{n},\forall x\in I_{b(\mathbf{a})},\ C_{\phi}^{-1}w_{\mathbf{a}}(x)\leq p_{\mathbf{a}}:=\mu(I_{\mathbf{a}})\leq C_{\phi}w_{\mathbf{a}}(x).$
Troughout the text well will use the notation $a_{n}\lesssim b_{n}$ if there exists some constant $C\geq 1$ such that $a_{n}\leq Cb_{n}$. We will write $a_{n}\simeq b_{n}$ if $a_{n}\lesssim b_{n}\lesssim a_{n}$. Gibbs estimates are then rewritten $w_{\mathbf{a}}(x)\simeq\mu(I_{\mathbf{a}})=p_{\mathbf{a}}$ for all $x$.
2.2. Moment bounds from the tail condition
By Birkhoff ergodic theorem, we know that the natural scale for the contractions $|g_{\mathbf{a}}^{\prime}(x)|$ is given by $e^{-\lambda n}$ for $\mu$-almost every $x$, where $\lambda$ is the Lyapunov exponent of $\mu$, is defined by
$\lambda:=-\int\tau(x)\,d\mu(x),$
where $\tau:I\to\mathbb{R}$ is the geometric potential $\tau(x)=-\ln|T^{\prime}(x)|$, $x\in I$. Proposition 2.1 quantifies deviation from this scale of contraction using properties of the Gurevich pressure $P(\phi+t(\tau-\lambda))$, which is defined for a potential $\psi$ as
$P(\psi):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{\mathbf{a}\in\Sigma^{n}}\sup_{x\in I_{b(\mathbf{a})}}e^{S_{n}\psi(g_{\mathbf{a}}(x))},$
where $S_{n}\psi(x)=\sum_{j=0}^{n-1}\psi(T^{k}(x))$. Notice that since $\mathcal{L}_{\phi}^{n}(1)=1$, we find that our normalized potential $\phi$ satify $P(\phi)=0$. It is also known that $\frac{d}{dt}P(\phi+t\tau)_{|t=0}=\int\tau d\mu$. The differentiability properties of the pressure allows us to bound the spreading of the random variable $e^{\lambda n}|g_{\mathbf{a}}^{\prime}(x)|$ in the following way:
Proposition 2.1.
Suppose $\mu$ satisfies the $(C_{\phi,\gamma_{0}},\gamma_{0})$-light tail condition. Then there exists $t_{\gamma_{0}}\in(0,1)$, $\varepsilon_{\gamma_{0}}>0$ such that:
$\forall t\in\left(-t_{\gamma_{0}},t_{\gamma_{0}}\right),\exists n_{0}(t),\forall n\geq n_{0}(t),\forall x\in I,\ \sum_{\mathbf{a}\in\Sigma^{n}\atop\mathbf{a}\rightarrow x}w_{\mathbf{a}}(x)(e^{\lambda n}|g_{\mathbf{a}}^{\prime}(x)|)^{t}\leq 2C_{\phi}e^{\varepsilon_{\gamma_{0}}nt^{2}},$
where
$\varepsilon_{\gamma_{0}}:=\sup_{|t|\leq t_{\gamma_{0}}}\frac{|P(\phi+t(\tau-\lambda))|}{t^{2}}<\infty.$
The optimal choice of $\varepsilon_{\gamma_{0}}$ comes from the second order term
$\frac{d^{2}}{dt^{2}}P(\phi+t(\tau-\lambda))|_{t=0}=\lim_{n\to\infty}\frac{1}{n}\int(S_{n}\tau)^{2}\,d\mu$
of the Taylor approximation for $t\mapsto P(\phi+t(\tau-\lambda))$ near 0. The formula given by the limit of Birkhoff sums of $\tau$ is proved e.g. in [37].
The proof of Proposition 2.1 can be deduced from the proof leading to large deviation principle, e.g. [25] and the references there-in. However, as our result is stated quantitatively in terms of moments, we included a direct proof using Sarig’s results from [43] for completeness.
Proof of Proposition 2.1 .
Let $\gamma_{0}\leq t<0$. Then if write $p_{a}:=\mu(I_{a})$ and recalling that $\kappa_{-}(a)\leq|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for all $x\in I_{b}$, where we have supressed the prefix notation of $b$ for admissibility, then by the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail property and using the Gibbs constant $C_{\phi}$ we have
$\sum_{\mathbf{a}\in\Sigma^{n}\atop\mathbf{a}\rightarrow x}w_{\mathbf{a}}(x)|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x)|^{t}\leq C_{\phi}\Big(\sum_{a\in\mathcal{A}}p_{a}(e^{\lambda}\kappa_{-}(a))^{t}\Big)^{n}$
$\leq C_{\phi}e^{t\lambda n}\Big(\sum_{a\in\mathcal{A}}p_{a}\kappa_{-}(a)^{-t}\Big)^{n}\leq C_{\phi}e^{t\lambda n}C_{\phi,\gamma_{0}}^{n}=C_{\phi}e^{(t\lambda+\log C_{\phi,\gamma_{0}})n}$
Moreover, if $t\geq 0$, then we can simply bound
$\sum_{\mathbf{a}\in\Sigma^{n}}w_{\mathbf{a}}(x)|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x)|^{t}\leq C_{\phi}\Big(\sum_{a\in\mathcal{A}}p_{a}(e^{\lambda}\kappa_{+})^{t}\Big)^{n}=C_{\phi}(e^{\lambda}\kappa_{+})^{tn}=C_{\phi}e^{t(\lambda+\log\kappa_{+})n},$
where note that by uniform $\kappa_{+}$-contraction the Lyapunov exponent of $\mu$ satisfies $\lambda\geq-\int\log\kappa_{+}\,d\mu(x)=-\log\kappa_{+}>1$ as $\kappa_{+}<1$, so the exponent on is non-negative. Since
$\frac{1}{n}\ln\sum_{\mathbf{a}\in\Sigma^{n}\atop\mathbf{a}\to x}w_{\mathbf{a}}(x)|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x)|^{t}\underset{n\rightarrow\infty}{\longrightarrow}P(\varphi+t(\tau-\lambda))$
this proves that for all $t\in[-\gamma_{0},\infty)$ the pressure is finite:
[…]
Since $\tau$ is $\alpha$-Hölder by the bounded distortion property and as the pressure $P(\varphi+t(\tau-\lambda))$ is finite for all $|t|\leq\gamma_{0}$, we know by [43, Theorem 4] that there exists $0<t_{\gamma_{0}}^{\prime}<\gamma_{0}$ such that the function $f:t\in(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})\mapsto P(\varphi+t(\tau-\lambda))$ is real analytic in $(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})$ and satisfies that $f(0)=P(\varphi)=0$ and $f^{\prime}(0)=\int_{I}(\tau-\lambda)d\mu=0$. Taylor estimate near 0 then gives some $t_{\gamma_{0}}^{\prime\prime}\in(0,1)$ such that $|f(t)|\leq\tfrac{3}{4}f^{\prime\prime}(0)t^{2}$ for $|t|\leq t_{\gamma_{0}}^{\prime\prime}$. Thus choosing $0<t_{\gamma_{0}}<t_{\gamma_{0}}^{\prime\prime}$ small enough, it follows that, for all $|t|\leq t_{\gamma_{0}}$, there exists $n_{0}(t)\geq 1$ such that
$\forall n\geq n_{0}(t),\ \sum_{\mathbf{a}\in\Sigma^{n}}w_{\mathbf{a}}(x)|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x)|^{t}\leq e^{f^{\prime\prime}(0)nt^{2}}.$
and we get the claim for $\varepsilon_{\gamma_{0}}=f^{\prime\prime}(0)$ as $\varepsilon_{\gamma_{0}}t_{\gamma_{0}}^{2}\geq|f(t_{\gamma_{0}})|$. We get a bound uniform in $x$ because of Gibbs estimates and bounded variations. ∎
2.3. Upper regularity of the Gibbs measure from the tail
The moment bound deduced in Proposition 2.1 allows us to prove the following Frostman condition or upper regularity bound for the measure $\mu$ allowing us to disregard $\mu$ concentrating near points at small scales. We note that if one would relax the light tail assumption to heavier tail assumptions, the upper regularity of $\mu_{\phi}$ is likely to fail.
Proposition 2.2 (Upper regularity).
There exists $C_{reg}\geq 1$, $s_{\mu}\in(0,1)$ such that:
$\forall x\in\mathbb{R},\ \forall\sigma\in(0,1),\ \mu([x-\sigma,x+\sigma])\leq C_{reg}\sigma^{s_{\mu}}.$
Proof.
First of all, $\sum_{\mathbf{a}\in\Sigma^{n}}\mu(I_{\mathbf{a}})=\mu(I)$, so that $(I_{\mathbf{a}})_{\mathbf{a}}$ is a partition of $I$ up to a zero-mesure set. Now consider some interval $J\subset I$. Define $\mathcal{W}_{n}(J):=\{\mathbf{a}\in\Sigma^{n},\ I_{\mathbf{a}}\cap J\neq 0\}$. We have, for all $n\geq 1$:
$J\subset\bigcup_{\mathbf{a}\in\mathcal{W}_{n}(J)}I_{\mathbf{a}}.$
Hence:
$\mu(J)\leq\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}\mu(I_{\mathbf{a}})\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}w_{\mathbf{a}}(x_{b(\mathbf{a})}).$
for any choice of reference points $x_{a}\in I_{a}$. Define further: $E_{n}:=\{\mathbf{a}\in\Sigma^{n},\ |g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\geq e^{-2\lambda n}\}.$ Proposition 2.1 then yields, for $t$ small enough and $n$ large enough (depending on $t$):
$\sum_{\mathbf{a}\in\Sigma^{n}\setminus E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})\leq\sum_{\mathbf{a}\in\Sigma^{n}\setminus E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})e^{-2\lambda nt}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{-t}=e^{-\lambda nt}\sum_{\mathbf{a}\in\Sigma^{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{-t}$
$\leq 2C_{\phi}e^{-\lambda nt}e^{\varepsilon_{\gamma_{0}}nt^{2}}\leq 2C_{\phi}e^{-n\tilde{\varepsilon}}$,
for some $\tilde{\varepsilon}>0$. We can then write:
$\mu(J)\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}w_{\mathbf{a}}(x_{b(\mathbf{a})})\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})+e^{-\tilde{\varepsilon}n}.$
Our second step is to bound $w_{\mathbf{a}}$ by a power of $|g_{\mathbf{a}}^{\prime}|$. First of all, recall that our moment condition is $\sum_{a\in\Sigma}\sup_{x}w_{ab}(x)|g_{ab}^{\prime}(x)|^{-\gamma}<\infty$ for some $\gamma>0$. It follows that there exists a constant $C$ such that:
$\forall ab\in\Sigma^{2},\forall x\in I_{b},\ \ w_{ab}(x)\leq C|g_{ab}^{\prime}(x)|^{\gamma}.$
Now, since $\sum_{ab\in\Sigma}|I_{ab}|<\infty$, we know that $g_{ab}^{\prime}(x)$ goes to zero when the word $ab$ goes far in the set $\Sigma^{2}$. In particular, there exists a finite set $\mathcal{B}\subset\Sigma^{2}$ such that:
$\forall ab\in\Sigma^{2}\setminus\mathcal{B},\ w_{ab}(x)\leq|g_{ab}^{\prime}(x)|^{\gamma/2}.$
Then, for every $ab\in\mathcal{B}$, we can write, since $\varphi\leq-\varepsilon_{\varphi}<0$:
$w_{ab}(x)=\exp(\varphi(g_{ab}(x)))\leq\exp(-\varepsilon_{\varphi})\leq\exp\Big(-\frac{\varepsilon_{\varphi}}{\max_{cd\in\mathcal{B}}|\ln g_{cd}^{\prime}|_{\infty}}\big|\ln|g_{ab}^{\prime}(x)|\big|\Big)=|g_{ab}^{\prime}(x)|^{\widehat{\varepsilon}}$,
for some $\widehat{\varepsilon}\in(0,1)$. We then find, reducing $\widehat{\varepsilon}$ if necessary:
$\forall ab\in\Sigma^{2},\forall x\in I_{b},\ w_{ab}(x)\leq|g_{ab}^{\prime}(x)|^{-\widehat{\varepsilon}}$.
We can then bound similar quantities for some word $\mathbf{a}$. We find, for any $\mathbf{a}\in\Sigma^{n}$, and for any $x\in I_{b(\mathbf{a})}$:
$w_{\mathbf{a}}(x)=w_{a_{1}a_{2}}(g_{a_{2}\dots a_{n}}(x))\dots w_{a_{n-1}a_{n}}(x)\leq|g_{a_{1}a_{2}}^{\prime}(g_{a_{2}\dots a_{n}}(x))^{\widehat{\varepsilon}}\dots g_{a_{n-1}a_{n}}^{\prime}(x)^{\widehat{\varepsilon}}|=|g_{\mathbf{a}}^{\prime}(x)|^{\widehat{\varepsilon}}$.
We then inject this bound into the previous sum to find, using Hölder’s inequality:
$\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})\leq\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{\widehat{\varepsilon}}\leq\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\Big)^{\widehat{\varepsilon}}\cdot\#(\mathcal{W}_{n}(J)\cap E_{n})^{1-\widehat{\varepsilon}}$.
Now notice the following. Since we are working with a partition of intervals, there can only exists at most $2$ words $\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}$ such that $I_{\mathbf{a}}\nsubseteq J$. So we can write
$\mathcal{W}_{n}(J)\cap E_{n}=X_{n}(J)\cup Y_{n}(J),$
where $\bigcup_{\mathbf{a}\in X_{n}(J)}I_{\mathbf{a}}\subset J$, and where $Y_{n}(J)\leq 2$. We can then bound the cardinal of $X_{n}(J)$, as follows:
$X_{n}(J)\leq e^{\lambda n/2}\sum_{\mathbf{a}\in X_{n}(J)}|I_{\mathbf{a}}|\leq|J|e^{\lambda n/2}.$
We then relate $n$ with $|J|$ by asking for this quantity to be bounded:
$n:=-\frac{2}{\lambda}\ln(|J|).$
In this case, $\mathcal{W}_{n}(J)\cap E_{n}\lesssim 1$. We finally conclude the proof in the following way, denoting
$\kappa=\sup_{a\in\Sigma}\|g_{a}^{\prime}\|_{\infty}\in(0,1)$:
$\mu(J)\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})+e^{-\tilde{\varepsilon}n}\lesssim\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}$
$\lesssim\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|I_{\mathbf{a}}|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}=\Big(2\kappa^{n}+\sum_{\mathbf{a}\in X_{n}(J)}|I_{\mathbf{a}}|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}$
$\lesssim\kappa^{\widehat{\varepsilon}n}+|J|^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}\lesssim|J|^{s_{\mu}},$
for some $s_{\mu}>0$ small enough. ∎
3. Discretised sum-product bounds under light tails
Our strategy is to reduce the Fourier decay bound to exponential sums of multiplicative random variables coming from products of the derivatives $g_{a}^{\prime}(x)$ of the IFS. To bound this, as in Bourgain and Dyatlov in [13], one can reduce this to a multiscale non-concentration bound via the discretised sum-product phenomenon [12] coming from additive combinatorics. However, to gain the multiscale non-concentration, previous works require smoothness from the underlying dynamics such as in the form of application of Dolgopyat’s method [41, 10, 9, 27, 30]. The main novelty of our paper lies in the analysis of the non-concentration bound so it can be studied also in the low regularity that cover $C^{1+\alpha}$ systems.
Next we give the necessary consequences of the sum-product theorem that we need in our setting that are similar to Bourgain and Dyatlov [13]. Since we also work with parabolic systems, the derivatives $g_{a}^{\prime}$ in the infinitely branched systems may have arbitrarily small values, we need to formulate the bounds using moments of the multiplicative random variables, which we can link to the light tail condition by the previous Section 2.
Firstly, let us recall a quantitative version of this sum-product phenomenon, where this version is a consequence of [36, Theorem 1.14], see also [13]:
Theorem 3.1.
Let $\gamma_{4}>0$. Then there exists $\varepsilon_{2}\in(0,1)$ such that for any $k\geq C\gamma_{4}^{-1}$ the following hold (where $C$ denotes an absolute constant). Let $\eta\in\mathbb{R}$, $|\eta|\geq 1$, let $C_{0}\geq 1$, and let $\mu_{1},\dots,\mu_{k}$ be Borel measures supported on $C_{0}^{-1},C_{0}]$ such that $\mu(\mathbb{R})\leq C_{0}$ so that
$\forall a\in\mathbb{R},\forall\sigma\in[|\eta|^{-1},|\eta|^{-\varepsilon_{2}}],\ \mu_{j}([a-\sigma,a+\sigma])\leq C_{0}\sigma^{\gamma_{4}}.$
Then, for some constant $C_{1}$ that only depends on $C_{0},\gamma$:
$\Big|\int_{[C_{0}^{-1},C_{0}]^{k}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1}(x_{1})\dots d\mu_{k}(x_{k})\Big|\leq C_{1}|\eta|^{-\exp(-C\gamma_{4}^{-1})}.$
We will generalise this result for the case where $\mu$ is supported on $\mathbb{R}^{*}_{+}$, under a polynomial moment condition.
Theorem 3.2.
Let $\gamma_{3}>0$. Then there exists $\varepsilon_{1}\in(0,1)$ and $k\geq 1$ such that the following holds: Let $\eta\in\mathbb{R}$, $|\eta|\geq 1$ and let $\mu_{1},\dots,\mu_{k}$ be Borel probability measures on $\mathbb{R}^{*}_{+}$ such that:
$\forall j\in\{1,\dots,k\},\ \forall a\in\mathbb{R},\ \forall\sigma\in[|\eta|^{-2},|\eta|^{-\varepsilon_{1}}],\ (\ln_{*}\mu_{j})([a-\sigma,a+\sigma])\leq 2\sigma^{\gamma_{3}}.$
Then, for some constant $c$ that only depends on $\gamma_{3}$, we find:
$\forall t\in(0,\varepsilon_{1}),\ \Big|\int_{(\mathbb{R}_{+}^{*})^{k}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1}(x_{1})\dots d\mu_{k}(x_{k})\Big|\leq c|\eta|^{-\varepsilon_{1}}+|\eta|^{-\frac{t^{3}}{2\sqrt{k}}}\prod_{j=1}^{k}M_{j}(t^{2}),$
where $M_{j}(t):=\int_{\mathbb{R}^{*}_{+}}e^{t|\ln x|}d\mu_{j}(x)$.
Proof.
Let $\gamma_{3}\in(0,1)$. Define $\gamma_{4}:=\gamma_{3}/2$. Then Theorem 3.1 gives us some $k\geq 1$ and some $\varepsilon_{2}\in(0,1)$ (associated to $\gamma_{4})$. Taking $\varepsilon_{2}$ smaller if necessary, let us assume $\varepsilon_{2}\leq e^{-C/\gamma_{4}}$. We will define $\varepsilon_{1}:=\big(\frac{\varepsilon_{2}}{4k}\big)^{6}$. Now, let $\eta\in\mathbb{R}$ be such that $|\eta|\geq 1$. Let $\mu_{1},\dots,\mu_{k}$ be measures that satisfies the hypothesis of Theorem 3.2 with respect to the quantities $(\gamma_{3},k,\varepsilon_{1},\eta)$. We wish to show that the bound on the integral holds.
The idea is to do a dyadic decomposition of the measures $\mu_{j}$. For each $n\in\mathbb{Z}$, we define $I_{n}:=[2^{n},2^{n+1}[$, and we define $\mu_{j,n}:=(\mu_{j})\mid_{I_{n}}$. We find
$\int_{(\mathbb{R}_{+}^{*})^{k}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1}(x_{1})\dots d\mu_{k}(x_{k})=\sum_{\vec{n}\in\mathbb{Z}^{k}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k}),$
where $\vec{n}=(n_{1},\dots,n_{k})$ and $I_{\vec{n}}=I_{n_{1}}\times\dots\times I_{n_{k}}$. We then cut the sum in two depending on if $|\vec{n}|_{\infty}:=\max_{j}|n_{j}|$ is large or small. Fix some $t\in(0,\varepsilon_{1})$. We have:
$\int_{(\mathbb{R}_{+}^{*})^{k}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1}(x_{1})\dots d\mu_{k}(x_{k})=$
$\underset{|\vec{n}|_{\infty}<t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k})\ +\underset{|\vec{n}|_{\infty}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k}).$
The sum for large $|\vec{n}|_{\infty}$ gives
$\Big|\underset{|\vec{n}|_{\infty}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k})\Big|\leq\underset{|\vec{n}|_{\infty}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\mu_{1}(I_{n_{1}})\dots\mu_{k}(I_{n_{k}}).$
The moment condition on the measures $\mu_{j}$ yields
$\forall j,\ \forall n\in\mathbb{Z},\ \mu_{j}(I_{n})\leq\int_{[2^{n},2^{n+1}[}2^{-t^{2}|n|}e^{t^{2}|\ln x|}d\mu_{j}(x)\leq 2^{-t^{2}|n|}M_{j}(t^{2}),$
which gives, injecting into the previous sum (denoting by $d\sigma$ the Lebesgue measure on the sphere, $M(t):=M_{1}(t)\dots M_{k}(t)$ and $|\vec{n}|_{p}:=(\sum_{j}|n_{j}|^{p})^{1/p}$):
$\Big|\underset{|\vec{n}|_{\infty}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k})\Big|\leq M(t^{2})\underset{|\vec{n}|_{\infty}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}2^{-t^{2}|\vec{n}|_{1}}$
$\leq M(t^{2})\underset{|\vec{n}|_{2}\geq t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}2^{-\frac{t^{2}|\vec{n}|_{2}}{\sqrt{k}}}\lesssim M(t^{2})\int_{\{x\in\mathbb{R}^{k},\ \|x\|_{2}\geq t\log_{2}|\eta|\}}2^{-\frac{t^{2}\|x\|_{2}}{\sqrt{k}}}dx$
$=M(t^{2})\int_{\mathbb{S}^{k-1}}\int_{t\log_{2}|\eta|}^{\infty}2^{-\frac{t^{2}r}{\sqrt{k}}}r^{k-1}drd\sigma\lesssim M(t^{2})\int_{t\log_{2}|\eta|}^{\infty}2^{-\frac{t^{2}r}{2\sqrt{k}}}dr\lesssim M(t^{2})\eta^{-\frac{t^{3}}{2\sqrt{k}}}.$
To conclude, it now suffices to bound the sum where $|\vec{n}|_{\infty}$ is small. We can rewrite it as
$\underset{|\vec{n}|_{\infty}<t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k})=\underset{|\vec{n}|_{\infty}<t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{[1,2]^{k}}e^{i\eta_{\vec{n}}x_{1}\dots x_{k}}d\nu_{1,n_{1}}(x_{1})\dots d\nu_{k,n_{k}}(x_{k}),$
where $\nu_{j,n}$ are the measures on $[1,2]$ defined by the relation $(m_{n})_{*}\nu_{j,n}=\mu_{j,n}$, for $m_{n}(x):=2^{n}x$, and where we denoted $\eta_{\vec{n}}:=\eta 2^{n_{1}+\dots n_{k}}$. Now fix some $\vec{n}\in\mathbb{Z}^{k}$ with $|\vec{n}|_{\infty}\leq t\log_{2}|\eta|$. We will check that the measures $\nu_{1,n_{1}},\dots,\nu_{k,n_{k}}$ satisfies the hypothesis of Theorem 3.1 for the quantities $(\gamma_{4},\varepsilon_{2},k,\eta_{\vec{n}})$. Notice that $|\eta|^{1/2}\leq|\eta_{\vec{n}}|\leq|\eta|^{2}$ since $t<\varepsilon_{1}<\frac{1}{(4k)^{4}}$. We check the following conditions:
  • Total mass condition:
    $\forall j,\ \nu_{j,n_{j}}(\mathbb{R})=\mu_{j}(I_{n_{j}})\leq 1.$
  • Support condition:
    $\forall j,\ \text{supp}(\nu_{j,n_{j}})\subset[1,2].$
  • Non-concentration condition:
    $\forall j,\ \forall a\in[1,2],\ \forall\sigma\in[|\eta_{\vec{n}}|^{-1},|\eta_{\vec{n}}|^{-\varepsilon_{2}}]\subset[\eta^{-2},\eta^{-\varepsilon_{2}/2}]:$
    $\nu_{j,n_{j}}([a-\sigma,a+\sigma])\leq\nu_{j,n_{j}}(\{x\in[1,2],\ \ln x\in[\ln a-2\sigma,\ln a+2\sigma]\})$
    $\mu_{j}(\{x\in I_{n_{j}},\ \ln x\in[n_{j}\ln(2)+\ln(a)-2\sigma,n_{j}\ln(2)+\ln(a)+2\sigma]\})\leq 4\sigma^{\gamma_{3}}$
    by the non-concentration hypothesis on $\ln_{*}\mu_{j}$.
Hence the conclusion of Theorem 3.1 apply in this setting. That is, for some constant $c>0$ that only depends on $\gamma_{3}$, we find the following bound:
$\Big|\int_{[1,2]^{k}}e^{i\eta_{\vec{n}}x_{1}\dots x_{k}}d\nu_{1,n_{1}}(x_{1})\dots d\nu_{k,n_{k}}\Big|\leq c|\eta_{\vec{n}}|^{-\varepsilon_{2}}\leq c|\eta|^{-\varepsilon_{2}/2}.$
We then inject this estimate in the sum, and find:
$\Big|\underset{|\vec{n}|_{\infty}<t\log_{2}|\eta|}{\sum_{\vec{n}\in\mathbb{Z}^{k}}}\int_{I_{\vec{n}}}e^{i\eta x_{1}\dots x_{k}}d\mu_{1,n_{1}}(x_{1})\dots d\mu_{k,n_{k}}(x_{k})\Big|\leq c(t\log_{2}|\eta|)^{k}|\eta|^{-\varepsilon_{2}/2}\lesssim_{\gamma_{3}}|\eta|^{-\varepsilon_{2}/4}\lesssim_{\gamma_{3}}\eta^{-\varepsilon_{1}},$
thus concluding the proof. ∎
We will apply the previous Theorem to the special case where the measures are countable sums of Dirac mass, which correspond to our present setting.
Theorem 3.3.
Let $\gamma_{2}\in(0,1)$. Then there exists $\varepsilon_{1}\in(0,1)$ and $k\geq 1$ such that the following holds. Let $\eta\in\mathbb{R}$, $|\eta|\geq 1$. Let $\mathcal{Z}_{1},\dots,\mathcal{Z}_{k}$ be a finite or countable sets. Choose, for each $j$, some nonnegative weights $(w_{\mathbf{b}}^{(j)})_{\mathbf{b}\in\mathcal{Z}_{j}}$ that sums to one, and some maps $\zeta_{j}:\mathcal{Z}_{j}\rightarrow\mathbb{R}_{+}^{*}$. Suppose that for all $j\in\{1,\dots,k\}$, the following non-concentration condition holds:
$\forall\sigma\in[\eta^{-2},\eta^{-\varepsilon_{1}}],\ \sum_{(\mathbf{b},\mathbf{c})\in\mathcal{Z}_{j}^{2}}w_{\mathbf{b}}^{(j)}w_{\mathbf{c}}^{(j)}\mathbf{1}_{[-\sigma,\sigma]}(\ln\zeta_{j}(\mathbf{b})-\ln\zeta_{j}(\mathbf{c}))\leq\sigma^{\gamma_{2}}.$
Then, there exists a constant $c>0$ that only depends on $\gamma$ such that
$\forall t>0,\ \Big|{{\sum_{\mathbf{b}_{1}\in\mathcal{Z}_{1}}}}\dots{\sum_{\mathbf{b}_{k}\in\mathcal{Z}_{k}}}w_{\mathbf{b}_{1}}^{(1)}\dots w_{\mathbf{b}_{k}}^{(k)}e^{i\eta\zeta_{1}(\mathbf{b}_{1})\dots\zeta_{k}(\mathbf{b}_{k})}\Big|\leq c|\eta|^{-\varepsilon_{1}^{2}}+\eta^{-\frac{t^{3}}{2\sqrt{k}}}\prod_{j=1}^{k}M_{j}(t^{2}),$
where $M_{j}(t):=\sum_{\mathbf{b}\in\mathcal{Z}_{j}}w_{\mathbf{b}}^{(j)}e^{t|\ln\zeta_{j}(\mathbf{b})|}.$
Proof.
Let $\gamma_{2}>0$. We then fix $\gamma_{3}:=\gamma_{2}/2$ and let $\varepsilon_{1}$ and $k\geq 1$ be given by Theorem 3.2 associated to $\gamma_{3}$. Let $\mathcal{Z}_{j}$, $(w_{\mathbf{b}}^{(j)})_{\mathbf{b}\in\mathcal{Z}_{j}}$ and $\zeta_{j}:\mathcal{Z}_{j}\rightarrow\mathbb{R}_{+}^{*}$ be such that they satisfy the hypothesis of Theorem 3.3. Define some measures $\mu_{j}$ by
$\mu_{j}:=\sum_{\mathbf{b}\in\mathcal{Z}_{j}}w_{\mathbf{b}}^{(j)}\delta_{\zeta_{j}(\mathbf{b})}.$
Let us check the non-concentration condition:
$\ln_{*}\mu_{j}([a-\sigma,a+\sigma])=\sum_{\mathbf{b}\in\mathcal{Z}_{j}}w_{\mathbf{b}}^{(j)}\mathbf{1}_{[a-\sigma,a+\sigma]}(\ln\zeta_{j}(\mathbf{b}))$
$=\Big(\sum_{\mathbf{b},\mathbf{c}\in\mathcal{Z}_{j}}w_{\mathbf{b}}^{(j)}w_{\mathbf{c}}^{(j)}\mathbf{1}_{[a-\sigma,a+\sigma]}(\ln\zeta_{j}(\mathbf{b}))\mathbf{1}_{[a-\sigma,a+\sigma]}(\ln\zeta_{j}(\mathbf{c}))\Big)^{1/2}$
$\leq\Big(\sum_{\mathbf{b},\mathbf{c}\in\mathcal{Z}_{j}}w_{\mathbf{b}}^{(j)}w_{\mathbf{c}}^{(j)}\mathbf{1}_{[2\sigma,2\sigma]}(\ln\zeta_{j}(\mathbf{b})-\ln\zeta_{j}(\mathbf{c}))\Big)^{1/2}\leq(2\sigma)^{\gamma_{2}/2}.$
The conclusion follows by applying the previous Theorem. ∎
4. Discretisation to exponential sums of random products
We finally start to prove Theorem 1.1. The proof is separated in two parts. The first part is to reduce the Fourier decay bound to the sum-product phenomenon (Theorem 3.3) for some well chosen maps $(\zeta_{j})$. This is the content of Section 4. The second part is to check that the non-concentration condition is actually satisfied for these maps $\zeta_{j}$ under (QNL). This is the content of Section 5.
4.1. Iterating into exponential sums of phase differences
Recall from Section 2 that the transfer operator associated to the potential $\phi$ is defined by
$\forall x\in I_{b},\ \mathcal{L}_{\phi}f(x)=\sum_{a\in\mathcal{A},ab\in\Sigma^{2}}e^{\phi(g_{ab}(x))}f(g_{ab}(x)),f:I\to\mathbb{C}.$
Let $\mu=\mu_{\phi}$ be the Gibbs measure associated to $\phi$, supported on the attractor $F$ of our IFS. If $\xi\in\mathbb{R}$ with $|\xi|$ large enough, choose the unique $n\in\mathbb{N}$ such that
$e^{((2k+1)\lambda+\varepsilon_{0})n}\leq|\xi|<e^{((2k+1)\lambda+\varepsilon_{0})(n+1)},$
where $\varepsilon_{0}>0$ is a small constant that we will chose later (see Definition 4.1), the point being that $\xi e^{-(2k+1)\lambda n}\simeq e^{\varepsilon_{0}n}$ is a term that we will chose to be slowly growing.
We will iterate $(2k+1)$ times the invariance relation $\int f\,d\mu=\int\mathcal{L}_{\phi}^{n}f\,d\mu$. Because of the relation $g_{ab}\circ g_{bc}=g_{abc}$, it is convenient, for a given word $a_{1}\dots a_{n+1}$, to define the notation $\mathbf{a}^{\prime}:=a_{1}\dots a_{n}$. We can then write $g_{\mathbf{a^{\prime}b}}=g_{\mathbf{a}}\circ g_{\mathbf{b}}$. Iterating the relation then gives
$\Big|\int\chi(x)e^{i\xi\psi(x)}\,d\mu(x)\Big|^{2}=\Big|\sum\limits_{\mathbf{A},\mathbf{B}}\int_{I_{b(\mathbf{A})}}w_{\mathbf{A}*\mathbf{B}}(x)\chi(g_{\mathbf{A}*\mathbf{B}}(x))e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2},$
where the sum is over blocks $\mathbf{A}=\mathbf{a}_{1}\dots\mathbf{a}_{k+1}\in\mathcal{A}^{n(k+1)},\mathbf{B}=\mathbf{b}_{1}\dots\mathbf{b}_{k}\in\mathcal{A}^{nk}$ of words $\mathbf{a}_{j},\mathbf{b}_{j}\in\Sigma^{n}$ such that the concatenation
$\mathbf{A}*\mathbf{B}:=\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}^{\prime}\mathbf{a}_{2}^{\prime}\mathbf{b}_{2}^{\prime}\dots\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}^{\prime}\mathbf{a}_{k+1}.$
is admissible (i.e. element of $\Sigma^{(2k+1)n}$). Recall that we use the weight notation
$w_{\mathbf{A}*\mathbf{B}}(x):=e^{S_{(2k+1)n}\phi(g_{\mathbf{A}*\mathbf{B}}(x)))}.$
We fix from now on some choices of reference points $x_{a}\in I_{a}$. We can then multiply and divide to get a form:
$\Big|\int\chi(x)e^{i\xi\psi(x)}\,d\mu(x)\Big|^{2}=\Big|\sum\limits_{\mathbf{A},\mathbf{B}}w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}\chi(g_{\mathbf{A}*\mathbf{B}}(x))e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2},$
which by Cauchy-Schwartz inequality and Gibbs property of $\mu$ is bounded by a constant multiple of
$\sum\limits_{\mathbf{A},\mathbf{B}}w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}\chi(g_{\mathbf{A}*\mathbf{B}}(x))e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}.$
Since $\chi$ is Hölder with $\|\chi\|_{\beta_{\chi}}\lesssim|\xi|^{\varepsilon_{\chi}}$, we have
$|\chi(g_{\mathbf{A}*\mathbf{B}}(x))-\chi(g_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})}))|\leq C|\xi|^{\varepsilon_{\chi}}|g_{\mathbf{A}*\mathbf{B}}(x)-g_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})|^{\beta_{\chi}}\leq C|\xi|^{\varepsilon_{\chi}}\kappa_{+}^{\beta_{\chi}(2k+1)n},$
so by the Gibbs property $\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}\lesssim 1$ we obtain a bound
$\sum\limits_{\mathbf{A},\mathbf{B}}w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})\chi(g_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})}))^{2}\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{0})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}+C^{\prime}|\xi|^{\varepsilon_{\chi}}\kappa_{+}^{2\beta_{\chi}(2k+1)n}.$ (4.1)
Since $\mathbf{A}*\mathbf{B}=\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}^{\prime}\mathbf{a}_{2}^{\prime}\mathbf{b}_{2}^{\prime}\dots\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}^{\prime}\mathbf{a}_{k+1}$, we know that for some $t_{j}\in F$:
$w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})=w_{\mathbf{a}_{1}}(t_{1})w_{\mathbf{b}_{1}}(t_{2})w_{\mathbf{a}_{2}}(t_{3})w_{\mathbf{b}_{2}}(t_{4})\dots w_{\mathbf{a}_{k}}(t_{2k-1})w_{\mathbf{b}_{k}}(t_{2k})w_{\mathbf{a}_{k+1}}(t_{(2k+1)}).$
Thus we can apply Gibbs property applied $2k+1$ times to obtain
$w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})\leq C^{2k+1}\mu(I_{\mathbf{a}_{1}})\mu(I_{\mathbf{b}_{1}})\mu(I_{\mathbf{a}_{2}})\mu(I_{\mathbf{b}_{2}})\dots\mu(I_{\mathbf{a}_{k}})\mu(I_{\mathbf{b}_{k}})\mu(I_{\mathbf{a}_{k+1}}).$
Writing $p_{\mathbf{A}}=\mu(I_{\mathbf{a}_{1}})\dots\mu(I_{\mathbf{a}_{k+1}})$ and $p_{\mathbf{B}}=\mu(I_{\mathbf{b}_{1}})\dots\mu(I_{\mathbf{b}_{k}})$, we can bound (4.1) by
$C^{2k+1}\|\chi\|_{\infty}^{2}\sum\limits_{\mathbf{A}}p_{\mathbf{A}}\sum_{\mathbf{B}:\mathbf{A}\ast\mathbf{B}\in\Sigma^{(2k+1)n}}p_{\mathbf{B}}\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}+C^{\prime}|\xi|^{\varepsilon_{\chi}}\kappa_{+}^{2\beta_{\chi}(2k+1)n}.$
Here as $|\xi|\lesssim e^{(2k+1)\lambda n+\varepsilon_{0}n}$ and $e^{-\lambda}\leq\kappa_{+}$, we have $|\xi|^{\varepsilon_{\chi}}\kappa_{+}^{2\beta_{\chi}(2k+1)n}\lesssim|\xi|^{-\beta^{\prime}}$ for some $\beta^{\prime}>0$ as long as $\varepsilon_{\chi}>0$ is small enough. Thus, fixing $\mathbf{A}$, we just need to have decay for
$\sum_{\mathbf{B}}p_{\mathbf{B}}\Big|\int_{b(\mathbf{A})}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}.$
Writing $\mathbf{A}\sharp\mathbf{B}=\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}^{\prime}\mathbf{a}_{2}^{\prime}\mathbf{b}_{2}^{\prime}\dots\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}$, we have $\mathbf{A}*\mathbf{B}=(\mathbf{A}\sharp\mathbf{B})^{\prime}\mathbf{a}_{k+1}$. We have that at any $x\in I_{b(\mathbf{A})}$:
$g_{\mathbf{A}*\mathbf{B}}^{\prime}(x)=g_{\mathbf{A}\sharp\mathbf{B}}^{\prime}(g_{\mathbf{a}_{k+1}}(x))g_{\mathbf{a}_{k+1}}^{\prime}(x),$
and
$w_{\mathbf{A}*\mathbf{B}}(x)=w_{\mathbf{A}\sharp\mathbf{B}}(g_{\mathbf{a}_{k+1}}(x))w_{{\mathbf{a}_{k+1}}}(x).$
By Gibbs property $\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\lesssim 1$, the Hölder property of $\phi$ gives us
$\Big|\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}-\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\Big|=\Big|\frac{w_{\mathbf{A}\sharp\mathbf{B}}(g_{\mathbf{a}_{k+1}}(x))}{w_{\mathbf{A}\sharp\mathbf{B}}(g_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})}))}-1\Big|\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}$
$\lesssim|S_{2kn}\phi(g_{\mathbf{A}\sharp\mathbf{B}}(g_{\mathbf{a}_{k+1}}(x)))-S_{2kn}\phi(g_{\mathbf{A}\sharp\mathbf{B}}(g_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})))|$
$\leq C|g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})|^{\beta}\lesssim\kappa_{+}^{\beta n}.$
Thus we obtain a bound for some universal $C^{\prime}>0$:
$\sum_{\mathbf{B}}p_{\mathbf{B}}\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}+C^{\prime}\kappa_{+}^{\beta n},$
so we just need to study
$\sum_{\mathbf{B}}p_{\mathbf{B}}\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}.$
Expand the square so we just need to bound
$\iint_{I_{b(\mathbf{A})}^{2}}\sum_{\mathbf{B}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\frac{w_{\mathbf{a}_{k+1}}(y)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}p_{\mathbf{B}}e^{i\xi(\psi(g_{\mathbf{A}*\mathbf{B}}(x))-\psi(g_{\mathbf{A}*\mathbf{B}}(y)))}\,d\mu(x)\,d\mu(y)$
$=\iint_{I_{b(\mathbf{A})}^{2}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\frac{w_{\mathbf{a}_{k+1}}(y)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi(\psi(g_{\mathbf{A}*\mathbf{B}}(x))-\psi(g_{\mathbf{A}*\mathbf{B}}(y)))}\,d\mu(x)\,d\mu(y).$
4.2. Linearisation of the phase
Let us now linearise the phase. By the mean value theorem, we have, for some $t=t_{\mathbf{A},\mathbf{B}}(x,y)$:
$\psi(g_{\mathbf{A}*\mathbf{B}}(x))-\psi(g_{\mathbf{A}*\mathbf{B}}(y)))=(\psi\circ g_{\mathbf{A}\sharp\mathbf{B}})^{\prime}(t)(g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)).$
By the chain rule, we can then find $t_{1}\in I_{\mathbf{a}_{1}},t_{2}\in I_{\mathbf{a}_{2}},\dots,t_{k+1}\in I_{\mathbf{a}_{k+1}}$ (depending on $t$) such that
$(\psi\circ g_{\mathbf{A}\sharp\mathbf{B}})^{\prime}(t)=\psi^{\prime}(t_{1})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1}).$
We can then rewrite our sum of interest as
$\sum_{\mathbf{B}}p_{\mathbf{B}}\Big|\int_{I_{b(\mathbf{A})}}\frac{w_{\mathbf{A}*\mathbf{B}}(x)}{w_{\mathbf{A}*\mathbf{B}}(x_{b(\mathbf{A})})}e^{i\xi\psi(g_{\mathbf{A}*\mathbf{B}}(x))}\,d\mu(x)\Big|^{2}$
$\iint_{I_{b(\mathbf{A})}^{2}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\frac{w_{\mathbf{a}_{k+1}}(y)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi(\psi(g_{\mathbf{A}*\mathbf{B}}(x))-\psi(g_{\mathbf{A}*\mathbf{B}}(y)))}\,d\mu(x)\,d\mu(y)$
$\iint_{I_{b(\mathbf{A})}^{2}}\frac{w_{\mathbf{a}_{k+1}}(x)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\frac{w_{\mathbf{a}_{k+1}}(y)}{w_{\mathbf{a}_{k+1}}(x_{b(\mathbf{A})})}\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi\psi^{\prime}(t_{1})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})(g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y))}\,d\mu(x)\,d\mu(y).$
To gain space, denote $W_{\mathbf{a}}(x,y):=\frac{w_{\mathbf{a}}(x)}{w_{\mathbf{a}}(x_{b(\mathbf{a})})}\frac{w_{\mathbf{a}}(y)}{w_{\mathbf{a}}(x_{b(\mathbf{a})})}$. It is more convenient to modify the signs in the phase so that everything becomes positive. To do so, notice now that this quantity is real. In particular, it is equal to it’s real part, which gives:
$\iint_{I_{b(\mathbf{A})}^{2}}W_{\mathbf{a}_{k+1}}(x,y)\sum_{\mathbf{B}}p_{\mathbf{B}}\cos\Big({\xi\psi^{\prime}(t_{1})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})(g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y))}\Big)d\mu(x)\,d\mu(y)$
$\iint_{I_{b(\mathbf{A})}^{2}}W_{\mathbf{a}_{k+1}}(x,y)\sum_{\mathbf{B}}p_{\mathbf{B}}\cos\Big({\xi|\psi^{\prime}(t_{1})||g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}\Big)d\mu(x)\,d\mu(y).$
And then triangle inequality and the fact that $|\text{Re}(z)|\leq|z|$ yields:
$\leq\iint_{I_{b(\mathbf{A})}^{2}}W_{\mathbf{a}_{k+1}}(x,y)\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}\cos\Big({\xi|\psi^{\prime}(t_{1})||g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}\Big)\Big|d\mu(x)\,d\mu(y)$
$\leq\iint_{I_{b(\mathbf{A})}^{2}}W_{\mathbf{a}_{k+1}}(x,y)\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}\Big|d\mu(x)\,d\mu(y)$
$\lesssim\iint_{I_{b(\mathbf{A})}^{2}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}\Big|d\mu(x)\,d\mu(y),$
where the last inequality comes from Gibbs estimates. Let us now renormalize the phases and get rid of the dependencies of the $(t_{j})_{j}$. Denote $x_{\mathbf{a}}:=g_{\mathbf{a}}(x_{b(\mathbf{a})})\in I_{\mathbf{a}}$ some reference points. For all $j\geq 1$, as $t_{j+1}\in I_{\mathbf{a}_{j+1}}$, we have by bounded $(C,\alpha)$-distortion property and mean value theorem for some $z_{\mathbf{a}_{j+1}}\in I$ that
$e^{-C\kappa_{+}^{\alpha n}}\leq\left|\frac{g_{\mathbf{a}_{j}^{\prime}\mathbf{b}_{j}}^{\prime}(t_{j+1})}{g_{\mathbf{a}_{j}^{\prime}\mathbf{b}_{j}}^{\prime}(x_{\mathbf{a}_{j+1}})}\right|\leq e^{C\kappa_{+}^{\alpha n}}.$
Similarly, by hypothesis we have that
$|\log\psi^{\prime}(x)-\log\psi^{\prime}(y)|\leq C_{\psi}|\xi|^{\tilde{\varepsilon}_{\psi}}|x-y|^{\alpha},$
for some small enough $\tilde{\varepsilon}_{\psi}>0$. One can choose it so small that $|\xi|^{\tilde{\varepsilon}_{\psi}}\leq\kappa_{+}^{-\alpha n/2}$. Then, for any $t_{1}\in I_{\mathbf{a}_{1}}$, we find that
$e^{-C_{\psi}\kappa_{+}^{\alpha n/2}}\leq e^{-C_{\psi}|\xi|^{\tilde{\varepsilon}_{\psi}}\kappa_{+}^{\alpha n}}\leq\left|\frac{\psi^{\prime}(t_{1})}{\psi^{\prime}(x_{\mathbf{a}_{1}})}\right|\leq e^{C_{\psi}|\xi|^{\tilde{\varepsilon}_{\psi}}\kappa_{+}^{\alpha n}}\leq e^{C_{\psi}\kappa_{+}^{\alpha n/2}}.$
Thus
$e^{-\tilde{C}\kappa_{+}^{\alpha n/2}}\leq e^{-kC\kappa_{+}^{\alpha n}-C_{\psi}\kappa_{+}^{\alpha n/2}}\leq\Big|\frac{\psi^{\prime}(t_{1})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})}{\psi^{\prime}(x_{\mathbf{a}_{1}})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(x_{\mathbf{a}_{2}})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(x_{\mathbf{a}_{k+1}})}\Big|\leq e^{kC\kappa_{+}^{\alpha n}+C_{\psi}\kappa_{+}^{\alpha n/2}}\leq e^{\tilde{C}\kappa_{+}^{\alpha n/2}}.$
Since by mean value theorem $|e^{x}-1|\leq|x|$ when $|x|\leq 1$, this implies for large enough $n$:
$\Bigg|\frac{|\psi^{\prime}(t_{1})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})|}{|\psi^{\prime}(x_{\mathbf{a}_{1}})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(x_{\mathbf{a}_{2}})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(x_{\mathbf{a}_{k+1}})|}-1\Bigg|\leq\tilde{C}\kappa_{+}^{\alpha n/2}.$
Thus if we denote
$\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y):=e^{-2\lambda nk}\xi|g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)||\psi^{\prime}(x_{\mathbf{a}_{1}})|,$
and
$\zeta_{j,\mathbf{A}}\left(\mathbf{b}\right):=e^{2\lambda n}|g_{\mathbf{a}_{j}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{a}_{j+1}})|,\quad j=1,2,\dots,k,$
then by the mean value theorem for some $z\in[x,y]$ we have
$|\xi\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|-\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})|$
$=|\xi||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|\left|\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})|-\psi^{\prime}(x_{\mathbf{a}_{1}})g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(x_{\mathbf{a}_{2}})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(x_{\mathbf{a}_{k+1}})\right|$
[…]
$\leq|\xi|\ |g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(x_{\mathbf{a}_{2}})\dots g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(x_{\mathbf{a}_{k+1}})g_{\mathbf{a}_{k+1}}^{\prime}(z)|\kappa_{+}^{\alpha n/3}.$
[…]
$\lesssim e^{(2k+1)\lambda n}g_{\mathbf{A}*\mathbf{B}}^{\prime}(z)\kappa_{+}^{\alpha n/6}.$
This is not a decaying term for all choice of words, but it is for most choices of words. More precisely, using the $1$-Lipschitz continuity of $x\mapsto e^{ix}$ and Proposition 2.1, we find:
$\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\xi\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}\Big|$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|$
$\qquad+\sum_{\mathbf{A},\mathbf{B}}p_{\mathbf{A}}p_{\mathbf{B}}\Big|e^{i\xi\psi^{\prime}(t_{1})|g_{\mathbf{a}_{1}^{\prime}\mathbf{b}_{1}}^{\prime}(t_{2})|\dots|g_{\mathbf{a}_{k}^{\prime}\mathbf{b}_{k}}^{\prime}(t_{k+1})||g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y)|}-e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|+2\sum_{\mathbf{A},\mathbf{B}}p_{\mathbf{A}}p_{\mathbf{B}}\mathbf{1}(g_{\mathbf{A}*\mathbf{B}}^{\prime}(x)e^{(2k+1)\lambda n}\geq\kappa_{+}^{-\alpha n/20})$
$\qquad+C\sum_{\mathbf{A},\mathbf{B}}p_{\mathbf{A}}p_{\mathbf{B}}\mathbf{1}(g_{\mathbf{A}*\mathbf{B}}^{\prime}(x)e^{(2k+1)\lambda n}\leq\kappa_{+}^{-\alpha n/20})e^{(2k+1)\lambda n}g_{\mathbf{A}*\mathbf{B}}^{\prime}(z)\kappa_{+}^{\alpha n/6}$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|$
$\qquad+C\sum_{\mathbf{C}\in\Sigma^{(2k+1)n}}p_{\mathbf{C}}\Big(e^{(2k+1)\lambda n}g_{\mathbf{C}}^{\prime}(z)\kappa_{+}^{\alpha n/20}\Big)^{t}+C\kappa_{+}^{\alpha n/10}$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|+C\kappa_{+}^{\alpha tn/20}e^{\varepsilon_{\gamma_{0}}nt^{2}}+C\kappa_{+}^{\alpha n/10}$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|+C(e^{-\frac{\alpha^{2}|\ln\kappa_{+}|^{2}}{\varepsilon_{\gamma_{0}}800}n}+\kappa_{+}^{\alpha n/10})$
by choosing $t:=\frac{\alpha|\ln\kappa_{+}|}{\varepsilon_{\gamma_{0}}20}$.
Thus, to conclude Fourier decay of the measure $\mu$, we just need to bound
$\sum_{\mathbf{A}}p_{\mathbf{A}}\iint_{I_{b(\mathbf{A})}^{2}}X_{x,y}(\mathbf{A})\,d\mu(x)\,d\mu(y),$
where $X_{x,y}(\mathbf{A})$ is the $\mathbb{R}_{+}$-valued random variable
$X_{x,y}(\mathbf{A}):=\Big|\sum_{\mathbf{B}}p_{\mathbf{B}}e^{i\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)\zeta_{1,\mathbf{A}}(\mathbf{b}_{1})\dots\zeta_{k,\mathbf{A}}(\mathbf{b}_{k})}\Big|.$
4.3. Applying the sum-product bound under non-concentration of derivatives
In the notation in Theorem 3.3, we can write our random variable as
$X_{x,y}(\mathbf{A})=\Big|{{\sum_{\mathbf{b}_{1}\in\mathcal{Z}_{1}}}}\dots{\sum_{\mathbf{b}_{k}\in\mathcal{Z}_{k}}}w_{\mathbf{b}_{1}}^{(1)}\dots w_{\mathbf{b}_{k}}^{(k)}e^{i\eta\zeta_{1}(\mathbf{b}_{1})\dots\zeta_{k}(\mathbf{b}_{k})}\Big|,$
where the map
$\zeta_{j}(\mathbf{b}):=\zeta_{j,\mathbf{A}}(\mathbf{b})=e^{2\lambda n}|g_{\mathbf{a}_{j}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{a}_{j+1}})|,$
defined on the sets $\mathcal{Z}_{j}:=\mathcal{Z}_{j,\mathbf{A}}=\{\mathbf{b}:\mathbf{a}_{j}^{\prime}\mathbf{b}^{\prime}\mathbf{a}_{j+1}\ \text{is admissible}\}$ and
$w_{\mathbf{b}_{j}}^{(j)}:=\mu(I_{\mathbf{b}_{j}})\in[0,1]\quad\text{and}\quad\eta:=\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y).$
Now, by Theorem 3.3, if there exists $0<\gamma<1$ and $\varepsilon_{1}>0$ such that if for all $j\in\{1,\dots,k\}$, the following non-concentration condition holds for when $|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|\geq 1$:
$\forall\sigma\in[|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|^{-2},|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|^{-\varepsilon_{1}}],\ \sum_{(\mathbf{b},\mathbf{c})\in\mathcal{Z}_{j,\mathbf{A}}^{2}}\mu(I_{\mathbf{b}})\mu(I_{\mathbf{c}})\mathbf{1}_{[-\sigma,\sigma]}(\zeta_{j,\mathbf{A}}(\mathbf{b})-\zeta_{j,\mathbf{A}}(\mathbf{c}))\leq\sigma^{\gamma},$ (4.2)
Then, there exists a constant $c>0$ that only depends on $\gamma$ such that
$\forall t>0,\ X_{x,y}(\mathbf{A})\leq c|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|^{-\varepsilon_{1}^{2}}+|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|^{-\frac{t^{3}}{2\sqrt{k}}}.\prod_{j=1}^{k}M_{j,\mathbf{A}}(t^{2}),$ (4.3)
where the $t$-moment $M_{j,\mathbf{A}}(t)$ is given by
$M_{j,\mathbf{A}}(t):=\sum_{\mathbf{b}\in\mathcal{Z}_{j,\mathbf{A}}}\mu(I_{\mathbf{b}})e^{t|\ln\zeta_{j,\mathbf{A}}(\mathbf{b})|}.$
We will now define the sets of well-distributed blocks and words on which the sum-product phenomenon applies, which requires us to finally fix parameters. The following Definition also reveals to us how the ultimate Fourier decay exponent depend on the hypotheses of the IFS:
Definition 4.1 (Choices of parameters).
Denote by $\Theta,\rho\in(0,1)$ the two constants given by our hypothesis (QNL). Recall that $\alpha\in(0,1)$ is an Hölder exponent of $\tau$ and that $\kappa_{+}:=\sup_{a\in\mathcal{A}}|g_{a}^{\prime}|_{\infty}\in(0,1)$. We then define
$\varepsilon_{0}:=\alpha|\ln\kappa_{+}|\cdot|\ln\rho|/10>0.$
Furthermore, we set $\gamma_{2}:=\Theta/4$. Theorem 3.3 then fixes some $k\in\mathbb{N}$ and $\varepsilon_{1}\in(0,1)$ (that are associated to $\gamma_{2}>0$).
With these fixed, we can now define
Definition 4.2 (Well-distributed words).
We say that a block is $\mathbf{A}=\mathbf{a_{0}}\dots\mathbf{a}_{k}\in\Sigma^{(k+1)n}$ is well-distributed if for all $j\in\llbracket 1,k\rrbracket$ and for all $x\in I$:
$\forall\sigma\in[e^{-4\varepsilon_{0}n},e^{-\varepsilon_{0}\varepsilon_{1}n/2}],\quad\sum_{(\mathbf{b},\mathbf{c})\in\mathcal{Z}_{j,\mathbf{A}}^{2}}w_{\mathbf{b}}(x)w_{\mathbf{c}}(x)\chi_{[-\sigma,\sigma]}(\ln\zeta_{j,\mathbf{A}}(\mathbf{b})-\ln\zeta_{j,\mathbf{A}}(\mathbf{c}))\leq\sigma^{\gamma_{2}}.$
Similarly, we call a couple of words $(\mathbf{a},\mathbf{d})\in(\Sigma^{n})^{2}$ well-distributed if, for all $x\in I$:
$\forall\sigma\in[e^{-4\varepsilon_{0}n},e^{-\varepsilon_{0}\varepsilon_{1}n/2}],\quad\underset{\mathbf{a}^{\prime}\mathbf{c}^{\prime}\mathbf{d}^{\prime}\in\Sigma^{3n}}{\underset{\mathbf{a}^{\prime}\mathbf{b}^{\prime}\mathbf{d}^{\prime}\in\Sigma^{3n}}{\sum_{(\mathbf{b},\mathbf{c})\in(\Sigma^{n+1})^{2}}}}w_{\mathbf{b}}(x)w_{\mathbf{c}}(x)\chi_{[-\sigma,\sigma]}(\ln g_{\mathbf{a}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{d}})-\ln g_{\mathbf{a}^{\prime}\mathbf{c}}^{\prime}(x_{\mathbf{d}}))\leq\sigma^{\gamma_{2}}.$
We will denote the set of well-distributed blocks by $\mathcal{W}_{n}^{k+1}$, and the set of well-distributed couples by $\mathcal{W}_{n}^{2}$.
We have the following crucial non-concentration property, whose proof is postponed to Section 5:
Proposition 4.3.
Define $\alpha_{reg}:=\varepsilon_{0}\varepsilon_{1}\gamma_{2}>0$. Then most blocks are well-distributed:
$\sum_{\mathbf{A}\notin\mathcal{W}_{n}^{k+1}}w_{\mathbf{A}}(x_{b(\mathbf{A})})\lesssim e^{-\alpha_{reg}n}.$
Assuming Proposition 4.3, we can now complete the proof of Theorem 1.1. In the following we will use the following notation:
$\kappa_{-}({\mathbf{a}_{k}})=\inf|g_{\mathbf{a}_{k}}^{\prime}|\quad\text{and}\quad\kappa_{+}({\mathbf{a}_{k}})=\sup|g_{\mathbf{a}_{k}}^{\prime}|.$
Let then $\tilde{\mathcal{W}}$ be those blocks $\mathbf{A}\in\Sigma^{(k+1)n}$ such that the last word $\mathbf{a}_{k+1}$ in the block $\mathbf{A}$ satisfies
$C_{\psi}e^{\lambda n}\kappa_{+}(\mathbf{a}_{k+1})\leq e^{\varepsilon_{0}n},$
where $\varepsilon_{\psi}>0$ comes from the condition
$\inf_{x}|\psi^{\prime}(x,\xi)|\gtrsim|\xi|^{-\varepsilon_{\psi}}=\exp(-\varepsilon_{\psi}c_{0}n),$
recall that $|\xi|=e^{c_{0}n}$ for $c_{0}=(2k+1)\lambda+\varepsilon_{0}$.
Finally, given any block $\mathbf{A}\in\Sigma^{(k+1)n}$, let $\mathcal{F}_{\mathbf{A}}$ be the set of pairs $(x,y)\in I_{b(\mathbf{A})}^{2}$ such that
$|x-y|>\exp(-(\lambda-\varepsilon_{\psi}c_{0})n)\kappa_{-}({\mathbf{a}_{k+1}})^{-1}\exp(-\varepsilon_{0}n/2).$
Lemma 4.4.
We can bound for any small enough $t>0$ that and large enough $n$:
$\sum_{\mathbf{A}}p_{\mathbf{A}}\iint_{I_{b(\mathbf{A})}^{2}}X_{x,y}(\mathbf{A})d\mu(x)\,d\mu(y)$
$\lesssim e^{-\alpha_{reg}n}+C_{\psi}^{t}2C_{\phi}e^{-(\varepsilon_{0}t+\varepsilon_{\gamma_{0}}t^{2})n}+2C_{\phi}e^{-(s_{\mu}(\lambda+\varepsilon_{\psi}c_{0}-\varepsilon_{0}/2)+t\lambda-\varepsilon_{\gamma_{0}}t^{2})n}$
$\qquad+\sum_{\mathbf{A}\in\mathcal{W}\cap\tilde{\mathcal{W}}}p_{\mathbf{A}}\sup_{(x,y)\in\mathcal{F}_{\mathbf{A}}}X_{x,y}(\mathbf{A}).$
Proof.
The proof combines the bounds from Proposition 5.1, Proposition 2.1 and Proposition 2.2 together with the bound $X_{x,y}(\mathbf{A})\leq 1$ due to $|e^{ix}|=1$ and that $\mu$ is a probability measure.
Firstly, by the non-concentration Proposition 5.1 and Gibbs property, we have
$\sum_{\mathbf{A}\notin\mathcal{W}}p_{\mathbf{A}}\lesssim e^{-\alpha_{reg}n},$
for some $\alpha_{reg}>0$, which gives the first term.
Secondly, by Proposition 2.1, we have for any small enough $t>0$ that
$\sum_{\mathbf{A}}p_{\mathbf{A}}\iint_{I_{b(\mathbf{A})}^{2}}X_{x,y}(\mathbf{A})\mathrm{\mathbf{1}}(\mathbf{a}_{k+1}:C_{\psi}e^{\lambda n}\kappa_{+}(\mathbf{a}_{k+1})>e^{\varepsilon_{0}n})\,d\mu(x)\,d\mu(y)$
$\leq C_{\psi}^{t}e^{-\varepsilon_{0}tn}\sum_{\mathbf{a}_{1}\dots\mathbf{a}_{k}}\mu(I_{\mathbf{a}_{1}})\dots\mu(I_{\mathbf{a}_{k}})\sum_{\mathbf{a}_{k+1}}\mu(I_{\mathbf{a}_{k+1}})(e^{\lambda n}\kappa_{+}(\mathbf{a}_{k+1}))^{t}$
$\leq C_{\psi}^{t}e^{-\varepsilon_{0}tn}2C_{\phi}e^{\varepsilon_{\gamma_{0}}t^{2}n}=C_{\psi}^{t}2C_{\phi}e^{-(\varepsilon_{0}t-\varepsilon_{\gamma_{0}}t^{2})n},$
where $\varepsilon_{\gamma_{0}}t^{2}<\varepsilon_{0}t$ as long as $t>0$ is small enough.
Finally, we can use the $s_{\mu}$-upper regularity of $\mu$ (Proposition 2.2) and that $\mu$ is a probability measure, and that $|e^{ix}|=1$ to get a bound:
$\sum_{\mathbf{A}}p_{\mathbf{A}}\iint_{\mathcal{F}_{\mathbf{A}}^{c}}X_{x,y}(\mathbf{A})\,d\mu(x)\,d\mu(y)$
$\leq\sum_{\mathbf{A}}p_{\mathbf{A}}\mu^{2\otimes}\big((x,y):|x-y|\leq\exp(-(\lambda-\varepsilon_{\psi}c_{0})n)\kappa_{-}({\mathbf{a}_{k+1}})^{-1}\exp(-\varepsilon_{0}n/2)\big)$
$\lesssim\sum_{\mathbf{A}}p_{\mathbf{A}}\exp(-s_{\mu}(\lambda-\varepsilon_{\psi}c_{0})n)\kappa_{-}({\mathbf{a}_{k+1}})^{-s_{\mu}}\exp(-s_{\mu}\varepsilon_{0}n/2)$
$\leq\exp(-ns_{\mu}(\lambda-\varepsilon_{\psi}c_{0}+\varepsilon_{0}/2))\exp(-t\lambda n)\sum_{\mathbf{a}_{1}\dots\mathbf{a}_{k}}\mu(I_{\mathbf{a}_{1}})\dots\mu(I_{\mathbf{a}_{k}})\sum_{\mathbf{a}_{k+1}}\mu(I_{\mathbf{a}_{k+1}})e^{t\lambda n}\kappa_{-}({\mathbf{a}_{k+1}})^{-t}$
$\leq\exp(-ns_{\mu}(\lambda-\varepsilon_{\psi}c_{0}+\varepsilon_{0}/2))\exp(-t\lambda n)2C_{\phi}\exp(\varepsilon_{\gamma_{0}}t^{2}n),$
as $\mu$ is a probability measure by Proposition 2.1 as long as $0<t<\min\{t_{\gamma_{0}},s_{\mu}\}$. Thus choosing $t>0$ and $\varepsilon_{\psi}>0$ small enough that
$s_{\mu}(\lambda-\varepsilon_{\psi}c_{0}+\varepsilon_{0}/2)+t\lambda-\varepsilon_{\gamma_{0}}t^{2}>0,$
we have exponential decay for this term. ∎
Let $\mathbf{A}\in\mathcal{W}\cap\tilde{\mathcal{W}}$ and $(x,y)\in\mathcal{F}_{\mathbf{A}}$. Then, since $c_{0}=(2k+1)\lambda+\varepsilon_{0}$, we can bound from below:
$|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|$ $=|e^{-2\lambda nk}\xi(g_{\mathbf{a}_{k+1}}(x)-g_{\mathbf{a}_{k+1}}(y))\psi^{\prime}(g_{\mathbf{a}_{1}}(x_{0}))|$
$\geq\exp((\lambda+\varepsilon_{0})n)(\inf|\psi^{\prime}|)\kappa_{-}({\mathbf{a}_{k+1}})|x-y|$
$\geq\exp((\lambda+\varepsilon_{0})n)\exp(-\varepsilon_{\psi}c_{0}n)\kappa_{-}({\mathbf{a}_{k+1}})|x-y|$
$\geq\exp(\varepsilon_{0}n)\exp((\lambda-\varepsilon_{\psi}c_{0})n)\kappa_{-}({\mathbf{a}_{k+1}})|x-y|,$
$\geq e^{\varepsilon_{0}n/2},$
and from above as $\mathbf{A}\in\tilde{\mathcal{W}}$ and $|x-y|\leq|I|$:
$|\eta_{\mathbf{a}_{k+1},\mathbf{a}_{1}}(x,y)|$ $\leq e^{-2\lambda nk}|\xi|(\sup|\psi^{\prime}|)(\sup|g_{\mathbf{a}_{k+1}}^{\prime}|)|x-y|$
$\lesssim e^{-2\lambda nk}e^{(2k+1)\lambda n+\varepsilon_{0}n}(\sup|\psi^{\prime}|)(\sup|g_{\mathbf{a}_{k+1}}^{\prime}|)|x-y|$
$\leq e^{\varepsilon_{0}n}C_{\psi}e^{\lambda n}\kappa_{+}({\mathbf{a}_{k+1}})$
$\leq e^{2\varepsilon_{0}n}.$
Thus by (4.3) we know that for all $t>0$:
$\sum_{\mathbf{A}\in\mathcal{W}\cap\tilde{\mathcal{W}}}p_{\mathbf{A}}\sup_{(x,y)\in\mathcal{F}_{\mathbf{A}}}X_{x,y}(\mathbf{A})\leq ce^{-\varepsilon_{0}\varepsilon_{1}^{2}n/2}+\sum_{\mathbf{A}}p_{\mathbf{A}}(\exp(\varepsilon_{0}n/2))^{-\frac{t^{3}}{2\sqrt{k}}}\prod_{j=1}^{k}M_{j,\mathbf{A}}(t^{2})$ (4.4)
so we have reduced Theorem 1.1 to the moment bound for $M_{j,\mathbf{A}}(t^{2})$. If $t\in(0,\varepsilon_{1})$, since $|\ln\zeta_{j,\mathbf{A}}\left(\mathbf{b}\right)|=|2\lambda n-\ln|g_{\mathbf{a}_{j}\mathbf{b}}^{\prime}\left(g_{\mathbf{a}_{j+1}}\left(x_{0}\right)\right)|$, we have by the chain rule and Gibbs property
$\mu(I_{\mathbf{a}_{j}})\mu(I_{\mathbf{b}})e^{t^{2}|\ln\zeta_{j,\mathbf{A}}(\mathbf{b})|}$ $\lesssim w_{\mathbf{a}_{j}^{\prime}\mathbf{b}}(x_{\mathbf{a}_{j+1}})\Big(e^{2\lambda n}g_{\mathbf{a}_{j}\mathbf{b}}^{\prime}\left(x_{\mathbf{a}_{j+1}}\right)\Big)^{t^{2}}$
Thus by Proposition 2.1 we have
$\sum_{\mathbf{a}_{1}\dots\mathbf{a}_{k}}\mu(I_{\mathbf{a}_{1}})\dots\mu(I_{\mathbf{a}_{k}})\prod_{j=1}^{k}M_{j,\mathbf{A}}(t^{2})$ $=\prod_{j=1}^{k}\sum_{\mathbf{a}_{j}}\mu(I_{\mathbf{a}_{j}})M_{j,\mathbf{A}}(t^{2})$
$=\prod_{j=1}^{k}\sum_{\mathbf{a}_{j}}\mu(I_{\mathbf{a}_{j}})\sum_{\mathbf{b}\in\mathcal{Z}_{j,\mathbf{A}}}\mu(I_{\mathbf{b}})e^{t|\ln\zeta_{j,\mathbf{A}}(\mathbf{b})|}$
$=\prod_{j=1}^{k}\sum_{\mathbf{a}_{j}}\sum_{\mathbf{b}}\mu(I_{\mathbf{a}_{j}})\mu(I_{\mathbf{b}})e^{t^{2}|\ln\zeta_{j,\mathbf{A}}(\mathbf{b})|}$
$\lesssim\Big(2C_{\phi}e^{2\varepsilon_{\gamma_{0}}nt^{4}}\Big)^{k}.$
Thus Proposition 2.1 gives
$\sum_{\mathbf{A}}p_{\mathbf{A}}(\exp(\varepsilon_{0}n/2))^{-\frac{t^{3}}{2\sqrt{k}}}\prod_{j=1}^{k}M_{j,\mathbf{A}}(t^{2})$
$=\sum_{\mathbf{a}_{k+1}}\mu(I_{\mathbf{a}_{k+1}})(\exp(\varepsilon_{0}n/2))^{-\frac{t^{3}}{2\sqrt{k}}}\sum_{\mathbf{a}_{1}\dots\mathbf{a}_{k}}\mu\left(I_{\mathbf{a}_{1}}\right)\cdots\mu\left(I_{\mathbf{a}_{k}}\right)\prod_{j=1}^{k}M_{j,\mathbf{A}}(t^{2})$
$\lesssim\sum_{\mathbf{a}_{k+1}}\mu(I_{\mathbf{a}_{k+1}})(\exp(\varepsilon_{0}n/2))^{-\frac{t^{3}}{2\sqrt{k}}}\Big(2C_{\phi}e^{2\varepsilon_{\gamma_{0}}nt^{4}}\Big)^{k}$
$=(2C_{\phi})^{k}\exp\Big(-n(\frac{t^{3}}{4\sqrt{k}}\varepsilon_{0}-2k\varepsilon_{\gamma_{0}}t^{4})\Big).$
Now we see that
$\frac{t^{3}}{4\sqrt{k}}\varepsilon_{0}-2k\varepsilon_{\gamma_{0}}t^{4}>0\quad\text{if and only if}\quad 0<t<\frac{1}{8k^{3/2}\varepsilon_{\gamma_{0}}}\varepsilon_{0},$
which is possible if $t>0$ is chosen small enough. This completes the proof of Theorem 1.1.
5. Reducing multiscale non-concentration to larger probability space
We are now left with proving Proposition 4.3. Recall its formulation:
Proposition 5.1.
Define $\alpha_{reg}:=\varepsilon_{0}\varepsilon_{1}\gamma_{2}>0$. Then most blocks are well-distributed:
$\sum_{\mathbf{A}\notin\mathcal{W}_{n}^{k+1}}w_{\mathbf{A}}(x_{b(\mathbf{A})})\lesssim e^{-\alpha_{reg}n},$
recall Definition 4.2 for the Definition of well-distributed blocks of words.
This estimate will be proved using by a sequence of reductions, adapting arguments found e.g. in [13, 41, 27, 10]. For clarity, we will drop the conditions on admissiblity of words in the sums, they are implicit. We denote $w_{\mathbf{a}}:=w_{\mathbf{a}}(x_{b(\mathbf{a})})$.
The first step is to reduce this estimate to an easier bound in a larger probability, using Markov’s inequality.
Lemma 5.2.
Suppose that, for $n$ large enough and for every $\sigma\in[e^{-4\varepsilon_{0}n},e^{-\varepsilon_{0}\varepsilon_{1}n/2}]$:
$\sum_{(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})\in(\Sigma^{n})^{4}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}\chi_{[-\sigma,\sigma]}(\ln g_{\mathbf{a}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{d}})-\ln g_{\mathbf{a}^{\prime}\mathbf{c}}^{\prime}(x_{\mathbf{d}}))\leq\sigma^{2\gamma_{2}}.$
Then the conclusion of Proposition 5.1 holds.
Proof.
First of all, notice that well-distributed blocks are concatenations of well-distributed couples. In other words:
$\mathcal{W}_{n}^{k+1}=\bigcap_{j=1}^{k}\left\{\mathbf{A}\in\Sigma^{(k+1)n}\mid\ (\mathbf{a}_{j-1},\mathbf{a}_{j})\in\mathcal{W}_{n}^{2}\right\}.$
We then find (using Gibbs estimates $w_{\mathbf{a}}\simeq\mu(I_{\mathbf{a}})$ and $w_{\mathbf{A}}\simeq w_{\mathbf{a}_{1}}\dots w_{\mathbf{a}_{k}}$):
$\sum_{\mathbf{A}\in\Sigma^{n(k+1)}\setminus\mathcal{W}_{n}^{k+1}}w_{\mathbf{A}}\leq\sum_{j=1}^{k}\underset{(\mathbf{a}_{j-1},\mathbf{a}_{j})\notin\mathcal{W}_{n}^{2}}{\sum_{\mathbf{A}\in\Sigma^{n(k+1)}}}w_{\mathbf{A}}\lesssim{\sum_{(\mathbf{a},\mathbf{c})\in(\Sigma^{n})^{2}\setminus\mathcal{W}_{n}^{2}}}w_{\mathbf{a}}w_{\mathbf{c}}.$
If we define, for all $l\in\mathbb{N}$, the set
$\mathcal{W}_{n,l}^{2}:=\Big\{(\mathbf{a},\mathbf{d})\in(\Sigma^{n})^{2},\ \sum_{\mathbf{b},\mathbf{c}\in\Sigma^{n}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{[-e^{-l},e^{-l}]}(\ln g_{\mathbf{a}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{d}})-\ln g_{\mathbf{a}^{\prime}\mathbf{c}}^{\prime}(x_{\mathbf{d}}))\leq e^{-(l+1)\gamma_{2}}\Big\},$
then we have the inclusion
$\mathcal{W}_{n}^{2}\supset\bigcap_{\varepsilon_{0}\varepsilon_{1}n/2\leq l\leq 4\varepsilon_{0}n}\mathcal{W}_{n,l}^{2},$
which gives, by Markov’s inequality:
${\sum_{(\mathbf{a},\mathbf{c})\in(\Sigma^{n})^{2}\setminus\mathcal{W}_{n}^{2}}}w_{\mathbf{a}}w_{\mathbf{c}}\leq\underset{\varepsilon_{0}\varepsilon_{1}n/2\leq l\leq 4\varepsilon_{0}n}{\sum_{l\in\mathbb{N}}}{\sum_{(\mathbf{a},\mathbf{d})\in(\Sigma^{n})^{2}\setminus\mathcal{W}_{n,l}^{2}}}w_{\mathbf{a}}w_{\mathbf{d}}$
$\leq\underset{\varepsilon_{0}\varepsilon_{1}n/2\leq l\leq 4\varepsilon_{0}n}{\sum_{l\in\mathbb{N}}}e^{(l+1)\gamma_{2}}{\sum_{(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})\in(\Sigma^{n})^{4}}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}\ \chi_{[-e^{-l},e^{-l}]}(e^{2\lambda n}g_{\mathbf{a}^{\prime}\mathbf{b}}^{\prime}(x_{\mathbf{d}})-e^{2\lambda n}g_{\mathbf{a}^{\prime}\mathbf{c}}^{\prime}(x_{\mathbf{d}}))$
$\leq\underset{\varepsilon_{0}\varepsilon_{1}n/2\leq l\leq 4\varepsilon_{0}n}{\sum_{l\in\mathbb{N}}}e^{(l+1)\gamma_{2}}e^{-2l\gamma_{2}}\lesssim e^{-\varepsilon_{0}\varepsilon_{1}\gamma_{2}n}.$
Recall that we can write $\ln g_{\mathbf{a}}^{\prime}(x)=-S_{n}\tau\circ g_{\mathbf{a}}$. The bound in the hypothesis of the previous Lemma can then be rewritten in terms of Birkhoff sums. We check in the next Lemma that the desired bound holds under (QNL).
Lemma 5.3.
Suppose (QNL). Then for every $\sigma\in[e^{-4\varepsilon_{0}n},e^{-\varepsilon_{0}\varepsilon_{1}n/2}]$:
$\sum_{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\in\Sigma^{n}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}\ \chi_{[-\sigma,\sigma]}(S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}}))\lesssim\sigma^{\Theta/2}.$
Proof.
We have, by Cauchy–Schwarz applied to the sum on $\mathbf{a},\mathbf{b},\mathbf{c}$:
$\Big(\sum_{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\in\Sigma^{n}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}\ \chi_{[-\sigma,\sigma]}(S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}}))\Big)^{2}$
$\lesssim\sum_{\mathbf{a},\mathbf{b},\mathbf{c}\in\Sigma^{n}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}\Big(\sum_{\mathbf{d}\in\Sigma^{n}}w_{\mathbf{d}}\chi_{[-\sigma,\sigma]}(S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}}))\Big)^{2}$
$\lesssim\sum_{\mathbf{a,b,c,d,e}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-\sigma,\sigma]}(S_{2n}\tau(g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}}))-S_{2n}\tau(g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}})))\chi_{[-\sigma,\sigma]}(S_{2n}\tau(g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{e}}))-S_{2n}\tau(g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{e}})))$
$\leq\sum_{\mathbf{a,b,c,d,e}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-2\sigma,2\sigma]}(S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{e}})+S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{e}})).$
Notice then that
$S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})=S_{n}\tau\circ g_{\mathbf{a}}(g_{\mathbf{b}}x_{\mathbf{d}})+S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}}).$
By bounded $(C,\alpha)$-distortions
$S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{e}})=S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{e}})+\mathcal{O}(\text{diam}(I_{\mathbf{b}})^{\alpha}),$
which gives, since $\text{diam}(I_{\mathbf{b}})^{\alpha}\lesssim\kappa_{+}^{\alpha n}\leq e^{-8\varepsilon_{0}n}\leq\sigma^{2}$ as $\varepsilon_{0}\leq\alpha|\ln\kappa_{+}|/8$:
$S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{b}}(x_{\mathbf{e}})+S_{2n}\tau\circ g_{\mathbf{a}^{\prime}\mathbf{c}}(x_{\mathbf{e}})$
$=S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{e}})+S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{e}})+\mathcal{O}(\sigma^{2}).$
In particular, for $n$ large enough, we find:
$\sum_{\mathbf{a,b,c,d,e}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-2\sigma,2\sigma]}(S_{2n}\tau\circ g_{\mathbf{a}\mathbf{b}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}\mathbf{c}}(x_{\mathbf{d}})-S_{2n}\tau\circ g_{\mathbf{a}\mathbf{b}}(x_{\mathbf{e}})+S_{2n}\tau\circ g_{\mathbf{a}\mathbf{c}}(x_{\mathbf{e}}))$
$\leq\sum_{\mathbf{a,b,c,d,e}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-3\sigma,3\sigma]}(S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{e}})+S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{e}}))$
$\leq\sum_{\mathbf{b,c,d,e}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-3\sigma,3\sigma]}(S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{e}})+S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{e}}))$
Under (QNL), this last sum is bounded by a multiple of $\sigma^{\Theta}+\rho^{n}\leq 2\sigma^{\Theta}$ since $\sigma$ is slowly decaying comparing to $\rho$: $\sigma^{\Theta}\geq e^{-4\varepsilon_{0}\Theta n}\geq\rho^{n}$ since $\varepsilon_{0}\leq|\ln(\rho)|/4$. This gives the desired bound since $\sigma^{\Theta/2}=\sigma^{2\gamma_{2}}$. ∎
This concludes the proof of Theorem 1.1. In the next section, we show how to check (QNL), first under (MNL) in a truly $C^{1+\alpha}$ IFS setting with finite branches, and then under (UNI) in a $C^{2}$-setting with infinite branches and the exponential moment condition.
6. Proofs of the non-concentration estimates
6.1. $C^{1+\alpha}$ IFSs: Using (MNL) for non-concentration
This subsection is devoted to the proof of Theorem 1.3, that is, the proof that (MNL) implies (QNL) in the setting where $|\mathcal{A}|<\infty$. To simplify, suppose that the dynamics is also fully branched: the subshift of finite type setting is similar.
Let us denote $0<\kappa_{+},\kappa_{-}<1$ such that for all $a\in\mathcal{A}$, $\kappa_{+}<g_{a}^{\prime}<\kappa_{-}$. Recall that we suppose that $\mu$ is the equilibrium state associated with $\delta\tau$ for $\delta:=\dim_{H}(F_{\Phi})$, so recall that in particular we have bounds:
$\forall\mathbf{c}\in\mathcal{A}^{*},\ \mu(I_{\mathbf{c}})\simeq|I_{\mathbf{c}}|^{\delta}$
and, by our one-dimensional setting:
$\forall\mathbf{c}\in\mathcal{A}^{*},\ |I_{\mathbf{c}}|\simeq|g_{\mathbf{c}}^{\prime}(0)|\quad;\quad\forall\mathbf{a},\mathbf{b}\in\mathcal{A}^{*},\ |I_{\mathbf{ab}}|\simeq|I_{\mathbf{a}}|\cdot|I_{\mathbf{b}}|.$
Recall we defined for $\mathbf{a}\in\mathcal{A}^{n}$ and $\mathbf{a}\to x,y$, $X_{\mathbf{a}}(x,y):=S_{n}\tau\circ g_{\mathbf{a}}(x)-S_{n}\tau\circ g_{\mathbf{a}}(y)$. Notice the relation:
$X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)=S_{n_{1}+n_{2}}\tau\circ g_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x)-S_{n_{1}+n_{2}}\tau\circ g_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(y)$
$(S_{n_{1}}\tau\circ g_{\tilde{\mathbf{a}}})(g_{\widehat{\textbf{a}}}x)-(S_{n_{1}}\tau\circ g_{\tilde{\mathbf{a}}})(g_{\widehat{\textbf{a}}}y)+S_{n_{2}}\tau\circ g_{\widehat{\mathbf{a}}}(x)-S_{n_{2}}\tau\circ g_{\widehat{\mathbf{a}}}(y)$
$X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x,g_{\widehat{\mathbf{a}}}y)+X_{\widehat{\mathbf{a}}}(x,y).$
Notice further that since $\tau\in C^{\delta}$, we have, uniformly in $\mathbf{a}\in\mathcal{A}^{*}$ :
$|X_{\mathbf{a}}(x,y)|\lesssim|x-y|^{\delta}.$
For the duration of this subsection 6.1, we will suppose that (MNL) holds. This is intuitively a very strong nonconcentration/spreading statement in $x$ of the function $X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})$, since the pushforward by this function of the measure $\mu$ flattens it to an absolutely continuous measure. (MNL) implies in particular a more familliar “UNI” statement everywhere and at every scales.
Lemma 6.1.
For all $N\geq 1$, there exists $\widehat{\mathbf{a}},\widehat{\mathbf{b}}\in\mathcal{A}^{N}$ such that for all $\textbf{c}\in\mathcal{A}^{*}$ and for all $y\in I_{\mathbf{c}}\cap F$, there exists $x\in I_{\mathbf{c}}\cap F$ such that
$|X_{\widehat{\mathbf{a}}}(x,y)-X_{\widehat{\mathbf{b}}}(x,y)|\geq\frac{1}{3C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$
Proof.
Let $N\geq 1$ and let $\mathbf{a},\mathbf{b}$ be such that (MNL) is satisfied. Fix any $y\in F$. Now consider $I:=[-\frac{1}{3C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}),\frac{1}{3C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})]$. Then
$\mu(x\in F,\ X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)\in I)\leq C_{\text{(MNL)}}|I|=\frac{2}{3}\mu(I_{\mathbf{c}}).$
This ensure that the set $\{x\in F,\ X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)\in I\}$ can not contain $I_{\mathbf{c}}\cap F$. ∎
Thanks to this seperation condition, we can show that we can always find a word $\widehat{\mathbf{a}}\in\mathcal{A}^{N}$ such that $X_{\widehat{\mathbf{a}}}$ can avoid some given small enough intervals.
Lemma 6.2.
Let $y\in F$. Consider some function $t:F\rightarrow\mathbb{R}$.
Then for all $N\geq 1$, for all $\mathbf{c}\in\mathcal{A}^{*}$, there exists $\widehat{\mathbf{a}}\in\mathcal{A}^{N}$ such that:
$\exists x\in I_{\mathbf{c}}\cap F,\ |X_{\widehat{\mathbf{a}}}(x,y)-t(x)|\geq\frac{1}{10C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$
Proof.
Let $y\in F$, and $t:F\rightarrow\mathbb{R}$. Let $N\geq 1$ and let $\widehat{\mathbf{a}},\widehat{\mathbf{b}}\in\mathcal{A}^{N}$ be given by the previous Lemma. Let $\mathbf{c}\in\mathcal{A}^{*}$. We then know that there exists $x\in I_{\mathbf{c}}$ such that $|X_{\widehat{\mathbf{a}}}(x,y)-X_{\widehat{\mathbf{b}}}(x,y)|\geq\frac{1}{3C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$ It follows that $|X_{\widehat{\mathbf{a}}}(x,y)-t(x)|\geq\frac{1}{10C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})$ or $|X_{\widehat{\mathbf{b}}}(x,y)-t(x)|\geq\frac{1}{10C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})$. ∎
Lemma 6.3.
Let $A\geq 1$. There exists $N:=N(A)\geq 1$ large enough such that the following holds.
Let $t\in C^{\delta}(F,\mathbb{R})$ such that $\|t\|_{C^{\delta}(F,\mathbb{R})}\leq A$. For all $\mathbf{c}\in\mathcal{A}^{*}$ and $y\in I_{\mathbf{c}}$:
$\exists(\widehat{\mathbf{a}},\widehat{\mathbf{c}})\in(\mathcal{A}^{N})^{2},\ \forall x\in I_{\mathbf{c}\widehat{\mathbf{c}}},\ \forall\tilde{\mathbf{a}}\in\mathcal{A}^{*},\ |X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)-t(x)|\geq\frac{1}{100C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$
Proof.
Let $A\geq 1$. Let $N\geq 1$, we will fix it large enough during the proof depending on $A$. Let $t\in C^{\delta}$ be such that $\|t\|_{C^{\delta}}\leq A$. Let $\mathbf{c}\in\mathcal{A}^{*}$, and let $y\in I_{\mathbf{c}}$.
We know already that there exists $\widehat{\mathbf{a}}\in\mathcal{A}^{N}$ and $x_{0}\in I_{\mathbf{c}}$ such that
$|X_{\widehat{\mathbf{a}}}(x_{0},y)-t(x_{0})|\geq\frac{1}{10C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$
To conclude, we just need to show that this bound still holds under perturbations. Let us define $\widehat{c}\in\mathcal{A}^{N}$ as the word for wich $x_{0}\in I_{\mathbf{c}\widehat{\mathbf{c}}}$. We have three terms to study.
  1. (1)
    First of all, for all $x\in I_{\mathbf{c}\widehat{\mathbf{c}}}$, we have a bound
    $|t(x)-t(x_{0})|\leq A|I_{\mathbf{c}\widehat{\mathbf{c}}}|^{\delta}\lesssim A\kappa_{-}^{\delta N}\mu(I_{\mathbf{c}})\leq\frac{1}{100C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})$
    if $N$ is large enough depending on $A$.
  2. (2)
    Then, for any $\tilde{\mathbf{a}}\in\mathcal{A}^{*}$, since $g_{\widehat{\mathbf{a}}}(x_{0}),g_{\widehat{\mathbf{a}}}(y)\in I_{\widehat{\mathbf{a}}\mathbf{c}}$:
    $|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{0},y)-X_{\widehat{\mathbf{a}}}(x_{0},y)|=|X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0}{},g_{\widehat{\mathbf{a}}}y)|\lesssim|I_{\mathbf{a}\mathbf{c}}|^{\delta}\lesssim\kappa_{-}^{\delta N}\mu(I_{\mathbf{c}})\leq\frac{1}{100C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})$
    if $N$ is large enough depending on $A$.
  3. (3)
    Finally, for any $\tilde{\mathbf{a}}\in\mathcal{A}^{*}$ and for all $x\in I_{\mathbf{c}\widehat{\mathbf{c}}}$:
    $|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{0},y)-X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)|=|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{0},x)|\lesssim|I_{\mathbf{c}\widehat{\mathbf{c}}}|^{\delta}\lesssim\kappa^{\delta N}\mu(I_{\mathbf{c}})\leq\frac{1}{100C_{\text{(MNL)}}}\mu(I_{\mathbf{c}})$
    if $N$ is chosen large enough depending on $A$.
Let us now fix $N=N(A)$ so that all three bounds holds. We can then write, for any $\tilde{\mathbf{a}}\in\mathcal{A}^{*}$ and $x\in I_{\mathbf{c}\widehat{\mathbf{c}}}$:
$|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)-t(x)|\geq|X_{\widehat{\mathbf{a}}}(x_{0},y)-t(x_{0})|$
$|t(x)-t(x_{0})|-|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{0},y)-X_{\widehat{\mathbf{a}}}(x_{0},y)|-|X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{0},y)-X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)|$
$\geq\frac{1}{100C_{\text{(MNL)}}}\mu(I_{\mathbf{c}}).$
which concludes the proof. ∎
Before concluding, let us recall that all the maps $X_{\mathbf{a}}(\cdot,y)$ are $C^{\delta}$, and this uniformly in the choice of $\mathbf{a}$. It follows that the previous Lemma holds for $t:=X_{\mathbf{a}}(\cdot,y)$ with a uniform $N$.
For the rest of the section let us define, for any $\sigma>0$:
$n(\sigma):=\max\{n\geq 1,\ \forall\mathbf{c}\in\mathcal{A}^{n}\ |\ \mu(I_{\mathbf{c}})\geq 100C_{\text{(MNL)}}\sigma\}.$
Notice that we have $\mu(I_{\mathbf{c}})\simeq|I_{\mathbf{c}}|^{\delta}\geq\kappa_{-}^{\delta n}$, so that $n(\sigma)\geq c_{0}|\ln(\sigma)|$ for some $c_{0}$.
As an immediate Corollary, we get the following Lemma.
Lemma 6.4.
Let $A\geq 1$. There exists $N:=N(A)\geq 1$ large enough such that the following holds.
Let $t\in C^{\delta}(F,\mathbb{R})$ such that $\|t\|_{C^{\delta}(F,\mathbb{R})}\leq A$. Let $\sigma>0$, and denote $I(x):=[t(x)-\sigma,t(x)+\sigma]$. For all $\mathbf{c}\in\mathcal{A}^{*}$ with $|\mathbf{c}|\leq n(\sigma)$, and for all $y\in I_{\mathbf{c}}$:
$\exists(\widehat{\mathbf{a}},\widehat{\mathbf{c}})\in(\mathcal{A}^{N})^{2},\ \forall x\in I_{\mathbf{c}\widehat{\mathbf{c}}},\ \forall\tilde{\mathbf{a}}\in\mathcal{A}^{*},\ X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,y)\notin I(x).$
The idea now, to prove our main estimate, is to repetitively apply our previous Lemma. Recall that $x_{\mathbf{c}}$ is some point in $I_{\mathbf{c}}$ and $w_{\mathbf{c}}(x)\simeq\mu(I_{\mathbf{c}})$.
Proposition 6.5.
Let $\sigma>0$ be small enough. Let $t\in C^{\delta}(F,\mathbb{R})$ such that $\|t\|_{C^{\delta}}$. Fix $A:=\|t\|_{C^{\delta}}+\sup_{\mathbf{a}\in\mathcal{A}^{\delta}}\sup_{y\in F}\|X_{\mathbf{a}}(\cdot,y)\|_{C^{\delta}}$.
Define $I(x):=[t(x)-\sigma,t(x)+\sigma]$. We have, for all $x_{0},y_{0},z_{0}\in F$ and for all $n\geq 1$:
$\sum_{\mathbf{a}\in\mathcal{A}^{n}}\sum_{\mathbf{c}\in\mathcal{A}^{n}}w_{\mathbf{a}}(x_{0})w_{\mathbf{c}}(y_{0})\mathbf{1}_{I(x_{\mathbf{c}})}\Big(X_{\mathbf{a}}(x_{\mathbf{c}},z_{0})\Big)\leq C_{0}(\sigma^{\gamma}+\rho^{n})$
for some $\gamma,\rho\in(0,1),C_{0}\geq 1$, depending on $A$.
Proof.
Fix $A:=\|t\|_{C^{\delta}}+\sup_{\mathbf{a}\in\mathcal{A}^{\delta}}\sup_{y\in F}\|X_{\mathbf{a}}(\cdot,y)\|_{C^{\delta}}$, and consider $N:=N(A)$ given by the previous Lemma. Define $m:=\min(n,n(\sigma))/(2N)$. We wish to show
$\sum_{\mathbf{a}\in\mathcal{A}^{n}}\sum_{\mathbf{c}\in\mathcal{A}^{n}}w_{\mathbf{a}}(x_{0})w_{\mathbf{c}}(y_{0})\mathbf{1}_{I(x_{\mathbf{c}})}\Big(X_{\mathbf{a}}(x_{\mathbf{c}},z_{0})\Big)\leq\rho^{m},$
for some $\rho\in(0,1)$. This will imply our desired bound, since then
$\rho^{m}=\max(\rho^{n/(2N)},\rho^{n(\sigma)/(2N)})\leq\rho^{n/(2N)}+\rho^{n(\sigma)/N}\lesssim\rho^{n/(2N)}+\sigma^{\gamma}$
for some $\gamma\in(0,1)$.
To prove this bound, we are going to use our previous Lemma $m$ times. Suppose that $n\geq N$ and that $\sigma$ is small enough so that $n(\sigma)\geq N$. First of all, we know that there exists $(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})\in(\mathcal{A}^{N})^{2}$ such that:
$\forall\tilde{\mathbf{a}}\in\mathcal{A}^{n-N},\ \forall x\in I_{\widehat{\mathbf{c}}},\ X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x,z_{0})\notin I(x).$
We can then write:
$\sum_{\mathbf{a},\mathbf{c}\in\mathcal{A}^{n}}w_{\mathbf{a}}(x_{0})w_{\mathbf{c}}(y_{0})\mathbf{1}_{I(x_{\mathbf{c}})}\Big(X_{\mathbf{a}}(x_{\mathbf{c}},z_{0})\Big)$
$\sum_{\widehat{\mathbf{a}},\widehat{\mathbf{c}}\in\mathcal{A}^{N}}w_{\widehat{\mathbf{a}}}(x_{0})w_{\widehat{\mathbf{c}}}(y_{0})\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}}y_{0})\mathbf{1}_{I(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},z_{0})\Big)$
$\underset{(\widehat{\mathbf{a}},\widehat{\mathbf{c}})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}},\widehat{\mathbf{c}}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}}(x_{0})w_{\widehat{\mathbf{c}}}(y_{0})\Bigg(\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}}y_{0})\mathbf{1}_{I(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},z_{0})\Big)\Bigg).$
[…]
for $t_{\widehat{\mathbf{a}}}(x):=t(x)-X_{\widehat{\mathbf{a}}}(x,z_{0})$. Notice that $\|t_{\widehat{\mathbf{a}}}\|_{C^{\delta}}\leq A$. We can then write
$\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}}y_{0})\mathbf{1}_{I(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},z_{0})\Big)$
$\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}}y_{0})\mathbf{1}_{I_{\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},g_{\widehat{\mathbf{a}}}z_{0})\Big),$
[…]
$\forall\tilde{\mathbf{a}}\in\mathcal{A}^{n-2N},\ \forall x\in I_{\widehat{\mathbf{a}}\widehat{\mathbf{c}}\widehat{\mathbf{c}}^{(2)}},\ X_{\tilde{\mathbf{a}}\widehat{\mathbf{a}}^{(2)}}(x,g_{\widehat{\mathbf{a}}}z_{0})\notin I_{\widehat{\mathbf{a}}}(x).$
It follows that:
$\underset{(\widehat{\mathbf{a}},\widehat{\mathbf{c}})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}},\widehat{\mathbf{c}}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}}(x_{0})w_{\widehat{\mathbf{c}}}(y_{0})\Bigg(\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}}y_{0})\mathbf{1}_{I_{\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},g_{\widehat{\mathbf{a}}}z_{0})\Bigg)$
$\underset{(\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{1}}(x_{0})w_{\widehat{\mathbf{c}}_{1}}(y_{0})\sum_{\tilde{\mathbf{a}}_{2},\tilde{\mathbf{c}}_{2}\in\mathcal{A}^{N}}w_{\widehat{\mathbf{a}}_{2}}(g_{\widehat{\mathbf{a}}_{1}}x_{0})w_{\widehat{\mathbf{c}}_{2}}(g_{\widehat{\mathbf{c}}_{1}}y_{0})\Bigg($
$\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-2N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}_{2}\widehat{\mathbf{a}}_{1}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}_{2}\widehat{\mathbf{c}}_{1}}y_{0})\mathbf{1}_{I_{\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},g_{\widehat{\mathbf{a}}}z_{0})\Big)\Bigg)$
$\underset{(\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{1}}(x_{0})w_{\widehat{\mathbf{c}}_{1}}(y_{0})\underset{(\widehat{\mathbf{a}}_{2},\widehat{\mathbf{c}}_{2})\neq(\widehat{\mathbf{a}}^{(2)},\widehat{\mathbf{c}}^{(2)})}{\sum_{\tilde{\mathbf{a}}_{2},\tilde{\mathbf{c}}_{2}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{2}}(g_{\widehat{\mathbf{a}}_{1}}x)w_{\widehat{\mathbf{c}}_{2}}(g_{\widehat{\mathbf{c}}_{1}}x)\Bigg($
$\sum_{\tilde{\mathbf{a}},\tilde{\mathbf{c}}\in\mathcal{A}^{n-2N}}w_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}_{2}\widehat{\mathbf{a}}_{1}}x_{0})w_{\tilde{\mathbf{c}}}(g_{\widehat{\mathbf{c}}_{2}\widehat{\mathbf{c}}_{1}}y_{0})\mathbf{1}_{I_{\widehat{\mathbf{a}}}(x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}})}\Big(X_{\tilde{\mathbf{a}}}(g_{\widehat{\mathbf{a}}}x_{\widehat{\mathbf{c}}\tilde{\mathbf{c}}},g_{\widehat{\mathbf{a}}}z_{0})\Big)\Bigg).$
The previous argument is then iterated $m$ times. After $m$ iterations, we get a nested sum that we can bound by:
$\underset{(\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}}_{1},\widehat{\mathbf{c}}_{1}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{1}}(x_{0})w_{\widehat{\mathbf{c}}_{1}}(y_{0})\underset{(\widehat{\mathbf{a}}_{2},\widehat{\mathbf{c}}_{2})\neq(\widehat{\mathbf{a}}^{(2)},\widehat{\mathbf{c}}^{(2)})}{\sum_{\tilde{\mathbf{a}}_{2},\tilde{\mathbf{c}}_{2}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{2}}(g_{\widehat{\mathbf{a}}_{1}}x_{0})w_{\widehat{\mathbf{c}}_{2}}(g_{\widehat{\mathbf{c}}_{1}}y_{0})\dots$
$\dots\underset{(\widehat{\mathbf{a}}_{m},\widehat{\mathbf{c}}_{m})\neq(\widehat{\mathbf{a}}^{(m)},\widehat{\mathbf{c}}^{(m)})}{\sum_{\widehat{\mathbf{a}}_{m},\widehat{\mathbf{c}}_{m}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}_{m}}(g_{\widehat{\mathbf{a}}_{1}\dots\widehat{\mathbf{a}}_{m-1}}x_{0})w_{\widehat{\mathbf{c}}_{m}}(g_{\widehat{\mathbf{c}}_{1}\dots\widehat{\mathbf{c}}_{m-1}}y_{0}).$
Now, denoting $\rho:=1-\inf_{\mathbf{a}\in\mathcal{A}^{N}}\inf_{x}w_{\mathbf{a}}(x)w_{\mathbf{c}}(x)$, we have
$\underset{(\widehat{\mathbf{a}},\widehat{\mathbf{c}})\neq(\widehat{\mathbf{a}}^{(1)},\widehat{\mathbf{c}}^{(1)})}{\sum_{\widehat{\mathbf{a}},\widehat{\mathbf{c}}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}}(x_{0})w_{\widehat{\mathbf{c}}}(y_{0})={\sum_{\widehat{\mathbf{a}},\widehat{\mathbf{c}}\in\mathcal{A}^{N}}}w_{\widehat{\mathbf{a}}}(x_{0})w_{\widehat{\mathbf{c}}}(y_{0})-w_{\widehat{\mathbf{a}}^{(1)}}(x_{0})w_{\widehat{\mathbf{c}}^{(1)}}(y_{0})$
$(\mathcal{L}_{\varphi}^{N})(1)^{2}-w_{\widehat{\mathbf{a}}^{(1)}}(x_{0})w_{\widehat{\mathbf{c}}^{(1)}}(y_{0})=1-w_{\widehat{\mathbf{a}}^{(1)}}(x_{0})w_{\widehat{\mathbf{c}}^{(1)}}(y_{0})\leq\rho.$
Using this bound inductively from the more nested sum to the first one then yields our desired bound:
$\sum_{\mathbf{a},\mathbf{c}\in\mathcal{A}^{n}}w_{\mathbf{a}}(x_{0})w_{\mathbf{c}}(y_{0})\mathbf{1}_{I(x_{\mathbf{c}})}\Big(X_{\mathbf{a}}(x_{\mathbf{c}},z_{0})\Big)\leq\rho^{m}.$
We are now ready to prove Theorem 1.3.
Proof.
Suppose (MNL). Then Proposition 6.1 applies. Thus, for any interval $I\subset\mathbb{R}$ of small enough diameter and for all $n$ large enough:
$\sum_{\mathbf{a,b,c,d}}w_{\mathbf{a}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}\chi_{[-\sigma,\sigma]}(S_{n}\tau\circ g_{\mathbf{a}}(x_{\mathbf{c}})-S_{n}\tau\circ g_{\mathbf{a}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{c}})+S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}}))$
$\sum_{\mathbf{b,d}}w_{\mathbf{b}}w_{\mathbf{d}}\sum_{\mathbf{a},\mathbf{c}}w_{\mathbf{a}}w_{\mathbf{c}}\mathbf{1}_{[X_{\mathbf{b}}(x_{\mathbf{c}},x_{\mathbf{d}})-\sigma,X_{\mathbf{b}}(x_{\mathbf{c}},x_{\mathbf{d}})+\sigma]}(X_{\mathbf{a}}(x_{\mathbf{c}},x_{\mathbf{d}}))$
$\lesssim\sum_{\mathbf{b,d}}w_{\mathbf{b}}w_{\mathbf{d}}\Big(\sigma^{\gamma}+\rho^{n}\Big)\lesssim\sigma^{\gamma}+\rho^{n},$
where we used the uniform $\delta$-Hölder regularity of the maps $X_{\mathbf{b}}(\cdot,x_{\mathbf{d}})$ to get a uniform bound from Proposition 6.1. This is the (QNL) bound. ∎
6.2. $C^{2}$ IFSs: Using UNI to prove the non-concentration
The goal of this subsection is to prove 1.2, that is, we will show that (QNL) holds under (UNI).
To this aim, we will state and prove a countable $C^{1+\alpha}$ branch version of the Tree Lemma presented in [28, Lemma 2.5.7] in the $C^{2}$ finite branch setting. This type of “tree Lemma” is useful as it allows us to avoid the use of spectral gap argument while still adapting some parts of a Dolgopyat style approach. The main difficulty of the countable compared to [28, Lemma 2.5.7] is the possible blow-ups of derivatives caused by the parabolicity or infinite number of branches, which prevents us to use same parameters as in the finite branch case. While we cannot work with the derivatives $\partial_{t}\log\left|g_{\mathbf{a}}^{\prime}(t)\right|_{|t=x}$ directly, it turns out that the simplicity of the tree Lemma allows us to conclude the necessary conclusion with the following $\alpha$-Hölder difference quotients of $\log\left|g_{\mathbf{a}}^{\prime}\right|$:
$\lambda_{\mathbf{a}}^{\alpha}\left(x,y\right)$ $\equiv\frac{X_{\mathbf{a}}(x,y)}{|x-y|^{\alpha}}=\frac{\log\left(\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|\right)-\log\left(\left|g_{\mathbf{a}}^{\prime}\left(y\right)\right|\right)}{\left|x-y\right|^{\alpha}},$
where $\mathbf{a}\in\Sigma^{N}$ for $N\in\mathbb{N}$ and $x\neq y$ with $\mathbf{a}\to x$ and $\mathbf{a}\to y$.
Recall that in our setting $\Phi$ satisfies the
  • (1)
    uniform $\kappa_{+}$-contraction:
    $\forall ab\in\Sigma^{2},x\in I_{b}:0<\kappa_{-}(ab)\leq|g_{ab}^{\prime}(x)|\leq\kappa_{+}<1.$
  • (2)
    bounded $(C,\alpha)$-distortions:
    $\mathrm{HA}\left(\Phi\right)$ $:=\underset{x,y\in I_{b}}{\sup_{ab\in\Sigma^{2}}}\left|\lambda_{ab}^{\alpha}\left(x,y\right)\right|<\infty.$
  • (3)
    Uniformly $(c_{0},\alpha)$-Non-Integrability (UNI): there exists infinitely many $N\in\mathbb{N}$ such that for all $c\in\mathcal{A}$, for any two $\mathbf{a},\mathbf{b}\in\Sigma^{N}$ with $b(\mathbf{a})=b(\mathbf{b})=c$, and for all $x,y\in I_{c}$ such that $x\neq y$, the following inequality holds:
    $\left|\lambda_{\mathbf{a}}^{\alpha}\left(x,y\right)-\lambda_{\mathbf{b}}^{\alpha}\left(x,y\right)\right|\geq c_{0}.$
Moreover, we chose the potential $\phi$ such that we have the following
  • (4)
    light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail some $\gamma>0$ we have $C_{\phi,\gamma_{0}}:=\sum_{ab\in\Sigma^{2}}\sup_{x\in I_{b}}w_{a}\left(x\right)\left|g_{a}^{\prime}\left(x\right)\right|^{-\gamma_{0}}<\infty$.
Our aim is to prove the following non-concentration property on a Cantor set defined by the $\lambda_{\mathbf{a}}^{\alpha}(x,y)$ when $\mathbf{a}$ varies:
Proposition 6.6 (Non-concentration of $\lambda_{\mathbf{a}}(x,y)$).
Let $c_{0}>0$ be the constant in the lower bound of the UNI condition, and $N\geq 0$ be sufficiently large such that it satisfies the UNI condition for $c_{0}$ and also that $A\kappa_{+}^{\alpha N}\leq\frac{c_{0}}{4}$, where $A\equiv\mathrm{HA}\left(\Phi\right)/\left(1-\kappa_{+}^{\alpha}\right)$. Let $b\in\mathcal{A}$.
For $n\geq 1,x\in I_{b},y,z\in F\cap I_{b}:y\neq z,t:F^{2}\to\mathbb{R},\sigma>0$ and the parameter $\alpha>0$, define the loss function $L_{x,y,z,t,\sigma,\alpha}\left(n\right)$
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $\equiv\underset{b(\mathbf{a})=b}{\sum_{\mathbf{a}\in\Sigma^{n}}}w_{\mathbf{a}}\left(x\right)1_{\left[t\left(x,y\right)-\sigma,t\left(x,y\right)+\sigma\right]}\left(\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)\right).$
Then: $\exists\gamma,\alpha_{N},c_{0}>0,\kappa\in\left(0,1\right):\forall t\in\mathbb{R},\sigma>0,n\geq 1,x\in I_{b},y\neq z,y,z\in I_{b}\cap F$,
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right).$
Explicitly,
$\gamma:=\min\Big\{\gamma_{0},\frac{\gamma_{0}}{2\alpha},\frac{1}{\alpha\lambda N},\frac{\alpha\lambda N}{\varepsilon_{\gamma_{0}}N\alpha^{2}},\frac{1}{2(\lambda N+C)},\frac{\varepsilon_{\mathrm{UNI}}}{4\alpha C_{\gamma^{\prime},\alpha}e^{2\alpha\gamma_{0}C}C_{\phi}}\Big\},$
where $\gamma^{\prime}=\min\{\frac{\gamma_{0}}{2\alpha},\frac{t_{\gamma_{0}}}{\alpha}\}$ and $C_{\alpha,\gamma^{\prime}}=-W(-\alpha\gamma^{\prime})/(\alpha\gamma^{\prime})>0$ for the $W$ is the Lambert W-function and where $(C,\alpha)$ are the constants of bounded distortions, $\lambda$ the Lyapunov exponent, $C^{\prime}>0$ is a universal constant, $C_{\phi}>0,0<t_{\gamma_{0}}<\gamma_{0}$ from Proposition 2.1 under light $(C_{\phi,\gamma_{0}},\gamma_{0})$ tail, and
$\varepsilon_{\mathrm{UNI}}:=\min\left\{\inf_{x\in I_{b}}w_{\mathbf{a}_{1}}\left(x\right),\inf_{x\in I_{b}}w_{\mathbf{b}_{1}}\left(x\right)\right\}>0,$
for words $\mathbf{a}_{1},\mathbf{b}_{1}\in\Sigma^{N}$ satisfying the $(c_{0},\alpha)$-UNI assumption of the IFS $\Phi$. Lastly, define the exponent $\alpha_{N}$ with these words $\mathbf{a}_{1},\mathbf{b}_{1}$:
$\alpha_{N}:=\min\Big\{\inf_{x\in[0,1]}-\frac{1}{N}\ln\Big(\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)\Big)\mid\mathbf{c}\in\{\mathbf{a}_{1},\mathbf{b}_{1}\}\Big\}.$
To prove Proposition 6.6, we need first the following important non-concentration on the fractal $F$ where UNI is used.
Lemma 6.7 (Application of UNI).
Assume the hypotheses of Proposition 6.6. Suppose $0<p<\frac{c_{0}}{4}$ is a fixed parameter and $t:F^{2}\to\mathbb{R}$. Then for every $c\in\mathcal{A}$ the following holds true: If $\mathbf{a}\in\Sigma^{N}$ and $\mathbf{b}\in\Sigma^{n-N}$ with $b\left(\mathbf{a}\right)=b\left(\mathbf{b}\right)=c,\mathbf{b}^{\prime}\mathbf{a}\in\Sigma^{n}$, then
$\forall x,y\in F\cap I_{c}:x\neq y\implies\lambda_{\mathbf{b}^{\prime}\mathbf{a}}^{\alpha}\left(x,y\right)\not\in\left[t\left(x,y\right)-p,t\left(x,y\right)+p\right]$
To prove Lemma 6.7, we first need the following Lemma that follows from the bounded distortions and uniform contraction:
Lemma 6.8.
We have
$\forall N\in\mathbb{N}:\forall\mathbf{a}\in\Sigma^{N}:\sup_{x\neq y\in I_{b(\mathbf{a})}}\left|\lambda_{\mathbf{a}}^{\alpha}\left(x,y\right)\right|\leq\frac{\mathrm{HA}\left(\Phi\right)\left(1-\kappa_{+}^{\alpha N}\right)}{1-\kappa_{+}^{\alpha}}.$
Proof.
The proof is done by induction on the variable $N$. The case $N=1$ is covered by hypothesis. Assume that for $1\leq N\leq k$ we have
$\forall N\in\mathbb{N}:\forall\mathbf{a}\in\Sigma^{N}:\sup_{x\neq y\in I_{b(\mathbf{a})}}\left|\lambda_{\mathbf{a}}^{\alpha}\left(x,y\right)\right|\leq\mathrm{HA}\left(\Phi\right)\sum_{j=0}^{N-1}\kappa_{+}^{\alpha j}.$
Next, for any such $\mathbf{c}=a\mathbf{b}\in\Sigma^{k+1}$ with $a\in\mathcal{A}$ and $\mathbf{b}\in\Sigma^{k}$ and $x\neq y\in I_{b(\mathbf{a})}$, the chain rule implies that
$\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)$ $=\lambda_{\mathbf{b}}^{\alpha}\left(x,y\right)+\frac{\log\left(\left|g_{a}^{\prime}\left(g_{\mathbf{b}}\left(x\right)\right)\right|\right)-\log\left(\left|g_{a}^{\prime}\left(g_{\mathbf{b}}\left(y\right)\right)\right|\right)}{\left|x-y\right|^{\alpha}},$
where
$\left|\frac{\log\left(\left|g_{a}^{\prime}\left(g_{\mathbf{b}}\left(x\right)\right)\right|\right)-\log\left(\left|g_{a}^{\prime}\left(g_{\mathbf{b}}\left(y\right)\right)\right|\right)}{\left|x-y\right|^{\alpha}}\right|$ $\leq\frac{\mathrm{HA}\left(\Phi\right)\left|g_{\mathbf{b}}\left(x\right)-g_{\mathbf{b}}\left(y\right)\right|^{\alpha}}{\left|x-y\right|^{\alpha}}\leq\mathrm{HA}\left(\Phi\right)\kappa_{+}^{\alpha k},$ (6.1)
by the mean value theorem. Hence, combining the induction hypothesis with (6.1) and taking supremum w.r.t. $x,y$ and $\mathbf{c}\in\Sigma^{k+1}$ shows that
$\forall\mathbf{c}\in\Sigma^{k+1}:\sup_{x\neq y\in I_{b(\mathbf{c})}}\left|\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)\right|$ $\leq\mathrm{HA}\left(\Phi\right)\sum_{j=0}^{k}\kappa_{+}^{\alpha j}=\frac{\mathrm{HA}\left(\Phi\right)\left(1-\kappa_{+}^{\alpha\left(k+1\right)}\right)}{1-\kappa_{+}^{\alpha}}.$
We can now prove Lemma 6.7:
Proof of Lemma 6.7.
Let $c\in\mathcal{A}$. By hypothesis we know that there exists two words $\mathbf{a},\mathbf{b}\in\Sigma^{N}$ such that $b(\mathbf{a})=b(\mathbf{b})=c$ and
$\forall x\neq y\in F\cap I_{c}:\left|\lambda_{\mathbf{a}}^{\alpha}\left(x,y\right)-\lambda_{\mathbf{b}}^{\alpha}\left(x,y\right)\right|\geq c_{0}.$ (6.2)
Let $x,y\in F\cap I_{c}$ be fixed and $t:F^{2}\to\mathbb{R}$ a map. One of the words, call it $\mathbf{c}$, satisfies
$\left|\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|\geq\frac{c_{0}}{2}.$
Otherwise we would arrive at a contraction in (6.2) after adding and subtracting $t\left(x,y\right)$ and using the triangle inequality. We are ultimately looking to show that the inequality
$\left|\lambda_{\mathbf{b}^{\prime}\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|\geq\frac{c_{0}}{4}.$
holds for any prefix $\mathbf{b}\in\Sigma^{n-N}$. Let $\mathbf{b}\in\Sigma^{n-N}$ be fixed, but arbitrary. By the chain rule we obtain that
$\lambda_{\mathbf{b}^{\prime}\mathbf{c}}^{\alpha}\left(x,y\right)$ $=\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)+\frac{\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(x\right)\right)\right|\right)-\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(y\right)\right)\right|\right)}{\left|x-y\right|^{\alpha}},$
and by applying exactly the same kind of argument as in the proof of Lemma 6.8, we get
$\left|\frac{\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(x\right)\right)\right|\right)-\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(y\right)\right)\right|\right)}{\left|x-y\right|^{\alpha}}\right|$ $\leq\frac{A\left|g_{\mathbf{c}}\left(x\right)-g_{\mathbf{c}}\left(y\right)\right|^{\alpha}}{\left|x-y\right|^{\alpha}}=A\left|g_{\mathbf{c}}^{\prime}\left(z\right)\right|^{\alpha}\leq A\kappa_{+}^{N\alpha}\leq\frac{c_{0}}{4},$
for some $z\in\left(x,y\right)\subset I_{c}$. Consequently, we obtain:
$\left|\lambda_{\mathbf{b}^{\prime}\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|$ $\geq\left|\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|-\left|\frac{\log\left(g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(x\right)\right)\right)-\log\left(g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{c}}\left(y\right)\right)\right)}{\left(x-y\right)^{\alpha}}\right|$
$\left|\lambda_{\mathbf{b}^{\prime}\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|$ $\geq\frac{c_{0}}{2}-A\kappa_{+}^{N\alpha}\geq\frac{c_{0}}{4}\implies\lambda_{\mathbf{b}^{\prime}\mathbf{c}}^{\alpha}\left(x,y\right)\not\in\left[t\left(x,y\right)-p,t\left(x,y\right)+p\right],$
for the chosen $x,y\in F\cap I_{c}$, with $\mathbf{b}\in\Sigma^{n-N}$ s.t. the word $\mathbf{b}^{\prime}\mathbf{c}$ is admissible, $n>N$, and $p\in\left(0,\frac{c_{0}}{4}\right)$.
With Lemma 6.7, we can finish the proof of Proposition 6.6 in the following:
Proof of Proposition 6.6.
The proof is done by induction on the variable $n$. Write $A\equiv\mathrm{HA}\left(\Phi\right)$. Let us first note that the loss function $L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ has the following autosimilarity property:
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $=\sum_{\mathbf{a}\in\Sigma^{k}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)L_{x\left(\mathbf{a}\right),y\left(\mathbf{a}\right),z\left(\mathbf{a}\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right),\sigma\left(\mathbf{a},y,z,\alpha\right),\alpha}\left(n-k\right),$
for any $1\leq k\leq n-1$. This can be seen as follows: Fix $1\leq k\leq n-1$ and partition a word $\mathbf{c}\in\Sigma^{n}$ into a suffix $\mathbf{a}\in\Sigma^{k}$ and a prefix $\mathbf{b}\in\Sigma^{n-k}$, so that $\mathbf{c}=\mathbf{b}^{\prime}\mathbf{a}$.
Next the inclusion condition
$\lambda_{\mathbf{b}^{\prime}\mathbf{a}}^{\alpha}\left(y,z\right)\in\left[t\left(y,z\right)-\sigma,t\left(y,z\right)+\sigma\right],$
for $y,z$ such that $\mathbf{b}^{\prime}\mathbf{a}\to y,z$ can be rewritten “in terms of a fixed $\mathbf{a}\in\Sigma^{k}$” with the use of the mean value theorem as follows:
$\lambda_{\mathbf{b}^{\prime}\mathbf{a}}^{\alpha}\left(y,z\right)$ $=\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)+\frac{\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{a}}\left(y\right)\right)\right|\right)-\log\left(\left|g_{\mathbf{b}}^{\prime}\left(g_{\mathbf{a}}\left(z\right)\right)\right|\right)}{\left|y-z\right|^{\alpha}}$
$=\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)+\lambda_{\mathbf{b}}^{\alpha}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\frac{\left|g_{\mathbf{a}}\left(y\right)-g_{\mathbf{a}}\left(z\right)\right|^{\alpha}}{\left|y-z\right|^{\alpha}}$
$=\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)+\lambda_{\mathbf{b}}^{\alpha}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha},$
for some $\eta\left(\mathbf{a},y,z\right)\in\left(z,y\right)\subset I_{b(\mathbf{a})}$. Hence,
$t\left(y,z\right)-\sigma$ $\leq\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)+\lambda_{\mathbf{b}}^{\alpha}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}\leq t\left(y,z\right)+\sigma\Longleftrightarrow$
$\frac{t\left(y,z\right)-\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)-\sigma}{\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}}$ $\leq\lambda_{\mathbf{b}}^{\alpha}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\leq\frac{t\left(y,z\right)-\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)+\sigma}{\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}}.$
Hence, we can set
$t\left(\mathbf{a},\alpha,\Phi\right)\left(y,z\right)$ $\equiv\frac{t\left(y,z\right)-\lambda_{\mathbf{a}}^{\alpha}\left(y,z\right)}{\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}},\ \sigma\left(\mathbf{a},y,z,\alpha\right)\equiv\frac{\sigma}{\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}},$
$x\left(\mathbf{a}\right)$ $\equiv g_{\mathbf{a}}\left(x\right),\ y\left(\mathbf{a}\right)\equiv g_{\mathbf{a}}\left(y\right),\ z\left(\mathbf{a}\right)\equiv g_{\mathbf{a}}\left(z\right).$
Thus,
$\lambda_{\mathbf{b}}^{\alpha}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\in\left[t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right)-\sigma\left(\mathbf{a},y,z,\alpha\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right)+\sigma\left(\mathbf{a},y,z,\alpha\right)\right],$
and consequently
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)=\sum_{\mathbf{c}\in\Sigma^{n}\atop\mathbf{c}\to x,y,z}w_{\mathbf{c}}\left(x\right)1_{\left[t\left(y,z\right)-\sigma,t\left(y,z\right)+\sigma\right]}\left(\lambda_{\mathbf{c}}^{\alpha}\left(y,z\right)\right)$
$=\sum_{\mathbf{a}\in\Sigma^{k}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\sum_{\mathbf{b}\in\Sigma^{n-k}\atop\mathbf{b}\to g_{\mathbf{a}}(x)}w_{\mathbf{b}}\left(g_{\mathbf{a}}\left(x\right)\right)1_{\left[t\left(\mathbf{a},\alpha,\Phi\right)\left(y,z\right)-\sigma\left(\mathbf{a},y,z,\alpha\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(y,z\right)+\sigma\left(\mathbf{a},y,z,\alpha\right)\right]}\left(\lambda_{\mathbf{b}}\left(g_{\mathbf{a}}\left(y\right),g_{\mathbf{a}}\left(z\right)\right)\right)$
$=\sum_{\mathbf{a}\in\Sigma^{k}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)L_{x\left(\mathbf{a}\right),y\left(\mathbf{a}\right),z\left(\mathbf{a}\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right),\sigma\left(\mathbf{a},y,z,\alpha\right),\alpha}\left(n-k\right).$
Next, let us consider the different cases of the induction proof:
Case 1. If $n\leq N$, then it follows that
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)\leq\sum_{\mathbf{a}\in\Sigma^{n}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)=1,$
and as $n-N\leq 0$ and $e^{-\alpha_{N}}<1$, we have that $\exp\left(-\alpha_{N}\left(n-N\right)\right)\geq 1$ and so
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right).$
We will then use Lemma 6.7 to cut unnecessary branches of the tree.
Case 2. If $N+1\leq n$, there are two subcases to consider.
Case 2(a). First, if $\sigma\geq\frac{c_{0}}{4}$, then $\frac{4}{c_{0}}\sigma\geq\frac{1}{\sigma}\sigma=1$. As the weights $w_{\mathbf{a}}\left(x\right)$ sum to one w.r.t. $\mathbf{a}\in\Sigma^{N}$, we get the trivial bound
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)\leq 1\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right).$
Case 2(b). Assume that $\sigma<\frac{c_{0}}{4}$. Due to proof technical reasons we will proceed iteratively by showing that the induction hypothesis applies for all $n$ such that $pN+1\leq n\leq\left(p+1\right)N$ for $p=1,2,\dots$. This in turn will conclude the claim for all $n\in\mathbb{N}$.
So suppose first that $N+1\leq n\leq 2N$. The autosimilar property of the loss function yields for $k=N$ that
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $=\sum_{\mathbf{a}\in\Sigma^{N}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)L_{x\left(\mathbf{a}\right),y\left(\mathbf{a}\right),z\left(\mathbf{a}\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right),\sigma\left(\mathbf{a},y,z,\alpha\right),\alpha}\left(n-N\right).$
We will next use Lemma 6.7 to cut unnecessary branches of the tree. Let $\mathbf{c}$ be such a suffix guaranteed by the Lemma 6.7 so that
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $=\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)L_{x\left(\mathbf{a}\right),y\left(\mathbf{a}\right),z\left(\mathbf{a}\right),t\left(\mathbf{a},\alpha,\Phi\right)\left(.,.\right),\sigma\left(\mathbf{a},y,z,\alpha\right),\alpha}\left(n-N\right).$ (6.3)
Since $n-N\leq N$, the previously written steps allow us to apply the induction hypothesis
$L_{x,y,z,t,\sigma,\alpha}\left(s\right)$ $\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(s-N\right)\right),$
to the value s=nNs=n-N, so that
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $\leq\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\left(\left(\frac{4\sigma\left(\mathbf{a},y,z,\alpha\right)}{c_{0}}\right)^{\gamma}+\exp(-\alpha_{N}(n-2N))\right).$ (6.4)
Since $\sigma\left(\mathbf{a},y,z,\alpha\right)=\sigma/\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{\alpha}$, we arrive to
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $\leq\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\left(\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}\left|g_{\mathbf{a}}^{\prime}\left(\eta\left(\mathbf{a},y,z\right)\right)\right|^{-\alpha\gamma}+\exp(-\alpha_{N}(n-2N))\right).$ (6.5)
where the sum can in the RHS of (6.5) can be bounded by the bounded $(C,\alpha)$-distortions by
$\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\left(\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}e^{\alpha\gamma C}\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}+\exp\left(-\alpha_{N}\left(n-2N\right)\right)\right).$
For any $0<\gamma\leq\gamma_{0}/\alpha$, denote
$\mathcal{Q}_{N,x}\left(\gamma\right)\equiv e^{\alpha\gamma_{0}C}\sum_{\mathbf{a}\in\Sigma^{N}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}=e^{\alpha\gamma\lambda N+\alpha\gamma_{0}C}\sum_{\mathbf{a}\in\Sigma^{N}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)(e^{\lambda N}\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|)^{-\alpha\gamma}.$
Next, an application of Taylor’s theorem shows that
$\forall\gamma^{\prime}\in\left(0,1\right):\exists C_{\gamma^{\prime}}>0:\forall x\in\left[1,\infty\right):\ln\left(x\right)\leq C_{\alpha,\gamma^{\prime}}x^{\alpha\gamma^{\prime}},$
where $C_{\alpha,\gamma^{\prime}}=-W(-\alpha\gamma^{\prime})/(\alpha\gamma^{\prime})>0$ and $W$ is the Lambert W-function. In turn, this shows that $\ln\left(\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|\right)\leq C_{\alpha,\gamma^{\prime}}\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma^{\prime}}$ so that
$w_{\mathbf{a}}\left(x\right)\left(-\alpha\ln\left(\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|\right)\right)\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}\leq\alpha C_{\gamma^{\prime},\alpha}w_{\mathbf{a}}\left(x\right)\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha(\gamma+\gamma^{\prime})}.$
Hence, as long as $\max\{\gamma,\gamma^{\prime}\}\leq\frac{\gamma_{0}}{2\alpha}$, then $\alpha(\gamma+\gamma^{\prime})\leq\gamma_{0}$, which shows that $\mathcal{Q}_{N,x}\left(\gamma\right)$ is differentiable w.r.t. $\gamma$, and we can bound the derivative
$\mathcal{Q}_{N,x}^{\prime}(\gamma)=e^{\alpha\gamma_{0}C}\sum_{\mathbf{a}\rightarrow x}w_{\mathbf{a}}(x)(-\alpha\ln|g_{\mathbf{a}}^{\prime}(x)|)|g_{\mathbf{a}}^{\prime}(x)|^{-\alpha\gamma}\leq\alpha C_{\gamma^{\prime},\alpha}e^{\alpha\gamma_{0}C}\mathcal{Q}_{N,x}(\gamma+\gamma^{\prime}).$
Now, by Proposition 2.1, as long as $\max\{\gamma,\gamma^{\prime}\}\leq t_{\gamma_{0}}/\alpha$ we have
$\mathcal{Q}_{N,x}(\gamma+\gamma^{\prime})$ $=e^{\alpha(\gamma+\gamma^{\prime})\lambda N+\alpha\gamma_{0}C}\sum_{\mathbf{a}\in\Sigma^{N}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)(e^{\lambda N}\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|)^{-\alpha(\gamma+\gamma^{\prime})}$
$\leq e^{\alpha(\gamma+\gamma^{\prime})\lambda N}2C_{\phi}e^{\varepsilon_{\gamma_{0}}N\alpha^{2}(\gamma+\gamma^{\prime})^{2}}$
$\leq 2C_{\phi}e^{\alpha\gamma_{0}C}$
as long as $\max\{\gamma,\gamma^{\prime}\}\leq\min\{\frac{1}{\alpha\lambda N},\frac{\alpha\lambda N}{\varepsilon_{\gamma_{0}}N\alpha^{2}}\}$ since then $\alpha(\gamma+\gamma^{\prime})\lambda N+\varepsilon_{\gamma_{0}}N\alpha^{2}(\gamma+\gamma^{\prime})^{2}\leq 1$. Thus by the mean value theorem
$\mathcal{Q}_{N,x}(\gamma)\leq 1+\gamma\mathcal{Q}_{N,x}^{\prime}(\gamma)\leq 1+2\gamma\alpha C_{\gamma^{\prime},\alpha}e^{2\alpha\gamma_{0}C}C_{\phi}.$
Now recall that $\mathbf{c}$ in Lemma 6.7 is one of the words $\mathbf{a}_{1},\mathbf{b}_{1}\in\Sigma^{N}$ in the $(c_{0},\alpha)$-UNI assumption of the IFS $\Phi$. Since
$\varepsilon_{\mathrm{UNI}}\leq w_{\mathbf{c}}(x)$
and with the choices $\gamma^{\prime}:=\min\{\frac{\gamma_{0}}{2\alpha},\frac{t_{\gamma_{0}}}{\alpha}\}$ and
$\gamma:=\min\Big\{\gamma_{0},\frac{\gamma_{0}}{2\alpha},\frac{1}{\alpha\lambda N},\frac{\alpha\lambda N}{\varepsilon_{\gamma_{0}}N\alpha^{2}},\frac{1}{2(\lambda N+C)},\frac{\varepsilon_{\mathrm{UNI}}}{4\alpha C_{\gamma^{\prime},\alpha}e^{2\alpha\gamma_{0}C}C_{\phi}}\Big\}.$
we have that
$\mathcal{Q}_{N,x}\left(\gamma\right)\leq 1+w_{\mathbf{c}}(x)$
so by defining
$\mathcal{R}\left(\gamma,N,\mathbf{c}\right)\equiv\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}e^{\alpha\gamma C}w_{\mathbf{a}}\left(x\right)\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}$
we have
$\mathcal{R}\left(\gamma,N,\mathbf{c}\right)<1+w_{\mathbf{c}}(x)-e^{\alpha\gamma C}w_{\mathbf{c}}\left(x\right)\left|g_{\mathbf{c}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}\leq 1+w_{\mathbf{c}}\left(x\right)\left(1-e^{\alpha\gamma C}\left|g_{\mathbf{c}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}\right).$ (6.6)
Since $\left|g_{\mathbf{c}}^{\prime}\left(x\right)\right|<1$, it follows that we have $1-e^{\alpha\gamma C}\left|g_{\mathbf{c}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}<0,$ so that $w_{\mathbf{c}}\left(x\right)\left(1-\left|g_{\mathbf{c}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}\right)<0$ so for our choice of $\gamma$ we indeed have have:
$\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}e^{\alpha\gamma C}w_{\mathbf{a}}\left(x\right)\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}$ $\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}.$ (6.7)
Lastly recall that
$\alpha_{N}$ $\leq\alpha_{N}\left(\mathbf{c}\right)\equiv\inf_{x\in[0,1]}-\frac{1}{N}\ln\left(\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)\right),$
so that,
$\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x}w_{\mathbf{a}}\left(x\right)\leq\exp\left(-\alpha_{N}N\right),$
and so
$\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\left(\exp(-\alpha_{N}(n-2N))\right)$ $\leq\exp\left(-\alpha_{N}\left(n-N\right)\right),$
which together with (6.7) shows that
$\sum_{\mathbf{a}\in\Sigma^{N}:\mathbf{a}\neq\mathbf{c}\atop\mathbf{a}\to x,y,z}w_{\mathbf{a}}\left(x\right)\left(\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}e^{\alpha\gamma C}\left|g_{\mathbf{a}}^{\prime}\left(x\right)\right|^{-\alpha\gamma}+\exp\left(-\alpha_{N}\left(n-2N\right)\right)\right)$ $\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right).$
Thus
$L_{x,y,z,t,\sigma,\alpha}\left(n\right)$ $\leq\left(\frac{4\sigma}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right),$
for all $n\in\mathbb{N}:N+1\leq n\leq 2N$.
Next, let us note that the precise part where we had to use the assumption $N+1\leq n\leq 2N$ was when we used the induction hypothesis to move from the equation (6.3) to the inequality (6.4). Crucial to this application was that $n-N\leq B$ for some $B$ such that the induction hypothesis applies for all $k=1,2,\dots,B$. Beyond this point, the assumption $N+1\leq n\leq 2N$ played no part in rest of the proof. Therefore, now that the induction hypothesis applies for all $k=1,\dots,2N$, we can redo the proof of the part b of the case 2 for $n$ s.t. $pN+1\leq n\leq\left(p+1\right)N$ for $p=2,3,\dots$. Hence the induction hypothesis applies for all $n\in\mathbb{N}$, which proves the claim. ∎
We now conclude this subsection by proving Theorem 1.2, using the non-concentration estimate given by Proposition 6.6.
Proof.
Suppose (UNI). Then Proposition 6.6 holds. Let us now reduce (QNL) to this nonconcentration bound. Let $I\subset\mathbb{R}$ be some interval. Consider the sum:
$\sum_{\mathbf{b,c,d,e}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-\sigma,\sigma]}(S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{d}})-S_{n}\tau\circ g_{\mathbf{b}}(x_{\mathbf{e}})-S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{d}})+S_{n}\tau\circ g_{\mathbf{c}}(x_{\mathbf{e}}))$
$\sum_{\mathbf{b,c,d,e}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-\sigma,\sigma]}\Big(\big(\lambda_{\mathbf{b}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})-\lambda_{\mathbf{c}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})\big)\big|x_{\mathbf{d}}-x_{\mathbf{e}}\big|^{\alpha}\Big)$
Define a neighbourhood of the diagonal in the following way:
$\mathcal{D}_{n}(\sigma^{1/2}):=\{(\mathbf{d},\mathbf{e})\in\Sigma^{2n},\ |x_{\mathbf{d}}-x_{\mathbf{e}}|\leq\sigma^{1/2}\}.$
Notice that, for every $(x,y)\in I_{\mathbf{d}}\times I_{\mathbf{e}}$ with $(\mathbf{d},\mathbf{e})\in\mathcal{D}_{n}(\sigma^{1/2})$, we have
$|x-y|\leq|x-x_{\mathbf{d}}|+|y-x_{\mathbf{e}}|+|x_{\mathbf{d}}-x_{\mathbf{e}}|\lesssim 2\kappa_{+}^{n}+\sigma^{1/2}.$
Summing characteristic functions then gives:
$\sum_{(\mathbf{d},\mathbf{e})\in\mathcal{D}_{n}(\sigma^{1/2})}\chi_{I_{\mathbf{d}}\times I_{\mathbf{e}}}\leq\chi_{\{(x,y),\ |x-y|\leq 2\kappa_{+}^{n}+\sigma^{1/2}\}}.$
Upper regularity of the measure $\mu$ and Gibbs estimates then yields:
$\sum_{(\mathbf{d},\mathbf{e})\in\mathcal{D}_{n}(\sigma^{1/2})}w_{\mathbf{d}}w_{\mathbf{e}}\leq\iint_{I\times I}\sum_{(\mathbf{d},\mathbf{e})\in\mathcal{D}_{n}(\sigma^{1/2})}\chi_{I_{\mathbf{d}}\times I_{\mathbf{e}}}d\mu^{2}(x,y)$
$\leq\mu^{\otimes 2}\big((x,y)\in I^{2},\ |x-y|\leq 2\kappa_{+}^{n}+\sigma^{1/2}\big)=\int_{I}\mu(B(x,2\kappa_{+}^{n}+\sigma^{1/2})d\mu(x)\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}.$
We can then conclude by throwing away the diagonal part, and then bounding the off-diagonal part. For the diagonal, we find:
$\sum_{\left(\mathbf{d},\mathbf{e}\right)\in\mathcal{D}_{n}\left(\sigma^{1/2}\right)}w_{\mathbf{d}}w_{\mathbf{e}}\sum_{\mathbf{b},\mathbf{c}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{[-\sigma,\sigma]}\Big(\big(\lambda_{\mathbf{b}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})-\lambda_{\mathbf{c}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})\big)\big|x_{\mathbf{d}}-x_{\mathbf{e}}\big|^{\alpha}\Big)\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}.$
It follows that we get a bound
$\sum_{\mathbf{b,c,d,e}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-\sigma,\sigma]}\Big(\big(\lambda_{\mathbf{b}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})-\lambda_{\mathbf{c}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})\big)\big|x_{\mathbf{d}}-x_{\mathbf{e}}\big|^{\alpha}\Big)$
$\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}+\sum_{\left(\mathbf{d},\mathbf{e}\right)\notin\mathcal{D}_{n}\left(\sigma^{1/2}\right)}w_{\mathbf{d}}w_{\mathbf{e}}\sum_{\mathbf{b},\mathbf{c}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{[-\sigma,\sigma]}\Big(\big(\lambda_{\mathbf{b}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})-\lambda_{\mathbf{c}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})\big)\big|x_{\mathbf{d}}-x_{\mathbf{e}}\big|^{\alpha}\Big)$
$\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}+\sum_{\mathbf{b,c,d,e}}w_{\mathbf{b}}w_{\mathbf{c}}w_{\mathbf{d}}w_{\mathbf{e}}\chi_{[-\sigma^{1/2},\sigma^{1/2}]}\big(\lambda_{\mathbf{b}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})-\lambda_{\mathbf{c}}^{\alpha}(x_{\mathbf{d}},x_{\mathbf{e}})\big)$
To bound this second sum, we write
$\sum_{\mathbf{b},\mathbf{c}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{\left[-\sigma^{1/2},\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right)=$
$\sum_{\mathbf{c}}w_{\mathbf{c}}\left(x\right)\sum_{\mathbf{b}}w_{\mathbf{b}}\chi_{\left[\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\sigma^{1/2},\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)+\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right).$
Notice then that by Proposition 6.6 we obtain for some $\gamma>0,c_{0}>0$:
$\sum_{\mathbf{b}}w_{\mathbf{b}}\chi_{\left[\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\sigma^{1/2},\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)+\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right)$
$\leq\left(\frac{4\sigma^{1/2}}{c_{0}}\right)^{\gamma}+\exp\left(-\alpha_{N}\left(n-N\right)\right),$
where $N\in\mathbb{N}$ is a large enough constant such that it satisfies the UNI condition for the IFS $\Phi$. Thus,
$\sum_{\mathbf{c}}w_{\mathbf{c}}\sum_{\mathbf{b}}w_{\mathbf{b}}\chi_{\left[\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\sigma^{1/2},\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)+\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right)$
$\leq\left(\frac{12\sigma^{1/2}}{c_{0}}\right)^{\gamma}+e^{{-\alpha_{N}\left(n-N\right)}}.$
Thus we can bound
$\sum_{\mathbf{d},\mathbf{e}}w_{\mathbf{d}}w_{\mathbf{e}}\sum_{\mathbf{b},\mathbf{c}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{\left[-3\sigma^{1/2},3\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right)$
$\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}+\sum_{\left(\mathbf{d},\mathbf{e}\right)\not\in\mathcal{D}_{n}\left(\sigma^{1/2}\right)}w_{\mathbf{d}}w_{\mathbf{e}}\sum_{\mathbf{b},\mathbf{c}}w_{\mathbf{b}}w_{\mathbf{c}}\chi_{\left[-\sigma^{1/2},\sigma^{1/2}\right]}\left(\lambda_{\mathbf{b}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)-\lambda_{\mathbf{c}}\left(x_{\mathbf{d}},x_{\mathbf{e}}\right)\right)$
$\lesssim\kappa_{+}^{s_{\mu}n}+\sigma^{s_{\mu}/2}+\sum_{\left(\mathbf{d},\mathbf{e}\right)\not\in\mathcal{D}_{n}\left(\sigma^{1/2}\right)}w_{\mathbf{d}}w_{\mathbf{e}}\ (\sigma^{\gamma/2}+e^{-\alpha_{N}(n-N)})\lesssim\sigma^{\min(s_{\mu},\gamma)/2}+\max(e^{-\alpha_{N}},\kappa_{+}^{s_{\mu}})^{n}.$
we thus proved (QLN) with constants $\Theta:=\min(s_{\mu},\gamma)/2>0$ and $\rho:=\max(e^{-\alpha_{N}},\kappa_{+}^{s_{\mu}})\in(0,1)$. ∎
7. Proof of the Measure Non-Linearity condition for $(\Phi_{\mu},-\delta_{\mu}\tau_{\mu})$
This section is devoted to the proof of Theorem 1.4. Recall some of the notations: we have fixed $C\geq 1$ and $\varepsilon>0$ and denoted
$\mathcal{M}_{C,\varepsilon}=\{\mu\in\mathcal{P}([-1,1]):\forall\text{ interval }I\text{ with }|I|\leq\tfrac{1}{100}\text{ we have }\mu(I)\leq C|I|^{\varepsilon}\}$
Given $\mu\in\mathcal{M}_{C,\varepsilon}$, define the branches of an IFS $\Phi_{\mu}$ by
$g_{\mu,1}(x):=\kappa\int_{0}^{x}\exp(\mu(-\infty,t])\,dt+b_{1},\quad x\in[0,1],$
and $g_{\mu,0}(x)=\kappa_{0}x+b_{0}$, where $0<\kappa_{0}\leq\kappa<1/10$, $b_{0},b_{1}\in\mathbb{R}$ is chosen such that the IFS $\Phi_{\mu}=\{g_{0,\mu},g_{1,\mu}\}$ satisfies the strong separation condition. Let $T_{\mu}$ be the associated expanding map on the attractor $F_{\mu}$ to $\Phi_{\mu}$, and write the geometric potential $\tau_{\mu}(x):=-\log|T_{\mu}^{\prime}(x)|$. Note by the Definition of each $\mu\in\mathcal{M}_{C,\varepsilon}$, the IFS $\Phi_{\mu}$ is $C^{1}$ and $\tau_{\mu}$ is $\alpha$-Hölder map for any $\alpha\leq\varepsilon$. Finally, we denoted the $\mu$-controlled transfer operators
$\mathcal{L}_{\mu}f(x):=\sum_{i}|g_{\mu,i}^{\prime}(x)|^{\delta_{\mu}}f(g_{\mu,i}(x)),$
where $\delta_{\mu}\in(0,1)$ is chosen as the unique number such that
$\mathcal{L}_{\mu}^{*}(\mu)(\mathbb{R})=1,$
which exists by the intermediate value theorem as $\mu$ is a probability measure. Then this gives us an operator $\Psi:\mathcal{M}_{C,\varepsilon}\to\mathcal{M}_{C,\varepsilon}$ defined by
$\Psi(\mu):=\mathcal{L}_{\mu}^{*}(\mu),\quad\mu\in\mathcal{M}_{C,\varepsilon}.$
Now the strategy to prove Theorem 1.4 is to construct a fixed point $\mu$ for the operator $\Psi$, that is, a measure satisfying
$\mathcal{L}_{\mu}^{*}(\mu)=\mu.$
Therefore, the pressure $P(\phi^{\mu})=0$ for $\phi^{\mu}:=\delta_{\mu}\tau_{\mu}$, which implies, as $\Phi_{\mu}$ is $C^{1+\alpha}$, that $\delta^{\mu}=\dim_{\mathrm{H}}(\mathrm{supp}\,\mu)$ and that $\mu$ is $\delta^{\mu}$ Ahlfors-David regular. This then allows us to have a convenient form for the geometric potential $\tau_{\mu}$: for some $c_{1}\in\mathbb{R}$, we have
$\log g_{1}^{\prime}(x)=c_{1}+\mu(-\infty,x)\in C^{\delta_{\mu}}.$
We can find this fixed point $\mu$ by adapting Schauder’s fixed point theorem to the operator $\Psi:\mathcal{M}_{C,\varepsilon}\to\mathcal{M}_{C,\varepsilon}$, which is possible by the following Lemma:
Lemma 7.1.
  • (0)
    As a map from $\mathcal{M}_{C,\varepsilon}$, the maps $\mu\mapsto g_{\mu,i}^{\prime}$ and $\mu\mapsto g_{\mu,i}$ are continuous
  • (1)
    The choice of $\delta^{\mu}$ is well-defined and continuous as a map of $\mu$, and that $\delta^{\mu}\geq 3\varepsilon$ for all $\mu\in\mathcal{M}_{C,\varepsilon}$
  • (2)
    $\Psi$ is well-defined: $\Psi(\mathcal{M}_{C,\varepsilon})\subset\mathcal{M}_{C,\varepsilon}$
  • (3)
    $\Psi$ is continuous on $\mathcal{M}_{C,\varepsilon}$ in the weak-$\ast$ topology
  • (4)
    $\mathcal{M}_{C,\varepsilon}$ is closed and convex
We postpone the proof of these properties the end of the section. Now, the fixed point property then allows us to prove Theorem 1.4 by following Lemma:
Lemma 7.2 (Measure Non-Linearity).
The IFS $\Phi_{\mu}$ and the fixed point $\mu$ satisfies MNL condition: There exists $C_{\text{(MNL)}}$ such that for all $N\geq 1$, there exists $\widehat{\mathbf{a}},\widehat{\mathbf{b}}\in\mathcal{A}^{N}$ such that
$\forall x_{0}\in F,\ \forall I\subset\mathbb{R},\ (X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0}))_{*}\mu(I)\leq C_{\text{(MNL)}}|I|.$
Proof.
Since $\log g_{1}^{\prime}(x)=c_{1}+\mu(-\infty,x)$ for some constant $c_{1}\in\mathbb{R}$, and since $\log g_{0}^{\prime}=c_{0}$, we know that for every $N$ there exists $\mathbf{a},\mathbf{b}\in\mathcal{A}^{N}$ such that
$S_{N}\tau_{\mu}\circ g_{\mathbf{a}}-S_{N}\tau_{\mu}\circ g_{\mathbf{b}}=F_{\mu}(x)+c_{0}.$
Indeed, for example words $\mathbf{a}:=0^{N}1$ and $\mathbf{b}:=0^{N-1}1$ gives us such property. The general fact that $(F_{\mu})_{*}\mu$ is equal to the Lebesgue measure on $[0,1]$ since $\mu$ doesn’t have atoms concludes our proof. ∎
So we are left with proving Lemma 7.1.
Proof of Lemma 7.1.
Property (0). Since any $\mu\in\mathcal{M}_{C,\varepsilon}$ has no atoms, we know that $\mu_{n}(-\infty,t]\to\mu(-\infty,t]$ if $\mu_{n}\to\mu$ in $\mathcal{M}_{C,\varepsilon}$. Thus by the dominated convergence theorem $g_{\mu_{n},i}(x)\to g_{\mu,i}(x)$. Similarly $g_{\mu_{n},i}^{\prime}$ is continuous as the formula involves also just applying $\mu_{n}(-\infty,t]\to\mu(-\infty,t]$ and dominated convergence theorem. In fact, convergence for the cumulative distribution function is uniform since $\mu$ has no atoms (see for example Exercise 5.2 in [38]).
Property (1). Write $\mathcal{A}=\{0,1\}$ and define
$\zeta_{\mu}(\alpha):=\int_{\mathbb{R}}\sum_{a\in\mathcal{A}}|g_{\mu,a}^{\prime}(x)|^{\alpha}\,d\mu(x)=\sum_{a\in\mathcal{A}}\int_{\mathbb{R}}|g_{\mu,a}^{\prime}(x)|^{\alpha}\,d\mu(x).$
Then $\zeta_{\mu}$ is continuous, decreasing and satisfies $\zeta_{\mu}(0)=2$ and
$\zeta_{\mu}(1)=\sum_{a\in\mathcal{A}}\int_{\mathbb{R}}|g_{\mu,a}^{\prime}(x)|\,d\mu(x)\leq\frac{6}{10}<1$
by the choice of $\kappa<1/10$. Thus there indeed exists unique $\delta^{\mu}$ such that $\zeta_{\mu}(\delta^{\mu})=1$.
Let us now show that $\mu\mapsto\delta^{\mu}$ is continuous map from $\mathcal{M}_{C,\varepsilon}$. Suppose $\mu_{n}\to\mu$. We have for any $n\in\mathbb{N}$ by the choices of $\delta^{\mu}$ we have $\zeta_{\mu_{n}}(\delta^{\mu_{n}})=1=\zeta_{\mu}(\delta^{\mu})$. Moreover, we know that $\zeta_{\mu_{n}}\to\zeta_{\mu}$ uniformly since $\mu$ has no atoms. Thus
$|\zeta_{\mu}(\delta^{\mu_{n}})-1|\leq|\zeta_{\mu_{n}}(\delta^{\mu_{n}})-1|+|\zeta_{\mu}(\delta^{\mu_{n}})-\zeta_{\mu_{n}}(\delta^{\mu_{n}})|\leq\|\zeta_{\mu}-\zeta_{\mu_{n}}\|_{\infty}$
so $|\delta^{\mu_{n}}-\delta^{\mu}|\to 0$ as $n\to\infty$ as $\zeta_{\mu}^{\prime}(\alpha)=\sum_{a}\int(\log|g_{\mu,a}^{\prime}|)|g_{\mu,a}^{\prime}|^{\alpha}\,d\mu$ giving $\zeta^{\prime}\in[-C,-c]$ for some constants $C,c>0$. Indeed, if $f\in C^{1}(\mathbb{R},\mathbb{R})$ be such that $f^{\prime}\leq-c$, then for any $h>0$ we have by Taylor’s approximation $f(x+h)\leq f(x)-c_{0}h$, so that if $f(x_{0})\leq\varepsilon$, we know that the zero $z_{0}$ of $f$ satisfies $z_{0}\in[x_{0},x_{0}+\varepsilon/c_{0}]$. Indeed, $f(x_{0}+\varepsilon/c_{0})\leq 0$ by the previous Taylor estimate and so by intermediate value theorem the zero is in the interval. Thus $\delta^{\mu}$ is continuous in $\mu$.
Finally, we notice that
$\zeta_{\mu}(3\varepsilon)=\int_{\mathbb{R}}\sum_{a\in\mathcal{A}}|g_{\mu,a}^{\prime}(x)|^{3\varepsilon}\,d\mu(x)\geq 3/2$
if we assume $\varepsilon<1/25$ as then $|g_{\mu,a}^{\prime}(x)|^{3\varepsilon}\geq 10^{-3\varepsilon}\geq 3/4$, so $\delta^{\mu}\geq 3\varepsilon$.
Property (2). We need to check $\Psi(\mathcal{M}_{C,\varepsilon})\subset\mathcal{M}_{C,\varepsilon}$ so fix $\mu\in\mathcal{M}_{C,\varepsilon}$. We know that $\mu$ is a probability measure as $\mu=\mathcal{L}_{\mu}^{*}(\mu)$ and $\mathcal{L}_{\mu}^{*}(\mu)(\mathbb{R})=1$ by the choice of $\delta^{\mu}$. Fix $h$ supported on $(1,\infty)$, which then gives
$\int h\,d\mathcal{L}_{\mu}^{*}(\mu)=\int\mathcal{L}_{\mu}(h)\,d\mu=0$
as the support of $\mathcal{L}_{\mu}(h)$ is also in $(1,\infty)$ as for all $x\in(-1,1)$ we have $g_{\mu,i}(x)\in(-1,1)$ and so $\mathcal{L}_{\mu}(h)(x)=0$. Hence we know that $\mathcal{L}_{\mu}^{*}(\mu)\in\mathcal{P}([-1,1])$. Let now $I\subset\mathbb{R}$ with $|I|\leq\frac{1}{100}$. Then
$\mathcal{L}_{\mu}^{*}(\mu)(I)=\int\mathcal{L}_{\mu}(\mathrm{\mathbf{1}}_{I})\,d\mu.$
As $|I|\leq\frac{1}{100}$, the interval $I$ can contain only one of the intervals $I_{a}$, $a\in\mathcal{A}$, say, $a_{0}\in\mathcal{A}$, so
$\mathcal{L}_{\mu}(\mathrm{\mathbf{1}}_{I})$ $=\int|g_{\mu,a_{0}}^{\prime}(x)|^{\delta^{\mu}}\mathrm{\mathbf{1}}_{I}(g_{\mu,a_{0}}(x))\,d\mu(x)\leq\kappa^{\delta^{\mu}}\mu(g_{\mu,a_{0}}^{-1}I)$
$\leq\kappa^{\delta^{\mu}}|\kappa_{-}^{-1}I|^{\varepsilon}\leq\kappa^{\delta^{\mu}}\kappa_{-}^{-\varepsilon}|I|^{\varepsilon}\leq|I|^{\varepsilon}$
as long as $\varepsilon>0$ here is small enough and using $\delta_{\mu}\geq 1/25$ we noticed earlier in proving (1). Thus indeed $\Psi(\mu)\in\mathcal{M}_{C,\varepsilon}$.
Property (3). To check the continuity of $\Phi$, let us notice that for a continuous $h$ we can write
$\int h\,d\mathcal{L}_{\mu}^{*}(\mu)=\sum_{a\in\mathcal{A}}\int|g_{\mu,a}^{\prime}(x)|^{\delta^{\mu}}h(g_{\mu,a}(x))\,d\mu(x)$
and here if $\mu_{n}\to\mu$ the integrand $|g_{\mu_{n},a}^{\prime}(x)|^{\delta^{\mu_{n}}}h(g_{\mu_{n},a}(x))$ converges uniformly to $|g_{\mu,a}^{\prime}(x)|^{\delta^{\mu}}h(g_{\mu,a}(x))$ as $\mu$ has no atoms. This implies that $\int h\,d\mathcal{L}_{\mu_{n}}^{*}(\mu_{n})\rightarrow\int h\,d\mathcal{L}_{\mu}^{*}(\mu)$.
Property (4). The set $\mathcal{M}_{C,\varepsilon}$ is convex as for any $t\in(0,1)$ and $\mu,\nu\in\mathcal{M}_{C,\varepsilon}$ we have for any interval $I$ that
$(t\mu+(1-t)\nu)(I)\leq t|I|^{\varepsilon}+(1-t)|I|^{\varepsilon}=|I|^{\varepsilon}.$
Finally, the set $\mathcal{M}_{C,\varepsilon}$ is closed as if $\mu_{n}\to\mu$, with $\mu_{n}\in\mathcal{M}_{C,\varepsilon}$, then for any interval $I$ we also have $\mu_{n}(I)\to\mu(I)$ as $\mu$ has no atoms implying that $\mu(I)\leq C|I|^{\varepsilon}$. ∎
8. Inducing approach to Fourier decay of parabolic measures
This section is devoted to the proof of Theorem 1.6, which gives Fourier decay for some invariant measures for the non-accelerated dynamics of the Manneville-Pommeau map $T_{\alpha}$.
To be able to use Theorem 1.1 to study Theorem 1.6, we need to check that the inverse branches $g_{a}$ of $T_{A}$ indeed satisfy the axioms (1)-(3). The conditions (1) and (2) are immediate, e.g. they were deduced in [19] for Lorenz-maps but for Manneville-Pommeau maps we can for example see that $T_{A}$ satisfies the following $(C,1)$-distortion bound: there exists a constant $C>0$ such that, for all $n\geq 1$ and any word $\mathbf{a}\in\mathbb{N}^{n}$:
$\left|(S_{n}\tau\circ g_{\mathbf{a}})^{\prime}\right|\leq C,$ (8.1)
[…]
$(S_{n}\tau\circ g_{\mathbf{a}})^{\prime}=\frac{(T_{A}^{n})^{\prime\prime}\circ g_{\mathbf{a}}\cdot g_{\mathbf{a}}^{\prime}}{(T_{A}^{n})^{\prime}\circ g_{\mathbf{a}}}=\text{Adl}(T_{A}^{n})\circ g_{\mathbf{a}}.$
Now in both cases of Manneville-Pommeau or Lorenz maps, the induced map $T_{A}$ also satisfies the following Uniform Non-Integrability condition:
Lemma 8.1 (UNI).
The $T_{A}$ associated to the MP map satisfies $(c_{0},1)$-UNI on $I$.
Proof of Lemma 8.1.
Let us first show for MP maps that there exists $c_{0}\in\left(0,1\right)$ and $N\geq 1$ such that, for all $n\geq N$, there exists $\mathbf{a},\mathbf{b}\in\Sigma^{n}$ such that for all $x\in I$:
$\left|\left(S_{n}\tau\circ g_{\mathbf{a}}-S_{n}\tau\circ g_{\mathbf{b}}\right)^{\prime}\left(x\right)\right|>c_{0}.$
Note that explicitly $g_{0}=f_{0}|_{I}$, so $g_{0}$ is linear map. Consider the words
$\mathbf{a}=0^{N-1}1\quad\text{and}\quad\mathbf{b}=0^{N},$
where the power indicates that the letter is repeated that many times to make a longer word. This implies for all $N\in\mathbb{N}$ and $x\in I$ that
$\left|\left(S_{N}\tau\circ g_{\mathbf{a}}-S_{N}\tau\circ g_{\mathbf{b}}\right)^{\prime}\left(x\right)\right|=\Big|\frac{d}{dx}\Big(-\log g_{1}^{\prime}(x)+\log 2^{N-1}-(-\log 2^{N})\Big)\Big|=\Big|\frac{d}{dx}\log g_{1}^{\prime}(x)\Big|.$
Now, if $x\in A_{n}$ we have $T_{A}(x)=(f_{1}^{-1})^{n}(f_{0}^{-1}(x))$ so in particular we know that $g_{1}(x)=(f_{0}\circ f_{1})(x)$ for all $x\in I$. Thus at $x\in I$ we have
$g_{1}^{\prime}(x)$ $=\frac{1}{2}f_{1}^{\prime}(x)=\frac{1}{2}\cdot\frac{1}{1+(\alpha+1)2^{\alpha}f_{1}(x)^{\alpha}},$
and thus
$\Big|\frac{d}{dx}\log g_{1}^{\prime}(x)\Big|=\frac{\alpha(\alpha+1)2^{\alpha}f_{1}(x)^{\alpha-1}}{(1+(\alpha+1)2^{\alpha}f_{1}(x)^{\alpha})^{2}}.$
Notice finally that for all $x\in I$, we have $f_{1}(x)\in I_{1}\subset(0,1/2]$. In particular, we get the bound:
$\forall x\in I,\ \frac{\alpha(\alpha+1)2^{\alpha}f_{1}(x)^{\alpha-1}}{(1+(\alpha+1)2^{\alpha}f_{1}(x)^{\alpha})^{2}}>c_{0},$
for some $c_{0}>0$ that only depends on $\alpha$. Now, using the mean value theorem (since $S_{n}\tau\circ g_{\mathbf{a}}$ is $C^{1}$ on $I$), one can rewrite the above into a condition that does not involve any derivatives:
$\forall(x,y)\in I^{2},\ |S_{n}\tau\circ g_{\mathbf{a}}(x)-S_{n}\tau\circ g_{\mathbf{b}}(x)-S_{n}\tau\circ g_{\mathbf{a}}(y)+S_{n}\tau\circ g_{\mathbf{b}}(y)|\geq c_{0}|x-y|,$
which is exactly the bound of $(c_{0},1)$-UNI. ∎
Remark.
As for UNI condition for Lorenz maps, we can argue this for certain examples as follows. Since the map has two non-linear branches, one has to use another argument. Here the main argument to get it comes from [5]. Combining Claim 2.1 and Claim 6.1 from [5] allows us to see that the sub-IFS $\{g_{0},g_{1}\}$ for the induced system satisfies UNI if it is not $C^{2}$ conjugated to an affine IFS. Now $C^{2}$ conjugacy to an affine IFS is not possible in general, and we do this for specific numbers here to get examples of Lorenz maps. Indeed, denoting by $x_{0},x_{1},x_{10}$ the fixed points of $g_{0},g_{1},g_{10}$ respectively, we have that $\log(g_{1}\circ g_{0})^{\prime}(x_{10})=\log g_{0}^{\prime}(x_{0})+\log g_{1}^{\prime}(x_{1})$. This is not possible in general, for example with simulations for $b_{0}=-1/2$, $b_{1}=1/2$, $a=1.1$ and $\alpha=0.25$ reveals $g_{0}(x)=-0.68301(0.5-x)^{4}$ with fixed point $x_{0}=-0.07429$, $g_{0}’(x)=2.73205(0.5-x)^{3}$ with $g_{0}’(x_{0})=0.51748$, $g_{1}(x)=0.68301(0.5+x)^{4}$ with fixed point $x_{1}=0.07429$, $g_{1}’(x)=2.73205(0.5+x)^{3}$ with $g_{1}’(x_{1})=0.51748$ so the product $g_{0}’(x_{0})g_{1}’(x_{1})\approx 0.26779$. Moreover, the composition $g_{1}\circ g_{0}(x)=0.683013(0.5-0.683013(0.5-x)^{4})^{4}$ has fixed point $x_{10}=0.03259$ and the derivative $(g_{1}\circ g_{0})’(x)=7.46411(0.5-0.683013(-0.5+x)^{4})^{3}(0.5-x)^{3}$ with $(g_{1}\circ g_{0})’(x_{10})\approx 0.077825$. Thus: $|\log(g_{1}\circ g_{0})’(x_{10})-\log g_{0}’(x_{0})-\log g_{1}’(x_{1})|\approx 1.23572>1$, which is a contradiction. Thus UNI holds for $\{g_{0},g_{1}\}$. By comparing the fixed points, it is likely though that UNI should hold for all choices of $\alpha$ without using specific examples but we do not pursue this here.
Thus we are in the setting of Theorem 1.1 to have that all equilibrium states for $T_{A}$ of light-tail have power Fourier decay. Then to go back to the parabolic system, we will just have to use the inducing structure in the following:
Proof of Theorem 1.6 assuming Theorem 1.1.
Let us first write the Manneville-Pommeau map case. Firstly, since
$\nu_{\phi}(E)=\sum_{a=0}^{\infty}\sum_{k=0}^{a}\mu_{\phi}(T_{\alpha}^{-k}(E)\cap A_{a}),\quad E\subset I,$
we can write
$\int\chi(x)e^{i\xi\psi(x)}\,d\nu_{\phi}(x)=\sum_{a=0}^{\infty}\sum_{k=0}^{a}\int_{I}\chi(T_{\alpha}^{k}(x))e^{i\xi\psi(T_{\alpha}^{k}(x))}\mathrm{\mathbf{1}}_{A_{a}}(x)\,d\mu_{\phi}(x).$
Applying the transfer operator associated to $\phi$ once, we see that this equals to
$\sum_{a=0}^{\infty}\sum_{k=0}^{a}\sum_{b=0}^{\infty}\int_{I}\chi(T_{\alpha}^{k}(g_{b}(x)))e^{i\xi\psi(T_{\alpha}^{k}(g_{b}(x)))}\mathrm{\mathbf{1}}_{A_{a}}(g_{b}(x))e^{\phi(g_{b}(x))}\,d\mu_{\phi}(x).$
Notice that as $g_{b}:I\to A_{b}$, we have that $\mathrm{\mathbf{1}}_{A_{a}}(g_{b}(x))=1$ if and only if $a=b$, so we arrive to
$\sum_{a=0}^{\infty}\sum_{k=0}^{a}\int_{I}\chi(T_{\alpha}^{k}(g_{a}(x)))e^{i\xi\psi(T_{\alpha}^{k}(g_{a}(x)))}e^{\phi(g_{a}(x))}\,d\mu_{\phi}(x).$
Define now maps $\psi_{a,k}:I\to\mathbb{R}$ and $\chi_{a,k}:I\to\mathbb{R}$ by the formulae
$\psi_{a,k}(x):=\psi(T_{\alpha}^{k}(g_{a}(x))),\quad\chi_{a,k}(x):=\chi(T_{\alpha}^{k}(g_{a}(x)))\frac{e^{\phi(g_{a}(x))}}{e^{\phi(g_{a}(3/4))}},\quad x\in I.$
Then we arrive to the formula
$\sum_{a=0}^{\infty}\sum_{k=0}^{a}e^{\phi(g_{a}(3/4))}\int_{I}\chi_{a,k}(x)e^{i\xi\psi_{a,k}(x)}d\mu_{\phi}(x).$
Notice that, for any $a$, we have $T_{\alpha}^{a}(g_{a}(x))=x$. Remember also that $|T_{\alpha}^{\prime}|\geq 1$. It follows that, for any $k\leq a$, we have a bound $|(T_{\alpha}^{k}\circ g_{\mathbf{a}})^{\prime}(x)|\leq 1$. In particular, for every $k\geq a$, we have a uniform bound $\|T_{\alpha}^{k}\circ g_{a}\|_{C^{\alpha}}\lesssim 1$. It follows that we can find uniform bounds
$\|\chi_{a,k}\|_{C^{\alpha}}$ $\lesssim 1,$
$\|\log\psi_{a,k}\|_{C^{1+\alpha}}$ $\lesssim 1.$
As $\inf|\psi^{\prime}|>c$ and $|T_{\alpha}^{\prime}|\geq 1$, we have
$|\psi_{a,k}^{\prime}(x)|\geq c|g_{a}^{\prime}(x)|\geq c^{\prime}|g_{a}^{\prime}(3/4)|,$
for some constant $c^{\prime}>0$ by the bounded distortion property of $T_{A}$. We will then cut the sum over $a$ depending on if $|g_{a}^{\prime}(3/4)|\geq|\xi|^{-\varepsilon_{\psi}}$, where $\varepsilon_{\psi}$ is small enough so that Theorem 1.1 gives us a uniform decay bound. Notice also that, because of the support condition on $\chi$, there exists some integer $K(\chi)\geq 1$ such that $\chi_{a,k}=0$ for any $k\geq K(\chi)$. We then find:
$\Big|\int\chi(x)e^{i\xi\psi(x)}\,d\nu_{\phi}(x)\Big|\leq\sum_{a=0}^{\infty}\sum_{k=0}^{a}e^{\phi(g_{a}(3/4))}\Big|\int_{I}\chi_{a,k}(x)e^{i\xi\psi_{a,k}(x)}d\mu_{\phi}(x)\Big|$
$\leq\underset{|g_{a}^{\prime}(3/4)|\geq|\xi|^{-\varepsilon_{\psi}}}{\sum_{a\geq 0}}\sum_{k=0}^{a}e^{\phi(g_{a}(3/4))}\Big|\int_{I}\chi_{a,k}(x)e^{i\xi\psi_{a,k}(x)}d\mu_{\phi}(x)\Big|+\underset{|g_{a}^{\prime}(3/4)|\leq|\xi|^{-\varepsilon_{\psi}}}{\sum_{a\geq 0}}\sum_{k=0}^{a}e^{\phi(g_{a}(3/4))}\|\chi_{a,k}\|_{\infty}$
$\lesssim{\sum_{a\geq 0}}K(\chi)e^{\phi(g_{a}(3/4))}|\xi|^{-\beta}+{\sum_{a\geq 0}}K(\chi)e^{\phi(g_{a}(3/4))}\mathbf{1}_{|g_{a}^{\prime}(3/4)|\leq|\xi|^{-\varepsilon_{\psi}}}.$
$\lesssim|\xi|^{-\beta}+{\sum_{a\geq 0}}e^{\phi(g_{a}(3/4))}\mathbf{1}_{|g_{a}^{\prime}(3/4)|\leq|\xi|^{-\varepsilon_{\psi}}}$
$\lesssim|\xi|^{-\beta}+{\sum_{a\geq 0}}e^{\phi(g_{a}(3/4))}|g_{a}^{\prime}(3/4)|^{-\gamma}|\xi|^{-\gamma\varepsilon_{\psi}}\lesssim|\xi|^{-\min(\beta,\gamma\varepsilon_{\psi})},$
which concludes the proof.
For Lorenz maps $f$ the proof is same due to the formula for the pushed measure given by the same expression as the Manneville-Pommeau maps. The main difference is the choice of $3/4$ which we now choose inside the inducing region constructed in [19] in a neighbourhood of 0. ∎
Acknowledgements
We thank Amir Algom, Simon Baker, Amlan Banaji, Jonathan Fraser, Thomas Jordan, Natalia Jurga, Osama Khalil, Frédéric Naud, Tuomas Orponen, Mike Todd and Caroline Wormell for useful discussions during the preparation of this manuscript. We thank in particular Simon Baker for suggesting to consider the Liverani-Saussol-Vaienti parametrisation for the Manneville-Pommeau map for studying the UNI and its generalisation.